### v2

# Viscous-inertial waves on the surface of the Sun: modeling, forward and inverse problems

Abstract
This paper develops a mathematical framework for studying the newly discovered solar inertial oscillations, offering promising new avenues for exploring the Sun’s dynamics. Under the assumption of purely toroidal motions, the stream function of the flow satisfies a fourth-order scalar equation governing inertial waves on the rotating Sun. We prove well-posedness of wave solutions under explicit conditions on differential rotation. Moreover, we study the inverse problem of simultaneously reconstructing viscosity and differential rotation parameters from either complete or partial surface observations. To this end, we verify the tangential cone condition, ensuring the convergence of iterative regularization methods, furthermore the local unique identifiability of the unknown parameters. Numerical experiments employing the Nesterov-Landweber iteration confirm robustness of the reconstruction across different observation schemes and noise levels.
Key words.
inverse problems, differential equations on manifolds, forth order elliptic differential equations, helioseismology, inertial waves, iterative regularization, tangential cone condition
Acknowledgement.
Support from Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1456/432680300 Mathematics of Experiment, Project C04 is gratefully acknowledged.
1 Introduction
Helioseismology is the study of the internal structure and dynamics of the Sun from observations of solar oscillations on the surface. Such inference is primarily done by analyzing the observed pressure wave (p-modes) and has led to many achievements including the determination of the Sun’s differential rotation [37]. However, the sensitivity of p-modes decreases with depth exacerbating the ill-posedness of the inversion. It is this ill-posedness that motivates the present article, and its accompanying analysis of regularized recovery of rotation and viscosity on the Sun.
The recent discovery of numerous inertial modes on the surface of the Sun [15] opens new possibilities for helioseismology. Inertial oscillations, whose frequencies are in the range of hundreds of nHz, are caused by the Coriolis force, exhibiting periodic behavior on the timescale of the solar rotation cycle. As a consequence, these modes require very long observation time to be identified. Indeed, their discovery relies on 10 years of data acquisition by the Helioseismic and Magnetic Imager on board the Solar Dynamics Observatory [34]. As many of these modes have maximum energy at the base of the convection zone [15], i.e., at a depth of about 0.71 solar radii, their study promises to give new insight into the dynamics of the deeper layers of the Sun. We will use these newly discovered waves as data of the inversion process.
Most of the observed modes are quasi-toroidal, i.e. almost divergence-free, with very small radial velocity component, giving rise to the use of a 2D model for inertial waves on a rotating sphere. In this framework, we can reduce the vectorial wave equation into a scalar equation employing a stream function, the divergence-free component of the Helmholtz decomposition of the flow. This idea was proposed in [40] and extended in [7] to take into account viscous stress, an important component in solar applications [12].
Inspired by these works, we will model the propagation of inertial-viscous waves on the surface of the rotating Sun by the stream function. The vectorial motion equation driven by Coriolis force and viscous stress will be turn into a fourth-order scalar equation in a low-dimensional setting. Moreover, we derive the separated equations for each longitudinal-wavenumber $m$ with realistic and physical meaningful $m$-dependent boundary conditions; this serves as the underlying model for the inverse parameter identification problem.
The problem of solar inversion using inertial waves is completely novel as these modes have been discovered only recently [15]. Here, we take a first step towards recovering the surface differential rotation and the viscosity from Dopplergrams. Doppler data are usually not used directly in helioseismology due to the high level of convective noise; indeed, solar excitations can be viewed as random turbulence, resulting in random oscillations. However, some inertial modes have been detected directly in the Doppler signal [14] motivating this study using direct measurements of inertial waves as input data.
Helioseismology predominantly solves linear inverse problem, under the assumption that perturbations are relatively small, as it is the case, for example, for meridional flows [13]. However, for parameters such as viscosity, where no suitable initial guess is available, this linearization assumption may no longer be valid. Here, following a related approach for pressure waves and correlation data in [27], we solve the nonlinear inversion for rotation and viscosity using inertial waves and iterative regularization reconstruction methods.
Contributions.
This article provides a complete treatment – modeling, forward problem and inverse problem – for the newly observed inertial waves on the Sun:
After a rigorous derivation of the forward model already used in [12], we prove existence, uniqueness and stability of both, the full and separated the inertial wave equations. This well-posedness is derived under explicit conditions on the relative smallness of variations of the the latitudinal-dependent rotation $\Omega$ compared to the product of the frequency and the third power of the viscous parameter $\gamma$. In passing, these results also allow us to apply analytic Fredholm theory, providing mathematical foundations to the recent study of inertial modes. The results of our analysis, both on the conditions for uniqueness and on the isolated occurrence of inertial modes, agree with state-of-the-art interpretations in inertial mode helioseismology.
For the inverse problem of the simultaneous identification of viscosity and differential rotation, we lay out foundations of iterative regularization methods by characterizing adjoint operators on a continuous level. Crucially, we provide convergence guarantees for these methods – for full and, to a certain extent, also for restricted observations. We achieve this by establishing the important tangential cone condition via a lifted regularity strategy, accounting for realistic $L^{2}$-data. In addition, we prove local unique identifiability of $\Omega$ when $\gamma$ is known, and vice versa, from full measurement.
Structure.
Section 2 introduces the modeling framework for inertial oscillations and outlines different observation strategies for solar data. The well-posedness and regularity of the resulting wave solutions are derived in Section 3. Section 4 addresses the inverse problem for viscosity and differential rotation, including sensitivity analysis and derivation of associated adjoint operators. In Section 5, we establish convergence guarantees of iterative reconstruction schemes and local unique identifiability of the unknown parameters based on the tangential cone condition. Finally, Section 6 presents numerical reconstruction results under different scenarios, and Section 7 concludes our findings with a summary and an outlook on future research directions.
2 Model and observation of inertial waves
2.1 Reduced-order modeling of inertial waves
In this section, we derive a physical model for dynamics of waves in the presence of latitudinal differential rotation and eddy viscosity on the Sun in a rotating frame.
Linearized Navier Stokes equations.
We begin with the equation of motion (momentum equation), in which the dynamics of a moving particle of density $\rho$ in an incompressible fluid is subject to linear viscous stress $\mathbf{\bm{\tau}}$ with viscosity $\gamma$, acoustic pressure $p$, gravity $\mathbf{g}$, and external source $\mathbf{f}$:
$\begin{aligned} &\rho\left(\frac{\partial\textbf{{v}}}{\partial t}+\nabla_{\!\textbf{{v}}}\textbf{{v}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{v}}\right)=\operatorname{\mathbf{div}}(\rho\gamma\mathbf{\bm{\tau}})-\operatorname{\mathbf{grad}}p+\rho\mathbf{g}+\rho\mathbf{f},\\ &\mathbf{\bm{\tau}}:=\operatorname{\mathbf{grad}}\textbf{{v}}+(\operatorname{\mathbf{grad}}\textbf{{v}})^{\top}.\end{aligned}$ (1a)
Here, $\textbf{{v}}=\textbf{{v}}(\textbf{{r}},t)\in\mathbb{R}^{3}$ is the vector velocity of a particle at position $\textbf{{r}}\in\mathbb{R}^{3}$ and time $t\in(0,T)$, while $\nabla_{\!\textbf{{v}}}:=\sum_{i=1}^{3}v_{i}\frac{\partial}{\partial x_{i}}$ applies componentwise. We work in the frame rotating at the angular frequency $\Omega_{\mathrm{{ref}}}$ around a fixed axis $\mathbf{\hat{a}}$ and set $\mathbf{\Omega}_{\mathrm{ref}}=\Omega_{\mathrm{{ref}}}\,\mathbf{\hat{a}}$. Equation (1a) corresponds to the Navier-Stokes equation in a rotating frame. The fictitious force $(2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{v}})$ is the inertial Coriolis force resulting from the transformation between the fixed and the rotating frame. The other additional fictitious forces, Euler force $\rho\left(\frac{d\mathbf{\Omega}_{\mathrm{ref}}}{dt}\times\textbf{{r}}\right)$ and centrifugal force $\rho\mathbf{\Omega}_{\mathrm{ref}}\times(\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{r}})$, are ignored as $\Omega_{\mathrm{{ref}}}$ is uniform in time and assumed to be small. For solar applications, one usually chooses the Carrington frame as reference frame with the angular velocity $\Omega_{\mathrm{{ref}}}=14.7^{0}/$day and the rotation axis $\mathbf{e}_{z}=[0,0,1]^{\top}$. Equation (1a) is complemented by the continuity equation (conservation of mass):
$\frac{\partial\rho}{\partial t}+\operatorname{div}(\rho\textbf{{v}})=0\,.$ (1b)
We linearize Eqs. (1) around a stationary background (unperturbed, equilibrium, mean) medium characterized by $\textbf{{u}}_{0}$}, $p_{0}$, $\rho_{0}$, and $\mathbf{g}_{0}$ such that
$\textbf{{v}}(\textbf{{r}},t)={\textbf{{u}}_{0}}(\textbf{{r}})+\textbf{{u}}(\textbf{{r}},t),\quad\rho(\textbf{{r}},t)=\rho_{0}(\textbf{{r}})+\rho^{\prime}(\textbf{{r}},t),$
$p(\textbf{{r}},t)=p_{0}(\textbf{{r}})+p^{\prime}(\textbf{{r}},t),\quad\mathbf{g}(\textbf{{r}},t)=\mathbf{g}_{0}(\textbf{{r}})+\mathbf{g}^{\prime}(\textbf{{r}},t),$
where u is the (perturbed) wave velocity, and ${\textbf{{u}}_{0}}(\textbf{{r}}):=(\mathbf{\Omega}(\textbf{{r}})-\mathbf{\Omega}_{\mathrm{ref}})\times\textbf{{r}}$ is assumed to be a differential rotation field with angular velocity $\mathbf{\Omega}(\textbf{{r}})=\Omega(r,\theta)\mathbf{e}_{z}$ depending only on radius $r$ and colatitude $\theta$ in spherical coordinates $r;\theta;\phi]\in[0,\infty)\times[0,\pi)\times[0,2\pi)$. We assume that the background medium solves the second-order quasilinear elliptic boundary value problem
$-\operatorname{\mathbf{div}}\left(\gamma\rho_{0}\left(\operatorname{\mathbf{grad}}{\textbf{{u}}_{0}}+(\operatorname{\mathbf{grad}}{\textbf{{u}}_{0}})^{\top}\right)\right)+\rho_{0}\nabla_{\!{\textbf{{u}}_{0}}}{\textbf{{u}}_{0}}+2\rho_{0}\mathbf{\Omega}_{\mathrm{ref}}\times{\textbf{{u}}_{0}}$
$\hskip 256.0748pt=-\operatorname{\mathbf{grad}}p_{0}+\rho_{0}\mathbf{g}_{0},$ (2a)
$\operatorname{div}(\rho_{0}{\textbf{{u}}_{0}})=0.$ (2b)
The background gravity force derives from a potential $\mathbf{g}_{0}=-\nabla\Phi_{0}$ that satisfies the Poisson equation $\Delta\Phi_{0}=4\pi G\rho_{0}$, where $G$ is the gravitational constant. To construct such a background model for the Sun, it is usually assumed that the background flows are weak and the background medium is built from the hydrostatic equilibrium, that is $\operatorname{\mathbf{grad}}p_{0}=\rho_{0}\mathbf{g}_{0}$. This assumption leads to background coefficients that depend only on depth and are given by standard solar models such as model S [8]. Starting from a radial model for density, it is possible to solve Poisson equation to obtain the gravitational potential and deduce the background pressure.
Assumption:
Waves have small amplitude in comparison to the mean flow, i.e. $|\textbf{{u}}|\ll|{\textbf{{u}}_{0}}|$. Moreover, the anelastic approximation $\rho^{\prime}\approx 0$ and the Cowling approximation $\mathbf{g}^{\prime}\approx 0$ (see [9]) hold true.
The anelastic approximation is justified if the flow speed is significantly smaller than the sound speed, which is adequate for inertial modes – a few meters per second versus kilometers per second for the surface sound speed. The system (1) thus becomes
$\rho_{0}\left(\frac{\partial\textbf{{u}}}{\partial t}+\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}+\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{u}}\right)=-\operatorname{\mathbf{grad}}p^{\prime}+\operatorname{\mathbf{div}}\left[\rho_{0}\gamma\bm{\tau}^{\prime}\right]+\rho_{0}\mathbf{f},$ (3a)
$\operatorname{div}(\rho_{0}\textbf{{u}})=0,$ (3b)
where $\bm{\tau}^{\prime}:=\operatorname{\mathbf{grad}}\textbf{{u}}+(\operatorname{\mathbf{grad}}\textbf{{u}})^{\top}$.
The flow field $\textbf{{u}}_{0}$}, equivalently $\Omega$, and the viscosity $\gamma$ will be the quantities of interest in the inverse problem.
Stratified equations.
We first rewrite the left-hand side of (3a) in spherical coordinates, i.e. in terms of the components $u_{r},u_{\theta},u_{\phi}$ of the expansion $\textbf{{u}}=u_{r}\mathbf{e}_{r}+u_{\theta}\mathbf{e}_{\theta}´+u_{\phi}\mathbf{e}_{\phi}$ into the local orthonormal basis vectors $\mathbf{e}_{r},\mathbf{e}_{\theta}$, and $\mathbf{e}_{\phi}$ as defined in the appendix A. Noting that ${\textbf{{u}}_{0}}=(\Omega(r,\theta)-\Omega_{\mathrm{{ref}}})\sin\theta\mathbf{e}_{\phi}$ and using the identities
$\mathbf{\Omega}\times\textbf{{u}}=\Omega\left(-\sin\theta\,u_{\phi}\mathbf{e}_{r}-\cos\theta\,u_{\phi}\mathbf{e}_{\theta}+(\sin\theta\,u_{r}+\cos\theta\,u_{\theta})\mathbf{e}_{\phi}\right),$
$\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}=(\Omega-\Omega_{\mathrm{{ref}}})(r,\theta)\frac{\partial\textbf{{u}}}{\partial\phi}+(\mathbf{\Omega}-\mathbf{\Omega}_{\mathrm{ref}})\times\textbf{{u}},$
$\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}=(\mathbf{\Omega}-\mathbf{\Omega}_{\mathrm{ref}})\times\textbf{{u}}+\left(\frac{\partial\Omega}{\partial r}(r,\theta)r\sin\theta\,u_{r}+\frac{\partial\Omega}{\partial\theta}(r,\theta)\sin\theta\,u_{\theta}\right)\mathbf{e}_{\phi},$
we obtain
$\left[\mathbf{e}_{r};\mathbf{e}_{\theta};\mathbf{e}_{\phi}\right]^{\top}\left(\frac{\partial\textbf{{u}}}{\partial t}+\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}+\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{u}}\right)$
$=$ $\mathbf{D}_{t}\begin{bmatrix}u_{r}\\ u_{\theta}\\ u_{\phi}\end{bmatrix}+\begin{bmatrix}0&0&-2\Omega\sin\theta\\ 0&0&-2\Omega\cos\theta\\ \left(2\Omega\sin\theta+\frac{\partial\Omega}{\partial r}r\sin\theta\right)&\left(2\Omega\cos\theta+\frac{\partial\Omega}{\partial\theta}\sin\theta\right)&0\end{bmatrix}\begin{bmatrix}u_{r}\\ u_{\theta}\\ u_{\phi}\end{bmatrix}$, (4)
where $\mathbf{D}_{t}:=\frac{\partial}{\partial t}+(\Omega-\Omega_{\mathrm{{ref}}})\frac{\partial}{\partial\phi}$ is the material derivative.
We now introduce the $3\times 2$-matrix $\mathbf{R}(\textbf{{r}}):=[\mathbf{e}_{\theta}(\textbf{{r}}),\mathbf{e}_{\phi}(\textbf{{r}})]$. Then multiplication by the $3\times 3$ matrix $\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}$ is the projection onto the tangential subspace and equivalent to the application of $\textbf{{r}}\times\textbf{{r}}\times$. By applying $\textbf{{r}}\times\textbf{{r}}\times$ to (3a), we wish to obtain an approximate equation for the horizontal velocity components $\textbf{{u}}_{\mathrm{h}}(\textbf{{r}}):=-\textbf{{r}}\times\textbf{{r}}\times\textbf{{u}}(\textbf{{r}})$ given by $\textbf{{u}}_{\mathrm{h}}=\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}\textbf{{u}}=u_{\theta}\mathbf{e}_{\theta}+u_{\phi}\mathbf{e}_{\phi}$. To this end, we need the following assumption:
Assumption:
Viscosity and density are radially symmetric, i.e. $\gamma=\gamma(r)$ and $\rho_{0}=\rho_{0}(r)$. The fluid is strongly stratified in the sense that the radial motion $u_{r}$ and its derivatives are small compared to the horizontal components $u_{\phi},u_{\theta}$.
Due to this assumption, we can neglect the first column of the matrix in Eq. (4), and the application of $\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}$ “kills” the first row of this matrix (the first element of the vector in (4) may be balanced by $\frac{\partial p^{\prime}}{\partial r}$). Therefore, we are left with the matrix
$\mathbf{M}(\textbf{{r}}):=\mathbf{R}(\textbf{{r}})\begin{bmatrix}0&-2\Omega(r,\theta)\cos\theta\\ 2\Omega(r,\theta)\cos\theta+\frac{\partial\Omega}{\partial\theta}(r,\theta)\sin\theta&0\end{bmatrix}\mathbf{R}(\textbf{{r}})^{\top}$
and obtain the following system of equations for the tangential vector field $\textbf{{u}}_{\mathrm{h}}$ on $r{\mathbb{S}^{2}}$} for a fixed $r>0$ with right hand side $\textbf{{f}}_{\!\mathrm{h}}:=\mathbf{R}\mathbf{R}^{\top}\textbf{{f}}$:
$-\rho_{0}\gamma\bm{\Delta}_{\mathrm{h}}\textbf{{u}}_{\mathrm{h}}+\rho_{0}\mathbf{D}_{t}\textbf{{u}}_{\mathrm{h}}+\rho_{0}\mathbf{M}\textbf{{u}}_{\mathrm{h}}=\operatorname{\mathbf{grad}_{\mathrm{h}}}p^{\prime}+\rho_{0}\textbf{{f}}_{\!\mathrm{h}},$ (5a)
$\operatorname{div_{\mathrm{h}}}\textbf{{u}}_{\mathrm{h}}=0.$ (5b)
employing the fact $\operatorname{\mathbf{grad}}(\rho_{0}(\operatorname{\mathbf{grad}}\textbf{{u}}_{\mathrm{h}})^{\top})=\operatorname{\mathbf{grad}_{\mathrm{h}}}(\rho_{0}\operatorname{div_{\mathrm{h}}}\textbf{{u}}_{\mathrm{h}})=0$ due to (5b).
Physically, the assumption of small radial motions is motivated by an identification of the observed inertial modes from [15] as the eigenvalues of the 3D linearized equations of momentum, mass and energy [5]. In this setup, the radial velocities of the eigenfunctions are small compared to the horizontal ones for most of the observed modes. This is also the case in nonlinear simulations of rotating convection in a spherical shell [6]. This assumption could possibly be justify mathematically within an asymptotic analysis as $r\to\infty$ or $\Omega\to 0$. This, however, would require to amend the system (3) by an atmospheric model and assumptions on the coefficient functions and is beyond the scope of this paper.
Reformulation as scalar fourth-order equation.
It follows from the Helmholtz decomposition (53) and Eq. (5b) that there exits a stream function $\Psi$ such that
$\rho_{0}\textbf{{u}}_{\mathrm{h}}=\operatorname{\mathbf{curl}_{\mathrm{h}}}\Psi.$
In order to obtain an equation for $\Psi$, we apply $\operatorname{curl_{\mathrm{h}}}$ to (5a). Noting that $\operatorname{curl_{\mathrm{h}}}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}\Psi=-\Delta_{\mathrm{h}}\Psi$ and using the identities
$\operatorname{curl_{\mathrm{h}}}\mathbf{D}_{t}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=-D_{t}\Delta_{\mathrm{h}}\Psi-\frac{1}{r^{2}}\frac{\partial\Omega}{\partial\theta}\frac{\partial^{2}\Psi}{\partial\theta\,\partial\phi},$
$\operatorname{curl_{\mathrm{h}}}\rho_{0}\gamma\bm{\Delta}_{\mathrm{h}}\textbf{{u}}_{\mathrm{h}}=-\gamma(\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}})^{2}\Psi=-\gamma\Delta_{\mathrm{h}}^{2}\Psi$ (6)
$\operatorname{curl_{\mathrm{h}}}\mathbf{M}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}+\frac{1}{r^{2}}\frac{\partial\Omega}{\partial\theta}\frac{\partial^{2}\Psi}{\partial\theta\,\partial\phi},\quad\alpha_{\Omega}(\theta):=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\right)$
with the analog $D_{t}:=\frac{\partial}{\partial t}+(\Omega-\Omega_{\mathrm{{ref}}})\frac{\partial}{\partial\phi}$ of $\mathbf{D}_{t}$, we arrive at the scalar fourth-order equation
$\gamma\Delta_{\mathrm{h}}^{2}\Psi-D_{t}\Delta_{\mathrm{h}}\Psi+\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}=f$ (7)
with $f:=\rho_{0}\operatorname{curl_{\mathrm{h}}}\textbf{{f}}_{\!\mathrm{h}}$, and $\alpha_{\Omega}$ is a linear function of $\Omega$ as in (2.1) for convenience.
If we are looking for time-harmonic solutions $\Psi(t,\textbf{{r}})=\Re\left(e^{i\omega t}\Psi_{\omega}(\textbf{{r}})\right)$, then the space-dependent part has to satisfy the fourth-order elliptic equation
$\gamma\Delta_{\mathrm{h}}^{2}\Psi+i\omega\Delta_{\mathrm{h}}\Psi-\beta_{\Omega}\Delta_{\mathrm{h}}\frac{\partial\Psi}{\partial\phi}+\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}=f$ (8)
with $\beta_{\Omega}(\theta):=\Omega(r,\theta)-\Omega_{\rm ref}$.
2.2 Separated equation as inversion model
Expanding $\Psi_{\omega}(\theta,\phi)=\frac{1}{\sqrt{2\pi}}\sum_{m=-\infty}^{\infty}\widehat{\Psi}_{\omega,m}(\theta)e^{im\phi}$ into a Fourier series in azimuth, the coefficients $\widehat{\Psi}_{\omega,m}$ satisfy the ordinary boundary value problems
$\begin{cases}\gamma\Delta_{m}^{2}\widehat{\Psi}_{\omega,m}(\theta)+i\omega\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)-im\beta_{\Omega}(\theta)\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)+im\alpha_{\Omega}(\theta)\widehat{\Psi}_{\omega,m}(\theta)=\hat{f}_{m}(\theta)\\ (\Gamma_{m}\widehat{\Psi}_{\omega,m})(\theta)=0,\qquad\theta\in P:=\{0,\pi\}\end{cases}$ (9)
with $\Delta_{m}$ defined by (58) in Appendix A, and the following boundary value operators derived in Appendix B
$\Gamma_{0}\Psi:=\begin{pmatrix}\Psi^{\prime}|_{P}\\ \Psi^{\prime\prime\prime}|_{P}\end{pmatrix},\quad\Gamma_{\pm 1}\Psi):=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime\prime}|_{P}\end{pmatrix},\quad\Gamma_{m}\Psi:=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime}|_{P}\end{pmatrix},\;|m|\geq 2.$ (10)
We remark that this equation reduces to Eq. (2.9) in [40] if $\gamma=0$ and to Eq. (8) in [12] if $\hat{f}_{m}=0$. However, vanishing Cauchy data are imposed at the poles in [12], which for $|m|\leq 1$ may yield solutions that do not correspond to the unseparated equation (8).
The system (9)-(10) is the underlying model for the parameter inversion problem considered in Section 4. Solving this inverse problem requires measured data, which is the newly discovered inertial waves in this study. The acquisition process will be detailed in the forthcoming section.
2.3 Observation of solar oscillations
The basic input of helioseismology are measurements of line-of-sight velocity based on Doppler shifts of certain spectral lines. These so-called Dopplergrams acquired since 1996 by ground- or space-based telescopes have high-spatial resolution, of at least $1000\times 1000$ pixels, and a temporal cadence of at least $1$ minute. A Fourier transform of the Dopplergram data $v_{\rm dop}(\theta,\phi,t)$ in longitude $\phi$ and time $t$ yields data $v_{\rm dop}^{m,\omega}(\theta)$. Due to the limited data coverage, as only at most half of the Sun is visible at a given time, the transform in $\phi$ can only be integrated on $0,\pi]$ instead of $0,2\pi]$, leading to a leakage of the signal; see, e.g. the review in [4].
Moreover an operator should be applied to convert the simulated $\hat{\Psi}$ into $v_{\rm dop}$ using for example that $v_{\rm dop}=\mathbf{l}\cdot\mathbf{u}$, where $\mathbf{l}$ is the line-of-sight vector. On a formal point of view, this could be written as a (linear) operator that transforms $\Psi$ onto $v_{\rm dop}$. We will not consider this aspect in this article, simply assuming that reprocessed signal $y^{\delta}$ is available and described by
$y^{\delta}(\theta)=\Psi(\theta)+\mathrm{noise}(\theta)\qquad\theta\in I:=(0,\pi)\quad\text{or}\quad\theta\in I^{\prime}:=(\epsilon,\pi-\epsilon),\epsilon\geq 0,$ (11)
here $\delta=\|\mathrm{noise}\|$ denotes the noise level. The case $\epsilon=0$ corresponds to full surface coverage, while $\epsilon>0$ takes into account that higher latitudes cannot easily be observed due to the line-of-sight projection.
It is usually not possible to work with the Doppler signal directly due to the high level of convective noise. However, some inertial modes of large amplitudes have been directly detected [14], enabling us to consider the recovery of viscosity and rotation from the measurements of $y^{\delta}$.
3 Inertial waves on the manifold $r{\mathbb{S}^{2}}$}
We first show the well-posedness of Eq. (8) associated to the differential operator
$\mathbf{B}(\Psi):=\gamma\Delta_{\mathrm{h}}^{2}\Psi+i\omega\Delta_{\mathrm{h}}\Psi-\beta_{\Omega}\Delta_{\mathrm{h}}\frac{\partial\Psi}{\partial\phi}+\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}$ (12)
with recalling that $\beta_{\Omega}(\theta)=\Omega(\theta)-\Omega_{\mathrm{ref}}$ and $\alpha_{\Omega}(\theta)=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\big{(}\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\big{)}$. We will work in the space $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ of functions in $H^{2}({r{\mathbb{S}^{2}}})$ with zero mean introduced in the Appendix A.
By partial integration with taking into account symmetry of the bi-Laplacian and Laplacian operators (see Appendix, Lemma 19), the corresponding sesquilinear form $\mathbf{A}:H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\times H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\to\mathbb{C}$ defined by $\mathbf{A}[\Psi,\psi]:=\langle\mathbf{B}(\Psi),\psi\rangle_{H_{\diamond}^{-2},H_{\diamond}^{2}}$ is given by
$\begin{aligned} \mathbf{A}[\Psi,\psi]&=\gamma\left\langle\Delta_{\mathrm{h}}\Psi,\Delta_{\mathrm{h}}\psi\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\Psi,\operatorname{\mathbf{grad}_{\mathrm{h}}}\psi\right\rangle+(\mathbf{A}^{\alpha}+\mathbf{A}^{\beta})[\Psi,\psi],\\ &\mathbf{A}^{\alpha}[\Psi,\psi]:=-\left\langle\alpha_{\Omega}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle,\qquad\mathbf{A}^{\beta}[\Psi,\psi]:=\left\langle\beta_{\Omega}\Delta_{\mathrm{h}}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle.\end{aligned}$ (13)
Since the coefficients $\alpha_{\Omega}$ and $\beta_{\Omega}$ are independent of $\phi$, this sesquilinear form is block diagonal with respect to the decomposition $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus_{m\in Z}^{\infty}V_{m}$ with the subspaces
$V_{m}:=H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$
introduced in (60). In other words, we have $\mathbf{A}(\Psi_{m},\Psi_{n})=0$ for $\Psi_{m}\in V_{m}$, $\Psi_{n}\in V_{n}$ and $m\neq n$. Hence, the restrictions $\mathbf{A}_{m}:=\mathbf{A}|_{V_{m}\times V_{m}}$ fulfill
$\mathbf{A}(\Psi,\psi)=\sum_{m\in\mathbb{Z}}\mathbf{A}_{m}(\Psi_{m},\psi_{m})$ (14)
if $\Psi=\sum_{m\in\mathbb{Z}}\Psi_{m}$ and $\psi=\sum_{m\in\mathbb{Z}}\psi_{m}$ with $\Psi_{m},\psi_{m}\in V_{m}$. We will also analyze the separated sesquilinear forms given by
$\mathbf{A}_{m}[\Psi_{m},\psi_{m}]$ (15)
$\quad=\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi_{m},\operatorname{\mathbf{grad}_{\mathit{m}\!}}\psi_{m}\right\rangle+(\mathbf{A}^{\alpha}_{m}+\mathbf{A}^{\beta}_{m})[\Psi_{m},\psi_{m}],$
$\mathbf{A}^{\alpha}_{m}[\Psi_{m},\psi_{m}]:=im\left\langle\alpha_{\Omega}\Psi_{m},\psi_{m}\right\rangle,\qquad\;\mathbf{A}^{\beta}_{m}[\Psi_{m},\psi_{m}]:=-im\left\langle\beta_{\Omega}\Delta_{m}\Psi_{m},\psi_{m}\right\rangle.$
with $\mathbf{A}^{\alpha}_{m},\mathbf{A}^{\beta}_{m}:V_{m}\times V_{m}\to\mathbb{C}$ as restrictions of $\mathbf{A}^{\alpha}$, $\mathbf{A}^{\beta}$ defined in (13). Here, $\Delta_{m}$, $\operatorname{\mathbf{grad}_{\mathit{m}\!}}$ are the decomposed operators as defined in (58), (59) in Appendix A.
3.1 Boundedness
It is straightforward to show that $\mathbf{A}^{\alpha}$ and $\mathbf{A}^{\alpha}_{m}$ are bounded if $\Omega\in W^{2,\infty}(0,\pi)$. However, we wish to reduce the regularity assumption on the unknown rotation and only assume
$\Omega\in H^{1}({r{\mathbb{S}^{2}}};m=0)\,.$ (16)
Under this reduced regularity assumption, establishing boundedness requires additional technical analysis. For later use in $\mathrm{G\mathring{a}rding}$ inequality, we also provide a bound by a weaker norm. In the following, we will frequently employ continuous embeddings between function spaces $X,Y$ denoted the embedding constants $C_{X\to Y}$, i.e., $\|\cdot\|_{Y}\leq C_{X\to Y}\|\cdot\|_{X}$.
Lemma 1.
If $\Omega\in C^{2}([0,\pi])$, then for any $m\in\mathbb{Z}$ and $\Psi_{m},\psi_{m}\in V_{m}\cap C^{\infty}({r{\mathbb{S}^{2}}})$ we have
$\mathbf{A}^{\alpha}[\Psi_{m},\psi_{m}]$ $=\left\langle\tilde{\alpha}_{\Omega}\operatorname{\mathbf{grad}_{\mathrm{h}}}\Psi_{m},\operatorname{\mathbf{curl}_{\mathrm{h}}}\psi_{m}\right\rangle$ (17)
where $\tilde{\alpha}_{\Omega}(\theta):=\Omega^{\prime}(\theta)\sin\theta+2\Omega(\theta)\cos\theta$. Moreover,
$\left|\mathbf{A}^{\alpha}[\Psi_{m},\psi_{m}]\right|\leq C\|\Omega\|_{H^{1}}\|\Psi_{m}\|_{H^{2}}\|\psi_{m}\|_{H^{3/2}}\leq C\|\Omega\|_{H^{1}}\|\Psi_{m}\|_{H^{2}}\|\psi_{m}\|_{H^{2}}$ (18)
where $C:=3C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}}$.
Under the regularity assumption (16), the right hand side of (17) has a unique continuous extension to a bounded sesquilinear form on $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$. And the inequality (18) is satisfied for all $\Psi,\psi\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$, that is,
$\left|\mathbf{A}^{\alpha}[\Psi,\psi]\right|\leq C\|\Omega\|_{H^{1}}\|\Psi\|_{H^{2}}\|\psi\|_{H^{3/2}}\leq C\|\Omega\|_{H^{1}}\|\Psi\|_{H^{2}}\|\psi\|_{H^{2}}.$ (19)
Proof.
Let $\Psi_{m},\psi_{m}\in V_{m}\cap C^{\infty}({r{\mathbb{S}^{2}}})$. Note that $\alpha_{\Omega}(\theta)=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\tilde{\alpha}_{\Omega}(\theta)$. As $\tilde{\alpha}_{\Omega}\in C^{1}([0,\pi])$, $\Psi_{0}(\theta)\frac{\partial\overline{\psi_{0}}}{\partial\phi}(\theta)=0$ and $\Psi_{m}(\theta)\frac{\partial\overline{\psi_{m}}}{\partial\phi}(\theta)=-im\Psi_{m}(\theta)\overline{\psi_{m}}(\theta)=0$ for $\theta\in\{0,\pi\}$, $m\neq 0$, we can perform a partial integration in $\theta$ to obtain
$\mathbf{A}^{\alpha}[\Psi_{m},\psi_{m}]$
$=\int_{0}^{2\pi}\!\int_{0}^{\pi}\tilde{\alpha}_{\Omega}\frac{\partial}{\partial\theta}\left(\Psi_{m}\frac{\overline{\psi_{m}}}{\partial\phi}\right)\,d\theta\,d\phi=\int_{0}^{2\pi}\!\int_{0}^{\pi}\tilde{\alpha}_{\Omega}\left(\Psi_{m}\frac{\partial^{2}\overline{\psi_{m}}}{\partial\phi\partial\theta}+\frac{\partial\Psi_{m}}{\partial\theta}\frac{\partial\overline{\psi_{m}}}{\partial\phi}\right)\,d\theta\,d\phi$
$=\int_{0}^{2\pi}\!\int_{0}^{\pi}\tilde{\alpha}_{\Omega}\left(-\frac{\partial\Psi_{m}}{\partial\phi}\frac{\partial\overline{\psi_{m}}}{\partial\theta}+\frac{\partial\Psi_{m}}{\partial\theta}\frac{\partial\overline{\psi_{m}}}{\partial\phi}\right)\,d\theta\,d\phi$ (20)
$=\int_{{r{\mathbb{S}^{2}}}}\frac{\tilde{\alpha}_{\Omega}(\theta)}{r^{2}\sin\theta}\left(-\frac{\partial\Psi_{m}}{\partial\phi}\frac{\partial\overline{\psi_{m}}}{\partial\theta}+\frac{\partial\Psi_{m}}{\partial\theta}\frac{\partial\overline{\psi_{m}}}{\partial\phi}\right)\,ds,$
where in the second line we have performed a further partial integration in $\phi$. By using (49), we then achieve (17). Next, the generalized Hölder’s inequality
$\left|\int_{{r{\mathbb{S}^{2}}}}abc\,ds\right|\leq\|a\|_{L^{2}}\|b\|_{L^{3}}\|c\|_{L^{6}}$ (21)
implies
$\left|\mathbf{A}^{\alpha}[\Psi_{m},\psi_{m}]\right|\leq\|\tilde{\alpha}_{\Omega}\|_{L^{2}}\|\operatorname{\mathbf{grad}_{\mathrm{h}}}\Psi_{m}\|_{L^{3}}\|\operatorname{\mathbf{grad}_{\mathrm{h}}}\psi_{m}\|_{L^{6}}.$
The bound $\|\tilde{\alpha}_{\Omega}\|_{L^{2}}\leq\|\frac{\partial\Omega}{\partial\theta}\|_{L^{2}}+2\|\Omega\|_{L^{2}}\leq 3\|\Omega\|_{H^{1}}$ together with the continuity of the embeddings $H^{1}({r{\mathbb{S}^{2}}})\hookrightarrow L^{6}({r{\mathbb{S}^{2}}})$ and $H^{1/2}({r{\mathbb{S}^{2}}})\hookrightarrow L^{3}({r{\mathbb{S}^{2}}})$ in (55b) yields (18).
Since $\mathbf{A}^{\alpha}$ separates in $V_{m}$ and the spaces $V_{m}$ are orthogonal in $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$, the bound (18) also holds true for finite linear combinations of $\Psi_{m}$ and $\psi_{m}$ of the form above. The last statement (19) follows from the density of such linear combinations (or even linear combinations of spherical hamonics) in $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$. ∎
Proposition 2.
Suppose that $\Omega$ satisfies (16). Then the sesquilinear form $\mathbf{A}$ defined by (13) is bounded on $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$, and the separated bilinear forms $\mathbf{A}_{m}$ defined by (15) are bounded on $V_{m}$. Both $\mathbf{A}$ and $\mathbf{A}_{m}$ depend continuously on $\Omega$ with respect to its $H^{1}$-norm.
Proof.
Boundedness of $\mathbf{A}^{\alpha}$ has been shown in Lemma 1. For $\mathbf{A}^{\beta}$, we again use the Hölder inequality (21) and the same Sobolev embeddings as above to obtain
$\begin{split}|\mathbf{A}^{\beta}(\Psi,\psi)|&\leq\|\beta\|_{L^{6}}\|\Delta_{\rm h}\Psi\|_{L^{2}}\|\operatorname{\mathbf{grad}_{\mathrm{h}}}\psi\|_{L^{3}}\\ &\leq C_{H^{1}\to L^{6}}C_{H^{1/2}\to L^{3}}\|\Omega-\Omega_{\mathrm{{ref}}}\|_{H^{1}}\|\Psi\|_{H^{2}}\|\psi\|_{H^{3/2}}.\end{split}$ (22)
Boundedness of the first two terms in (13) is obvious, and boundedness of $\mathbf{A}_{m}$ follows from that of $\mathbf{A}$, using (14) and the pairwise orthogonality of the spaces $V_{m}$. ∎
3.2 Fredholm property and well-posedness
Theorem 3.
Let $\gamma>0$ and $\Omega\in H^{1}({r{\mathbb{S}^{2}}};m=0)$ as in Proposition 2.
  1. 1.
    The differential operator $\mathbf{B}:H^{2}_{\diamond}({r{\mathbb{S}^{2}}})\to H^{-2}_{\diamond}({r{\mathbb{S}^{2}}})$ defined in (12) is Fredholm of index 0.
  2. 2.
    Moreover, $\mathbf{B}$ is boundedly invertible at sufficiently large frequencies satisfying
    $|\omega|>\frac{4}{\gamma^{3}}(C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}})^{4}\big{(}\|\Omega-\Omega_{\mathrm{{ref}}}\|^{2}_{H^{1}}+9\|\Omega\|^{2}_{H^{1}}\big{)}^{2}.$ (23)
Proof.
[…]
$\sqrt{2}|\mathbf{A}[\Psi,\Psi]|\geq|\Re\{\mathbf{A}[\Psi,\Psi]\}|+|\Im\{\mathbf{A}[\Psi,\Psi]\}|$
$\geq\gamma\|\Delta_{\mathrm{h}}\Psi\|^{2}_{L^{2}}+|\omega|\|\operatorname{\mathbf{grad}_{\mathrm{h}}}\Psi\|^{2}_{L^{2}}-|\mathbf{A}^{\alpha}[\Psi,\Psi]|-|\mathbf{A}^{\beta}[\Psi,\Psi]|$
$\geq\frac{3\gamma}{4}\|\Psi\|^{2}_{H_{\diamond}^{2}}+|\omega|\|\Psi\|^{2}_{H_{\diamond}^{1}}-\frac{2}{\gamma}(C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}})^{2}\big{(}\|\Omega-\Omega_{\mathrm{{ref}}}\|^{2}_{H_{\diamond}^{1}}+9\|\Omega\|^{2}_{H_{\diamond}^{1}}\big{)}\|\Psi\|^{2}_{H_{\diamond}^{3/2}}.$
Since the embeddings $H^{1}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ and $H^{3/2}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ are compact, $\mathbf{B}$ is a Fredholm operator of index 0; see, e.g., [33, Sec. 8.2.4].
Part 2: Applying the interpolation inequality $\|\Psi\|_{H_{\diamond}^{3/2}}^{2}\leq\|\Psi\|_{H_{\diamond}^{1}}\|\Psi\|_{H_{\diamond}^{2}}$ in the last inequality and then Young’s inequality $ab\leq\frac{1}{\gamma}a^{2}+\frac{\gamma}{4}b^{2}$ yields
$\sqrt{2}|\mathbf{A}[\Psi,\Psi]|$
$\quad\geq\frac{\gamma}{2}\|\Psi\|^{2}_{H_{\diamond}^{2}}+\left(|\omega|\!-\!\frac{4}{\gamma^{3}}(C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}})^{4}\big{(}\|\Omega\!-\!\Omega_{\mathrm{{ref}}}\|^{2}_{H^{1}}\!+\!9\|\Omega\|^{2}_{H^{1}}\big{)}^{2}\right)\|\Psi\|^{2}_{H_{\diamond}^{1}}.$
If $|\omega|$ satisfies (23), then $\mathbf{A}$ is coercive, so $\mathbf{B}$ is boundedly invertible by the Lax-Milgram theorem [11, Section 6.2, Theorem 1]. ∎
We now indicate the dependence of $\mathbf{B}$ on the frequency $\omega$ explicitely by $\mathbf{B}_{\omega}$ and also consider complex values of $\omega$. This allows us to apply analytic Fredholm theory and study resonances:
Corollary 4.
There exists a discrete set $\Lambda\subset\mathbb{C}$ without accumulation points such that $\mathbf{B}_{\omega}$ is boundedly invertible in $L(H_{\diamond}^{2}({r{\mathbb{S}^{2}}}),H_{\diamond}^{-2}({r{\mathbb{S}^{2}}}))$ for $\omega\in\mathbb{C}\setminus\Lambda$, and $\dim\operatorname{ker}\mathbf{B}_{\omega}\in\mathbb{N}$ for $\omega\in\Lambda$.
The elements of $\operatorname{ker}\mathbf{B}_{\omega}$ for $\omega\in\Lambda$ are called inertial modes and have been studied intensively in [12] and other publications.
Proof of Corollary 4.
It follows from part 1 in the proof of Theorem 3 that $\mathbf{C}(\omega):H^{2}({r{\mathbb{S}^{2}}})\to H^{2}({r{\mathbb{S}^{2}}})$ defined via $\mathbf{C}(\omega):=(I-\Delta_{\mathrm{h}})^{-2}\mathbf{B}_{\omega}-I$ is a compact operator for all $\omega\in\mathbb{C}$. Moreover, $\omega\mapsto\mathbf{C}(\omega)$ is affinely linear and in particular holomorphic.
Therefore, we are in the framework of the analytic Fredholm theorem (see [33, Thm. 7.92]). Note that $\mathbf{B}_{\omega}$ is boundedly invertible if and only of if $I+\mathbf{C}(\omega)$ is boundedly invertible and that the kernels of both operators coincide. The first alternative of this theory, that is $(I+\mathbf{C}(\omega))^{-1}$ does not exist for any $\omega\in\mathbb{C}$, is excluded by the second part of Theorem 3. Therefore, the second alternative holds true and provides the statements of this corollary. ∎
3.3 Separated equations
Let us summarize some consequences for the separated operators $\mathbf{B}_{m}:V_{m}\to V_{m}^{*}=H^{-2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ defined by
$\left\langle\mathbf{B}_{m}\Psi_{m},\psi_{m}\right\rangle=\mathbf{A}_{m}[\Psi_{m},\psi_{m}],\qquad\Psi_{m},\psi_{m}\in V_{m}.$ (24)
Corollary 5.
For any parameters $\omega,\gamma,\Omega$ for which $\mathbf{B}$ is boundedly invertible, in particular for sufficiently large $|\omega|$ satisfying (23) as in Theorem 3, the separated operators $\mathbf{B}_{m}:V_{m}\to V_{m}^{*}$ are boundedly invertible for all $m$.
Proof.
It follows from (14) that $\|\mathbf{B}_{m}\|\leq\|\mathbf{B}\|$, so $\mathbf{B}_{m}$ is bounded. If $\mathbf{B}$ is boundedly invertible, then if follows from (14) and the inf-sup characterization of $\|\mathbf{B}_{m}^{-1}\|$ that $\|\mathbf{B}_{m}^{-1}\|\leq\|\mathbf{B}^{-1}\|$, and $\mathbf{B}_{m}$ is boundedly invertible as well. This completes the proof. ∎
Remark 6.
Repeating the argument of the proof of Theorem 3, it can also be shown that the separated operators $\mathbf{B}_{m}$ are Fredholm of index 0. Moreover, Corollary 4 holds true with $\mathbf{B}_{\omega}$ replaced by $\mathbf{B}_{m}=\mathbf{B}_{m,\omega}$.
Proposition 7 (Uniqueness for small $\Omega^{\prime}$).
For any $m\in\mathbb{Z}$ the operator $\mathbf{B}_{m}:V_{m}\to V_{m}^{*}$ is boundedly invertible if the derivative $\Omega^{\prime}$ (not necessarily differential rotation $\Omega$ itself) is small in comparison to the viscosity $\gamma$, in the sense that
$\|\Omega^{\prime}\|_{L^{2}({r{\mathbb{S}^{2}}})}\frac{|m|}{r}C_{H^{2}_{\diamond}\to L^{3}_{\diamond}}C_{H^{1}_{\diamond}\to L^{6}_{\diamond}}<\gamma.$ (25)
Proof.
Since $\mathbf{A}_{m}$ is bounded by Proposition 2, the Lax-Milgram lemma reduces the proof to verifying the coercivity of $\mathbf{A}$. To this end, we estimate its real part under the smallness condition (25). We will employ the (equivalent) norm $\|u\|_{H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)}:=\|\Delta u\|_{L^{2}({r{\mathbb{S}^{2}}})}$ for $u\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$. As $\Re\mathbf{A}^{\alpha}_{m}[\Psi,\Psi]=0$, one obtains
$\Re\mathbf{A}_{m}[\Psi,\Psi]$ $=\gamma\|\Delta_{m}\Psi\|^{2}_{L^{2}({r{\mathbb{S}^{2}}})}-\Re\big{(}im\big{\langle}\beta_{\Omega}\Delta_{m}\Psi,\Psi\big{\rangle}\big{)}.$
Using partial integration, the identity $\operatorname{\mathbf{grad}_{\mathit{m}\!}}\beta_{\Omega}=\frac{1}{r}\Omega^{\prime}\mathbf{e}_{\theta}$, Hölder’s inequality (21) and the embeddings (63), we can bound
$\Re\big{(}im\big{\langle}\beta_{\Omega}\Delta_{m}\Psi,\Psi\big{\rangle}\big{)}$
$\quad\leq|m|\left|\Im\big{\langle}\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega}\Psi)\big{\rangle}\right|=|m|\left|\Im\big{\langle}\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega})\Psi\big{\rangle}\right|$
$\quad\leq|m|\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi\|_{L^{6}}\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\beta_{\Omega}\|_{L^{2}}\|\Psi\|_{L^{3}}\leq\frac{|m|}{r}C_{H^{2}_{\diamond}\to L^{3}_{\diamond}}C_{H^{1}_{\diamond}\to L^{6}_{\diamond}}\|\Omega^{\prime}\|_{L^{2}({r{\mathbb{S}^{2}}})}\|\Psi\|_{H^{2}_{\diamond}}^{2}.$
Therefore, $\Re\mathbf{A}_{m}[\Psi,\Psi]\geq c\|\Delta_{m}\Psi\|^{2}_{L^{2}({r{\mathbb{S}^{2}}})}$ with $c:=\gamma-\frac{|m|}{r}C_{H^{2}_{\diamond}\to L^{3}_{\diamond}}C_{H^{1}_{\diamond}\to L^{6}_{\diamond}}\|\Omega^{\prime}\|_{L^{2}}>0$ ensured by the assumption (25). This completes the proof of coercivity. ∎
3.4 Additional regularity
Proposition 8 (Lifted regularity).
Let $m\in\mathbb{Z}$ and suppose that (23) or (25) hold true such that $\mathbf{B}_{m}:H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)\to H^{-2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ is boundedly invertible. Further assume that source and rotation have regularity
$f_{m}\in L^{2}({r{\mathbb{S}^{2}}};m),\qquad\Omega\in H^{2}({r{\mathbb{S}^{2}}};m=0).$ (26)
Then the solution $\Psi_{m}:=\mathbf{B}_{m}^{-1}f_{m}$ as well as $\tilde{\Psi}_{m}:=(\mathbf{B}_{m}^{\star})^{-1}f_{m}$ with the adjoint operator $\mathbf{B}_{m}^{\star}:H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)\to H^{-2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ have higher regularity
$\Psi_{m}\in H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m).$ (27)
Moreover, with $\Psi_{m}(\theta,\phi)=\hat{\Psi}_{m}(\theta)e_{m}(\phi)$, the boundary values $\Gamma_{m}\hat{\Psi}_{m}$ are well defined for all $m\neq 0$ and $\Gamma_{m}\hat{\Psi}_{m}=0,$ and similarly for $\tilde{\Psi}_{m}$.
Finally, the restrictions of $\mathbf{B}_{m}$ and $\mathbf{B}_{m}^{\star}$ have the forms
$\left.\begin{array}[]{l}\mathbf{B}_{m}=\gamma\Delta_{m}^{2}+i\omega\Delta_{m}-im\beta_{\Omega}\Delta_{m}+im\alpha_{\Omega}\\ \mathbf{B}_{m}^{\star}=\gamma\Delta_{m}^{2}-i\omega\Delta_{m}+im\Delta_{m}(\beta_{\Omega}\cdot)-im\alpha_{\Omega}\end{array}\right\}:H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)\to L^{2}({r{\mathbb{S}^{2}}};m)$
and they have bounded inverses $\mathbf{B}_{m}^{-1},(\mathbf{B}_{m}^{\star})^{-1}:L^{2}({r{\mathbb{S}^{2}}};m)\to H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)$.
Proof.
The solution $\Psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ satisfies
$\begin{aligned} \forall\psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m):\quad&\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle=\left\langle\tilde{f}_{m},\psi_{m}\right\rangle\\ &\mbox{with }\tilde{f}_{m}:=f_{m}+i\omega\Delta_{m}\Psi_{m}-im\beta_{\Omega}\Delta_{m}\Psi_{m}+im\alpha_{\Omega}\Psi_{m}.\end{aligned}$ (28)
We shall claim that $\tilde{f}_{m}\in L^{2}({r{\mathbb{S}^{2}}};m)$. As $\alpha_{\Omega}=\Omega^{\prime\prime}\sin+3\Omega^{\prime}\cot-2\Omega\sin$, hence $\Omega\in H^{2}({r{\mathbb{S}^{2}}};0)$ implies
$\|\alpha_{\Omega}\Psi_{m}\|_{L^{2}}$ $\leq\|(\Omega^{\prime\prime}-2\Omega\sin)\Psi_{m}\|_{L^{2}}+\left\|3\Omega^{\prime}\cos\frac{\Psi_{m}}{\sin}\right\|_{L^{2}}$
$\leq(\|\Omega^{\prime\prime}\|_{L^{2}}+2\|\Omega\|_{L^{2}})\|\Psi_{m}\|_{L^{\infty}}+3\|\Omega^{\prime}\|_{L^{4}}\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi_{m}\|_{L^{4}}$
with noting that $\Psi_{m}/\sin$ is $\phi$-component of $\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi_{m}$ as defined in (59) for $m\neq 0$; for the case $m=0$, the term $\alpha_{\Omega}$ does not appear in the equation. Using the embeddings (63) and setting $C_{\alpha}:=2C_{H^{2}\to L^{\infty}}+3C^{2}_{H^{1}\to L^{4}}$, we obtain
$\|\alpha_{\Omega}\Psi_{m}\|_{L^{2}}\leq C_{\alpha}\|\Omega\|_{H^{2}}\|\Psi_{m}\|_{H^{2}}.$ (29)
With this we can derive the following norm bound:
$\|\tilde{f}_{m}\|_{L^{2}}$ $\leq\|f_{m}\|_{L^{2}}+|\omega|\|\Psi_{m}\|_{H^{2}}+|m|\|\beta_{\Omega}\|_{L^{\infty}}\|\Psi_{m}\|_{H^{2}}+|m|C_{\alpha}\|\Omega\|_{H^{2}}\|\Psi_{m}\|_{H^{2}}$
$\leq\|f_{m}\|_{L^{2}}+\big{(}|\omega|+|m|(C_{H^{2}\to L^{\infty}}\|\Omega\|_{H^{2}}+\|\Omega_{\mathrm{{ref}}}\|_{L^{\infty}}+C_{\alpha}\|\Omega\|_{H^{2}})\big{)}\|\Psi_{m}\|_{H^{2}}.$
It follows from the isometry property (62) that weak solutions $\Psi\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ to the biharmonic equation $\Delta_{\mathrm{h}}^{2}\Psi=f$ with $f\in L^{2}({r{\mathbb{S}^{2}}})$ belong to $H^{4}({r{\mathbb{S}^{2}}})$. Due to the separability of $\Delta_{\mathrm{h}}$, weak solutions $\Psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ to the separated equation $\Delta_{m}^{2}\Psi_{m}=\tilde{f}_{m}$ with $\tilde{f}_{m}\in L^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ also belong to $H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)$ with $\|\Psi_{m}\|_{H^{4}}=\|\tilde{f}_{m}\|_{L^{2}}$. Together with $\|\Psi_{m}\|_{H^{2}}\leq\|\mathbf{B}_{m}^{-1}\|\|f_{m}\|_{H^{-2}}\leq\|\mathbf{B}_{m}^{-1}\|\|f_{m}\|_{L^{2}}$ this shows that $\|\Psi_{m}\|_{H^{4}}\leq C\|f\|_{L^{2}}$, i.e., $\mathbf{B}_{m}:H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)\to L^{2}({r{\mathbb{S}^{2}}};m)$ is boundedly invertible. The proof of the analogous statement for $\mathbf{B}_{m}^{\star}$ only requires a different treatment of the term involving $\beta_{\Omega}$. Here, we use the fact that $H^{2}({r{\mathbb{S}^{2}}})$ is a Banach algebra (see [2, p. 115] or [3, Thm 1.4]) to show that $\|\beta_{\Omega}\Psi_{m}\|_{H^{2}}\leq C\|\beta_{\Omega}\|_{H^{2}}\|\Psi_{m}\|_{H^{2}}$.
Concerning the boundary values, we use the continuous embedding $H^{4}({r{\mathbb{S}^{2}}})\hookrightarrow C^{2}({r{\mathbb{S}^{2}}})$ in (55c) and Lemma 18 to show that $\Gamma_{m}\hat{\Psi}_{m}$ is well-defined and vanishes. ∎
3.5 Well-posedness on the restricted domain
Inspired by the practical measurement scenario in Section 2.3, where only data on the restricted latitude range $I^{\prime}=(\epsilon,\pi-\epsilon)$ can be observed, this section derives well-posedness, regularity of the wave operators on this restricted domain. This result will be later employed in Section 5.2 for convergence guarantee of reconstruction method under restricted measurements.
We sketch how an analysis similar to that of Propositions 7 and 8 can be carried out, whose key elements are symmetry of the bi-Laplacian and Laplacian operators and partial integrations. According to Remark 20, symmetry of Laplacian operators on function spaces supported on $I^{\prime}$ is guaranteed with homogeneous Cauchy boundary data, i.e. $\Gamma_{2}$ in (10).
To this end, we define the following function spaces on $I^{\prime}$ with norms inherited from $H^{s}({r{\mathbb{S}^{2}}};m)$:
$\begin{split}&H^{s}(I^{\prime};m):=\left\{\Psi|_{I^{\prime}}:\Psi\in H^{s}({r{\mathbb{S}^{2}}};m)\right\},s\geq 0,\\ &H^{s}_{0}(I^{\prime};m):=\left\{\Psi|_{I^{\prime}}:\Psi\in H^{s}({r{\mathbb{S}^{2}}};m),\Psi|_{\partial I^{\prime}}=\Psi^{\prime}|_{\partial I^{\prime}}=0\right\},s\geq 2.\end{split}$ (30)
Note that these spaces coincide with the standard Sobolev spaces $H^{s}(I^{\prime})$ and $H^{s}_{0}(I^{\prime})$ with equivalent norms. Correspondingly, the dual spaces are, respectively, given by $H^{-s}_{0}(I^{\prime};m):=H^{s}(I^{\prime};m)^{\prime}=H^{-s}_{0}(I^{\prime})$ and $H^{-s}(I^{\prime};m):=H^{s}_{0}(I^{\prime};m)^{\prime}=H^{-s}(I^{\prime})$, again with equivalent norms.
Proposition 9 (Well-posedness on $I^{\prime}$).
Given $\gamma>0$, and $\Omega\in H^{1}({r{\mathbb{S}^{2}}};m=0)$ satisfying (23) or (25). Assume for any $m\in\mathbb{Z}$ the Cauchy boundary condition
$\Psi|_{\partial I^{\prime}}=u_{1}\in\mathbb{C}^{2}\quad\text{and}\quad\Psi^{\prime}|_{\partial I^{\prime}}=u_{2}\in\mathbb{C}^{2}.$ (31)
Let $u\in C^{\infty}(I)$ with $u|_{\partial I^{\prime}}=u_{1}$, $u^{\prime}|_{\partial I^{\prime}}=u_{2}$ and $g:=\mathbf{B}_{m}u$. We claim:
  1. (i)
    The operator $\mathbf{B}_{m}:H_{0}^{2}(I^{\prime};m)\to H^{-2}(I^{\prime};m)$ is boundedly invertible. The wave solution is $\Psi=\mathbf{B}_{m}^{-1}(f_{m}-g)+u\in H^{2}(I^{\prime};m)$.
  2. (ii)
    Furthermore, if $\Omega\in H^{2}({r{\mathbb{S}^{2}}};m=0)$ as in Proposition 8, then the linear bounded operator $\mathbf{B}_{m}$, $\mathbf{B}_{m}^{\star}$ have bounded inverses $\mathbf{B}_{m}^{-1},(\mathbf{B}_{m}^{\star})^{-1}:L^{2}(I^{\prime};m)\to H^{4}_{0}(I^{\prime};m)$.
Proof.
The splitting $\Psi=\mathbf{B}_{m}^{-1}(f_{m}-g)+u$ allows us to study the differential operators $\mathbf{B}_{m}$ on the state space $H^{2}_{0}(I^{\prime};m)$ with homogeneous boundaries as defined in (30). Symmetry of the bi-Laplacian and Laplacian operators with homogeneous Cauchy boundary condition (31) shown in Remark 13 enables the sequilinear form $\mathbf{A}$ in (13). Next, the integration-by-part (3.1) in Lemma 1 holds under Cauchy boundary condition, implying boundedness of $\mathbf{A}$, $\mathbf{A}^{\alpha}$ as in Proposition 2, and well-posedness of the unseparated equation in the same setting as Theorem 3. Then, for the separated equations, following the proof of Proposition 7, in which the partial integration also holds on the subdomain $I^{\prime}$, we obtain the same uniqueness result.
Concerning the lifted regularity result as in Proposition 8, by setting the auxiliary source $\tilde{f}_{m}$ as in (28) and partial integration, we (formally) have $\gamma\left\langle\Delta_{m}^{2}\Psi_{m},\psi_{m}\right\rangle=\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle=\left\langle\tilde{f}_{m},\psi_{m}\right\rangle$ for all $\psi_{m}\in H^{2}_{0}(I^{\prime};m)$. Using the fact the weak solutions of the bi-Laplace equations are also strong solutions, one obtains boundedness of the fourth order term $\|\Delta_{m}^{2}\Psi_{m}\|_{L^{2}}$, given higher regularity of $\Omega$ and $f$ as in Proposition 8. Alternatively, one can apply general elliptic regularity results as, e.g., in [38, Sec. 5.3.4]. ∎
4 Inverse problems for viscosity and rotation
4.1 Inversion with different observation strategies
In this section, we address the inverse problem of recovering the differential rotation function $\Omega$ and the viscosity coefficient $\gamma$ in the separated equations. To achieve this, we require additional information, namely full or partial measurements of the state, as described in Section 2.3:
$\begin{array}[]{l}L:V_{m}\to Y\\ (L\Psi)(\theta):=\Psi(\theta)\end{array}\qquad Y=\begin{cases}L^{2}(I,r^{2}\sin),&\text{full measurements}\\ L^{2}(I^{\prime},r^{2}\sin),&\text{partial measurements}\end{cases}$ (34)
with the weighted $L^{2}$-data spaces. We formulate the parameter identification problem within the Hilbert space framework established by the well-posedness results in Sections 3.2-3.3:
$\gamma\in\mathbb{R}^{+},\quad\Omega\in X:=H^{1}({r{\mathbb{S}^{2}}},m=0),\quad\Psi\in V_{m}:=H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m),\quad y\in Y,$ (35)
By introducing the parameter-to-state map
[…]
together with the measurement operator we can formulate the forward operator $F$ for the parameter identification as
$F:\mathcal{D}(S)\to Y\qquad F:=L\circ S.$ (37)
Note that $X$ is the space of real-valued functions, while $V_{m}$ and $Y$ are spaces of complex-valued functions; together, they form Gelfand triples.
4.2 Sensitivity and adjoints
Compactness of the embedding $V_{m}\hookrightarrow L^{2}(I,r^{2}\sin)$ yields compactness of the measurement operator $L$, thus of the forward operator $F$. In such cases, the inverse coefficient problem becomes ill-posed, necessitating regularization for stable reconstruction. Given the inherent nonlinearity of the problem, iterative methods such as Landweber-type and Newton-type algorithms are typically employed for reconstruction (see Section 5). For noisy data $y^{\delta}\in Y$, the former is given by
$p_{k+1}=p_{k}-F^{\prime}(p_{k})^{*}(F(p_{k})-y^{\delta})\qquad k\leq K(y^{\delta},\delta),$
where the unknown parameter is $p:=(\gamma,\Omega)$. The stopping index $K(y^{\delta},\delta)$ serves as the regularization parameter and is determined according to the discrepancy principle; see Section 6 for further details. This formulation highlights that the main components of iterative gradient-based methods: sensitivity analysis and the derivation of the adjoint; these are the primary focus of this section.
To this end, we will add the dependence on the unknown parameters $p=(\gamma,\Omega)\in\mathcal{D}(S)$ in parenthesis, i.e. $\mathbf{B}_{m}(p)=\mathbf{B}_{m}:V_{m}\to V_{m}^{*}.$ Since the mapping $p\mapsto\mathbf{B}_{m}(p)$ is affine linear, its derivative with respect to $p$ is independent of $p$ and we can omit it in our notation. Hence, we simply denote the derivative of $p\mapsto\mathbf{B}_{m}(p)$ by
$\mathbf{B}_{m}^{\prime}:\mathbb{R}\times X\to\mathcal{L}(V_{m},V_{m}^{*}),\qquad\mathbf{B}_{m}^{\prime}(\delta\gamma,\delta\Omega)\Psi:=(\delta\gamma)\Delta_{m}^{2}\Psi+im\alpha_{\delta\Omega}\Psi-im(\delta\Omega)\Psi.$
More precisely, $\mathbf{B}_{m}^{\prime}(\delta\gamma,\delta\Omega)$ is the operator induced by the sesquilinear form $\mathbf{A}^{\prime}_{m}(\delta\gamma,\delta\Omega):V_{m}\times V_{m}\to\mathbb{C}$ defined by
$\mathbf{A}^{\prime}_{m}(\delta\gamma,\delta\Omega)[\Psi,\psi]:=(\delta\gamma)\left\langle\Delta_{m}\Psi,\Delta_{m}\psi\right\rangle+im(\alpha_{\delta\Omega}-\delta\Omega)\langle\Psi,\psi\rangle.$
We wish to emphasize that a similar analysis can be carried out for the unseparated equation with obvious modifications.
Lemma 10 (Sensitivity).
The parameter-to-state map $S$ in (36) is Fréchet differentiable, and for any $p\in\mathcal{D}(S)$ and $\delta p\in\mathbb{R}\times X$ we have
$S^{\prime}[p]\delta p=-\mathbf{B}_{m}(p)^{-1}\mathbf{B}_{m}^{\prime}(\delta p)\Psi\in V_{m},$ (38)
where $\Psi:=\mathbf{B}_{m}(p)^{-1}f$ is the weak solution to the primal equation (9).
Proof.
The result follows from the differentiability statement in the implicit function theorem (see, e.g., [41, §4.7]) applied to the operator $G:\mathcal{D}(S)\times V_{m}\to V_{m}^{*}$, $G(p,\Psi):=\mathbf{B}_{m}(p)\Psi-f$. Here, one interprets $S$ as the implicitly defined function $G(p,S(p))=0$ and employs linearity of $G$ in the second argument as well as bounded invertibility of $\mathbf{B}_{m}(p):V_{m}\to V_{m}^{*}$. ∎
Differentiability of $F$ is straightforward from that of $S$ and boundedness of the linear measurement operator $L$ defined in (34). We now derive the adjoint for different observation senarios.
Proposition 11 (Adjoint – full observation).
Let $L$ be the full measurement operator in (34). Denote by $I_{X}:X^{*}\to X$ the Riesz isomorphism between dual spaces.
The Hilbert space adjoint of the derivative of $F$ in (37) is given by
$F^{\prime}[p]^{*}:Y\to\mathbb{R}\times X\qquad F^{\prime}[p]^{*}y=-\left[\begin{smallmatrix}1&0\\ 0&I_{X}\end{smallmatrix}\right]\,\Re\left([\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}(\mathbf{B}_{m}(p)^{-1})^{\star}y\right)$ (39)
where $\Psi:=\mathbf{B}_{m}(p)^{-1}f$ is again the solution to the separated equation (9) and
$[\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}:V_{m}\to\mathbb{R}\times X^{*},\quad(\mathbf{B}_{m}(p)^{-1})^{\star}:V_{m}^{*}\to V_{m}$ (40)
denotes the Banach space adjoints of $\mathbf{B}_{m}^{\prime}(\cdot)\Psi:\mathbb{R}\times X\to V_{m}^{*}$ and $\mathbf{B}_{m}(p)^{-1}:V_{m}^{*}\to V_{m}$.
Proof.
Since the parameter space is real, we impose on $Y$ the real-valued inner product, the real part of the canonical complex-valued inner product, as
$(f,g)_{Y}:=(\Re\left(f\right),\Re\left(g\right))_{Y}+(\Im\left({f}\right),\Im\left({g}\right))_{Y}=\Re\left((f,g)_{Y}^{\mathbb{C}}\right).$
The linearization (39) and Banach space adjoints (40) enable us to deduce
$\Big{(}F^{\prime}[p]\delta p\,,\,y\Big{)}_{Y}^{\mathbb{C}}=\left\langle S^{\prime}[p]\delta p\,,y\right\rangle_{L^{2}}^{\mathbb{C}}=-\left\langle\mathbf{B}_{m}(p)^{-1}\mathbf{B}_{m}^{\prime}(\delta p){\Psi}\,,y\right\rangle_{L^{2}}^{\mathbb{C}}$
$=-\left\langle\mathbf{B}_{m}^{\prime}(\delta p)\Psi\,,\,(\mathbf{B}_{m}(p)^{-1})^{\star}y\right\rangle_{V_{m}^{*},V_{m}}^{\mathbb{C}}=-\left\langle\delta p\,,\,[\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}(\mathbf{B}_{m}(p)^{-1})^{\star}y\right\rangle_{\mathbb{R}\times X,\mathbb{R}\times X^{*}}^{\mathbb{C}}.$
Then with the real-valued inner product and the Riesz isomorphism $I_{X}$, we arrive at
$\Big{(}F^{\prime}[p]\delta p\,,\,y\Big{)}_{Y}$ $=\left\langle\delta p\,,\,-\Re\left([\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}(\mathbf{B}_{m}(p)^{-1})^{\star}y\right)\right\rangle_{\mathbb{R}\times X,\mathbb{R}\times X^{*}}$
$=\Big{(}\delta p\,,\,-\left[\begin{smallmatrix}1&0\\ 0&I_{X}\end{smallmatrix}\right]\Re\left([\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}(\mathbf{B}_{m}(p)^{-1})^{\star}y\right)\Big{)}_{\mathbb{R}\times X},$
in agreement with the expression (39). ∎
We next derive the explicit expression for the Hilbert space adjoint (39).
Corollary 12 (Explicit form).
The adjoint for full observation in Proposition 11 takes the explicit form
$\begin{split}&F^{\prime}[\gamma,\Omega]^{*}y=\begin{bmatrix}\Re\left(\int_{0}^{\pi}(\Delta_{m}^{2}\Psi)\overline{z}r^{2}\sin(\cdot)\,ds\right)\\[3.87498pt] mI_{X}\Im\left({\dfrac{\sin(\cdot)}{r^{2}}\dfrac{d}{d\theta}\left(\dfrac{1}{\sin(\cdot)}\dfrac{d(\overline{\Psi}z)}{d\theta}\right)-\left(\Delta_{m}\overline{\Psi}\right)z}\right)\end{bmatrix}\end{split}$ (41)
with the adjoint state $z:=(\mathbf{B}_{m}(p)^{-1})^{\star}y$.
Proof.
To compute $\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}$, take $(\delta\gamma,\delta\Omega)\in\mathbb{R}\times X$ and $z\in V_{m}$ and note that
$\left\langle\mathbf{B}_{m}^{\prime}(\delta\gamma,\delta\Omega){\Psi},z\right\rangle$ $=\left\langle\delta\gamma\Delta_{m}^{2}\Psi-im(\delta\Omega)\Delta_{m}\Psi+im\alpha_{\delta\Omega}\Psi\,,\,z\right\rangle$
$=\delta\gamma\int_{0}^{\pi}(\Delta_{m}^{2}\Psi)\overline{z}r^{2}\sin(\cdot)\,d\theta+\left\langle\delta\Omega\,,\,im(\Delta_{m}\overline{\Psi})z\right\rangle-\left\langle\alpha_{\delta\Omega}\,,\,im\overline{\Psi}z)\right\rangle$
To characterize the quantity involving $\alpha_{\delta\Omega}$, we may assume w.l.o.g. that $m\neq 0$ since $\alpha$ is irrelevant for $m=0$. Using the fact the $\Psi$ and $z$ satisfy the boundary conditions $\Gamma_{m}(\Psi)=\Gamma_{m}(z)=0$ under our regularity assumption on $\Omega$ according to Proposition 8, we can carry out the following partial integrations:
$\left\langle\alpha_{\delta\Omega}\,,\,\overline{\Psi}z\right\rangle$
$\quad=\int_{0}^{\pi}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\right)\Psi\overline{z}\,d\theta=\int_{0}^{\pi}\frac{-1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\frac{d}{d\theta}\left(\Psi\overline{z}\right)\,d\theta$
$\quad=\int_{0}^{\pi}(\delta\Omega)(\theta)\sin^{2}\theta\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\Psi\overline{z}\right)\right)\,d\theta=\left\langle\delta\Omega\,,\,\frac{\sin}{r^{2}}\frac{d}{d\theta}\left(\frac{1}{\sin}\frac{d}{d\theta}\left(\overline{\Psi}z\right)\right)\right\rangle$
Inserting these formulas into (39) yields (41). ∎
Remark 13 (Adjoint for partial and real measurements).
Note that the partial observation operator is the restriction operator $R\Psi:=\Psi|_{I^{\prime}}$. Its Hilbert space adjoint is the extension-by-zero operator
$R^{*}:L^{2}(I^{\prime},r^{2}\sin)\to L^{2}(I,r^{2}\sin)\qquad R^{*}\Psi)(\theta):=\begin{cases}\Psi(\theta),&\theta\in I^{\prime},\\ 0,&\theta\in I\setminus I^{\prime}.\end{cases}$
The corresponding forward operator relates to the forward operator with full measurement via $F_{\mathrm{part}}[p]=RF_{\mathrm{full}}[p]$. Thus, $F_{\mathrm{part}}^{\prime}[p]^{*}=F_{\mathrm{full}}^{\prime}[p]^{*}R^{*}$ with $F_{\mathrm{full}}^{\prime}[p]^{*}$ as in Corollary 12.
Similarly, if only the real part of the data can be observed, analogous formulas hold true with $R$ replaced by the real-part operator $\Re:L^{2}_{\mathbb{C}}(I,r^{2}\sin)\to L^{2}_{\mathbb{R}}(I,r^{2}\sin)$, of which the Hilbert space adjoint is the canonical embedding of $L^{2}_{\mathbb{R}}(I,r^{2}\sin)$ into $L^{2}_{\mathbb{C}}(I,r^{2}\sin)$.
5 Tangential cone condition
As discussed in Section 4, the inverse problem under consideration is ill-posed, necessitating the use of regularization to achieve stable reconstruction of the target coefficients. Iterative gradient-based regularization methods, such as Landweber iteration, Newton-type schemes, and their variants, fundamentally rely on three core elements: the well-posedness of the forward operator $F$ (Section 3), adjoint of the linearized operator (Section 4.2), and guarantees of convergence. This final requirement is typically ensured by making specific structural assumptions about $F$, particularly regarding its nonlinearity and the uniform boundedness of its derivative; these issues are the main focus of this section.
To address this, we consider the tangential cone condition (TCC), a celebrated criterion first introduced in [16] that ensures local convergence of iterative regularization algorithms. Intuitively, if the forward map $F$ is excessively nonlinear, there is no general guarantee that gradient descent steps will remain within the vicinity of the true solution. The tangential cone condition offers a quantitative tool for assessing the nonlinearity of $F$, especially in the context of compact operators, thereby paving the way for the successful application of iterative regularization methods. Notably, satisfaction of this condition also implies a characterization of the solution set via null space of the linearized operator, uniqueness of the minimum-norm solution, and, combined with the null space condition, ensures that the iterative reconstruction sequence converges to this solution [22].
For a general model $F:X\ni p\mapsto y\in Y$, TCC is stated in the following (strong) form, that is the inequality
$\|F(p)-F(\tilde{p})-F^{\prime}(p)(p-\tilde{p})\|_{Y}\leq C_{tc}\|p-\tilde{p}\|_{X}\|F(p)-F(\tilde{p})\|_{Y}\,\,\,\text{for all }p,\tilde{p}\in B^{X}_{R}(p^{\dagger})$
holds for some constant $C_{tc}>0$ and radius $R>0$ of the ball around a ground truth $p^{\dagger}\in\mathcal{D}(F)$; see also [25] for the weak forms. In our setting, TCC reads as
$\|F(\gamma,\Omega)-F(\tilde{\gamma},\tilde{\Omega})-F^{\prime}[\gamma,\Omega]((\gamma,\Omega)-(\tilde{\gamma},\tilde{\Omega}))\|_{Y}$ (42)
$\hskip 113.81102pt\leq C_{tc}\|(\gamma,\Omega)-\tilde{\gamma},\tilde{\Omega})\|_{\mathbb{R}\times X}\|F(\gamma,\Omega)-F(\tilde{\gamma},\tilde{\Omega})\|_{Y}$
$\hskip 227.62204pt\text{for all }(\gamma,\Omega),(\tilde{\gamma},\tilde{\Omega})\in B_{R}^{\mathbb{R}\times X}(\gamma^{\dagger},\beta^{\dagger}).$
It is worth noting that, in addition to the TCC, several other conditions can be employed to characterize the structure of the forward map and provide convergence guarantees. Notable examples include range-invariance [24], convexity, and the Polyak-Lojasiewicz condition [31].
Although the tangential cone condition is a powerful tool, it can be challenging to verify in practice; for a general verification strategy with full data, we refer the reader to [23]. Contributions in this area span a wide range of applications, including elliptic inverse problems [18], elastography [21, 20], electrical impedance tomography [26], full waveform inversion [10], and more recently, neural network-based inverse problems [35, 1].
In this work, we contribute to the literature by establishing the tangential cone condition for full measurements, introducing a lifted regularity strategy developed in Section 3.4 to account for realistic measured data in $L^{2}$. Furthermore, we extend our verification to the setting of restricted measurements.
5.1 Full measurement
Lemma 14 (Uniform boundedness).
Given the setting in Proposition 8. The operators $\mathbf{B}_{m}(\gamma^{\dagger},\Omega^{\dagger})^{-1}$ are bounded from $H^{-4}_{\diamond}({r{\mathbb{S}^{2}}};m)$ to $L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ for all $(\gamma^{\dagger},\Omega^{\dagger})\in\mathcal{D}(S)\cap(\mathbb{R}^{+}\times H^{2}({r{\mathbb{S}^{2}}};m=0))$. Moreover, they are locally uniformly bounded in the sense that there exists $\overline{R}(\gamma^{\dagger},\Omega^{\dagger})>0$ such that
$N^{\gamma^{\dagger},\Omega^{\dagger}}(R):=\sup\left\{\|\mathbf{B}_{m}(\gamma,\Omega)^{-1}\|_{H^{-4}\to L^{2}}:(\gamma,\Omega)\in B_{R}^{\mathbb{R}\times X}(\gamma^{\dagger},\Omega^{\dagger})\right\}$
is finite for all $R\leq\overline{R}(\gamma^{\dagger},\Omega^{\dagger})$, and $B_{R}^{\mathbb{R}\times X}(\gamma^{\dagger},\Omega^{\dagger})\subset\mathcal{D}(S)$.
Proof.
Boundedness follows from the lifted regularity result in Proposition 8 and that $\mathbf{B}_{m}(\gamma^{\dagger},\Omega^{\dagger}):L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)\to H^{-4}_{\diamond}({r{\mathbb{S}^{2}}};m)$ is the Banach adjoint of $\mathbf{B}_{m}(\gamma^{\dagger},\Omega^{\dagger})^{\star}:H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)\to L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$. The local uniformity of norm bounds can be seen from the fact that the norms of $\gamma$, $\Omega$ appear in an affine liner manner in all upper bounds in the proof of Proposition 8. ∎
Theorem 15 (TCC – full measurement).
Given the setting in Proposition 8. For $m\neq 0$, the forward operator $F$ with full measurements satisfies the TCC (42).
Proof.
Recall $F=S$ for the full measurements . For any $p^{\dagger}=(\gamma^{\dagger},\Omega^{\dagger})\in\mathcal{D}(S)$, we have to show that there exists $R,C_{\text{tc}}>0$ such that
[…]
with $h:=\tilde{p}-p$, for all $p,\tilde{p}\in B_{R}^{\mathbb{R}\times X}(\gamma^{\dagger},\beta^{\dagger})$. Using the identity $\mathbf{B}_{m}(p+h)-\mathbf{B}_{m}(p)=\mathbf{B}_{m}^{\prime}(h)$ and twice the identity $T^{-1}-S^{-1}=T^{-1}(S-T)S^{-1}$ for invertible operators $S$ and $T$, we obtain
$\begin{split}S(p+h)-S(p)-S^{\prime}[p]h&=\mathbf{B}_{m}(p+h)^{-1}f-\mathbf{B}_{m}(p)^{-1}f+\mathbf{B}_{m}(p)^{-1}\mathbf{B}_{m}^{\prime}(h)S(p)\\ &=[-\mathbf{B}_{m}(p+h)^{-1}+\mathbf{B}_{m}(p)]^{-1}]\mathbf{B}_{m}^{\prime}(h)S(p)\\ &=\mathbf{B}_{m}(p)^{-1}\mathbf{B}_{m}^{\prime}(h)\mathbf{B}_{m}(p+h)^{-1}\mathbf{B}_{m}^{\prime}(h)\mathbf{B}_{m}(p)^{-1}f\\ &=-\mathbf{B}_{m}(p)^{-1}\mathbf{B}_{m}^{\prime}(h)(S(p+h)-S(p)).\end{split}$
Therefore, it suffices to show that $\|\mathbf{B}_{m}(p)^{-1}\mathbf{B}_{m}^{\prime}(h)\|_{L^{2}\to L^{2}}\leq C_{\mathrm{tc}}\|h\|_{\mathbb{R}\times H^{2}}$.
We first estimate $\|\mathbf{B}_{m}^{\prime}(h)\|_{L^{2}\to H^{-4}}$. Let $h=:(\delta\gamma,\delta\Omega)$, for $w\in L^{2}({r{\mathbb{S}^{2}}};m)$ and $\varphi\in H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)$, we obtain the following bound on the inner product:
$\left\langle\mathbf{B}_{m}^{\prime}(h)w,\varphi\right\rangle=\left\langle\delta\gamma w,\Delta_{m}^{2}\varphi\right\rangle-\left\langle im(\delta\Omega)w,\Delta_{m}\varphi\right\rangle+\left\langle imw,\alpha_{\delta\Omega}\varphi\right\rangle$
$\leq$ $|\delta\gamma|\|w\|_{L^{2}}\|\varphi\|_{H^{4}}+|m|\|\delta\Omega\|_{L^{\infty}}\|w\|_{L^{2}}\|\varphi\|_{H^{2}}+C_{\alpha}|m|\|\delta\Omega\|_{H^{2}}\|w\|_{L^{2}}\|\varphi\|_{H^{2}}$
$\leq$ $C_{\text{emb}}\|h\|_{\mathbb{R}\times\widetilde{X}}\|w\|_{L^{2}}\|\varphi\|_{H^{4}}$
where $C_{\mathrm{emb}}:=\sqrt{2}\max\big{\{}1,|m|C_{H^{4}\to H^{2}}(C_{H^{2}\to L^{\infty}}+C_{\alpha})\big{\}}$ with $C_{\alpha}$ as in (29). This shows that
$\|\mathbf{B}_{m}^{\prime}(h)\|_{L^{2}\to H^{-4}}\leq C_{\text{emb}}\|h\|_{\mathbb{R}\times\widetilde{X}}$.
Then, together with the locally uniform boundedness in Lemma 14, we achieve
$\begin{split}\|\mathbf{B}_{m}(p)^{-1}\mathbf{B}_{m}^{\prime}(h)\|_{L^{2}\to L^{2}}&\leq\|\mathbf{B}_{m}(p)^{-1}\|_{H^{-4}\to L^{2}}\|\mathbf{B}_{m}^{\prime}(h)\|_{L^{2}\to H^{-4}}\\ &\leq N^{\gamma^{\dagger},\Omega^{\dagger}}(R)C_{\text{emb}}\|h\|_{\mathbb{R}\times H^{2}}.\end{split}$ (43)
In this step, we observe the importance of the lifted regularity laid out in Section 3.4, which is the key to Lemma 14. This proves the TCC (42) in $B^{X}_{R}(p^{\dagger})$. ∎
With TCC, we further obtain unique identifiability of $\Omega$ given $\gamma$ and vice versa. To this end, we denote by $\mathrm{Arg}(z)\in[0,2\pi)$ the principal value of the argument of a complex number $z$.
Corollary 16 (Unique identifiability).
Let $m\neq 0$ and assume full measurement.
  1. (i)
    Assume $\gamma$ is known. Suppose that
    $\Delta_{m}\Psi\neq 0\text{ a.e.~on }I,\quad\mathrm{Arg}(\Psi)-\mathrm{Arg}(\Delta_{m}\Psi)\notin\{-\pi,0,\pi\}\text{ a.e.~on }I.$ (44)
    Then $\Omega$ is uniquely determined locally in $B_{R}^{X}(\Omega^{\dagger})$.
  2. (ii)
    i) Assume $\Omega$ is known. Suppose that
    $\Delta_{m}^{2}\Psi\neq 0\text{ on some non-null subset of }I.$ (45)
    Then $\gamma$ is uniquely determined by full measurement, locally in $B_{R}^{X}(\Omega^{\dagger})$.
Proof.
(i) We first prove $\mathcal{N}(F^{\prime}(\Omega^{\dagger}))=\{0\}$, where $F=L\circ S$ with the embedding $L$. By injectivity of the embedding, it suffices to show $\mathcal{N}(S^{\prime}(\Omega^{\dagger}))=\{0\}$. Recall from Lemma 10 the linearized state equation
$S^{\prime}[\Omega^{\dagger}]\delta\Omega=-\mathbf{B}_{m}(\Omega^{\dagger})^{-1}(-im(\delta\Omega)\Delta_{m}\Psi^{\dagger}+im\alpha_{\delta\Omega}\Psi^{\dagger}),$
where $\Psi^{\dagger}=S(\Omega^{\dagger})$. Due to invertibility of $\mathbf{B}_{m}(\Omega^{\dagger})$, the linearized state $S^{\prime}[\Omega^{\dagger}]\delta\Omega=0\in L^{2}(\Omega)$ if and only if $\alpha_{\delta\Omega}\Psi^{\dagger}=(\delta\Omega)\Delta_{m}\Psi^{\dagger}$ a.e. on $I$.
This is clearly true if $\alpha_{\delta\Omega}\Psi^{\dagger}=(\delta\Omega)\Delta_{m}\Psi^{\dagger}=0$ a.e. on $I$, which would then imply $\delta\Omega=0$ a.e on $I$ by the first assumption in (44).
If this is not the case, there exists a non-null subset of $I$ where $\alpha_{\delta\Omega}\Psi^{\dagger}=(\delta\Omega)\Delta_{m}\Psi^{\dagger}\neq 0$ a.e. This allows taking $\mathrm{Arg}$ here; noting that $\delta\Omega$, $\alpha_{\delta\Omega}$ are real functions, it necessarily holds that $\mathrm{Arg}(\Psi)-\mathrm{Arg}(\Delta_{m}\Psi)\in\{-\pi,0,\pi\}$ a.e. on this non-null subset, contradicting the second part of assumption (44). Thus, $\mathcal{N}(F^{\prime}(\Omega^{\dagger}))=\{0\}$.
From this, along with the characterization of the solution set given the TCC in [16, Proposition 2.1], we deduce $\Omega^{*}-\Omega^{\dagger}\in\mathcal{N}(F^{\prime}(\Omega^{\dagger}))=\{0\}$ for any solution $\Omega^{*}$, $\Omega^{\dagger}$ of the inverse problem. We thus claim the unique identifiability of $\Omega$ given full measurement.
(ii) Assumption (45) implies that $F^{\prime}[\gamma^{\dagger}]\delta\gamma=\delta\gamma\Delta_{m}^{2}\Psi^{\dagger}=0\in L^{2}(\Omega)$ holds if and only if $\delta\gamma=0\in\mathbb{R}$. The remainder of the argument proceeds as above. ∎
5.2 Partial measurement including Cauchy data
We now turn to the case of restricted measurements, where observations are limited to $I^{\prime}=(\epsilon,\pi-\epsilon)$ as described in Section 2.3. As a simplification, we additionally assume that Cauchy data at the boundary of $I^{\prime}$ can be measured exactly.
We follow the analysis for the full Sun, while specifically leveraging the well-posedness and regularity results for the restricted domain established in Section 3.5, Proposition 9.
Theorem 17 (TCC – partial measurement).
Consider the setting in Proposition 9, (ii). If one imposes the constraint
$\Psi|_{\partial I^{\prime}}=\Psi^{\dagger}|_{\partial I^{\prime}}\quad\text{and}\quad\Psi^{\prime}|_{\partial I^{\prime}}=(\Psi^{\dagger})^{\prime}|_{\partial I^{\prime}},$ (46)
where $\Psi^{\dagger}$ is the exact state, then for $m\neq 0$, the forward operator $F$ with partial measurements satisfies the TCC (42).
Proof.
The proof is almost identical to that of Theorem 15. Note that S(p+h)S(p)S(p+h)-S(p) and $S^{\prime}(p)$ both satisfy homogeneous Cauchy boundary conditions on $\partial I^{\prime}$. This allows us to carry out partial integration as before and prove that
$\|\mathbf{B}_{m}(p)^{-1}\mathbf{B}_{m}^{\prime}(h)\|_{L^{2}(I^{\prime})\to L^{2}(I^{\prime})}$ $\leq\|\mathbf{B}_{m}(p)^{-1}\|_{H^{-4}(I^{\prime})\to L^{2}(I^{\prime})}\|\mathbf{B}_{m}^{\prime}(h)\|_{L^{2}(I^{\prime})\to H^{-4}(I^{\prime})}$
$\leq N^{\gamma^{\dagger},\Omega^{\dagger}}(R)C_{\text{emb}}\|h\|_{\mathbb{R}\times H^{2}(I^{\prime};m=0)}$
$\leq N^{\gamma^{\dagger},\Omega^{\dagger}}(R)C_{\text{emb}}\|h\|_{\mathbb{R}\times H^{2}({r{\mathbb{S}^{2}}};m=0)}$
for any $p$, $p+h$ in the ball $B^{X}_{R}(p^{\dagger})$. ∎
It is worth remarking that the constraint assumption in (46), which requires knowledge of the exact solution on the boundary, is an essential component to the proof of Theorem 17; relaxing this condition poses an interesting question for future research.
6 Numerical experiments
We implemented the accelerated Nesterov-Landweber method to simultaneously reconstruct the scalar viscosity and latitudinal differential rotation parameters, $(\gamma,\Omega)$. A convergence analysis of this method has been carried out in [19, 29] under the tangential cone condition, which we have been verified for our problem in Section 5. The Nesterov-Landweber iteration proceeds as follows: initiate $p_{-1}=p_{0}:=(\gamma_{\mathrm{int}},\Omega_{\mathrm{int}})$, then run
$z_{k}$ $=p_{k}+\frac{k-1}{k+\alpha-1}\left(p_{k}-p_{k-1}\right)\qquad$ $\alpha=3$ (47)
$p_{k+1}$ $=z_{k}-\mu_{k}F^{\prime}(z_{k})^{*}\left(F(z_{k})-y^{\delta}\right)$ $k\leq K(y^{\delta},\delta).$
Here, the forward map $F$ corresponds to the wave equation at single frequency $\omega$ and longitudinal wavenumber $m$. We allow for different observation strategies: full data, leaked data, and real part measurements. The Hilbert space adjoints $F^{\prime}(z_{k})^{*}$ are derived in Section 4.2. The inversion takes place in the following function spaces: $\gamma\in\mathbb{R}$, $\Omega\in H^{2}({r{\mathbb{S}^{2}}},m=0)$, thus ensuring convergence of the iterative reconstruction.
The numerical wave solution uses a fourth-order finite difference scheme [39] on a uniform grid of 100 points spanning the latitude range $(0,\pi)$. To generate ground truths while avoiding inverse crimes, we explicitly choose analytic $\Psi_{m}$$\gamma$ and $\Omega$ with Neumann (Fig. 1) or Dirichlet (Fig. 2) boundary, then explicitly obtain the corresponding source $f$. The discrete measurement $\underline{y}^{\delta}$ is corrupted by independent additive complex Gaussian noise with variance $\sigma^{2}$ on 100 equidistant measurement points and $\sigma$ chosen to have relative noise levels $\|\underline{y}^{\delta}-\underline{y}\|_{2}/\|\underline{y}\|_{2}\in\{0.01,0.05,0.2\}$.
We initialize the algorithm with $\gamma_{\mathrm{int}}=-3\gamma_{\mathrm{true}}$ and $\Omega_{\mathrm{int}}=0$, corresponding to no prior knowledge of the ground truth, and iterate with a step size $\mu_{k}$ determined by backtracking line search. The stopping criterion is the first instance when the residual falls below the threshold $\tau\delta$ (with $\tau=1.03$), which in practice yields better results than choosing $\tau>2$ [19]. The algorithm typically completes 200 iterations within one second on an i7-1255U CPU (4.70 GHz).
Figure 1: 1% data noise. Top to bottom: full and leaked data (filled area) at different levels: 20%, 50%, missing imaginary part.
Figure 2: Leaked data. Top to bottom: 1%, 5%, 10%, 20% relative noise level.
Full data.
Figure 1, row 1 shows the simultaneous recovery of $(\gamma,\Omega)$ from full measurement of the state $\Psi$ at $(\omega,m)=(3,3)$ and with $1\%$ data noise. Despite of an initial guess far from the ground truth, the reconstruction closely approximates the exact parameters. The reconstructed wave field also matches the true state in both real and imaginary parts.
Leaked data at poles.
Figure 1, rows 2–3 present the practical case where no data are available near the poles. In our numerical experiments, we did not impose additional Cauchy boundary data in Theorem 17, and simply computed the adjoint as described in Remark 13. With 20% of the data leaked, the reconstruction quality remains nearly as high as with full data. For 50% data loss, the recovery of $\Omega$ moderately deteriorates. Nonetheless, the algorithm exhibits robust performance, consistent with our convergence results in Section 5.2.
Real part measurement.
Figure 1, row 4 considers the scenario where only the real part of $\Psi$ is measured. Compared to row 3 (which also has 50% data leakage, but full complex data), the absence of the imaginary component further impairs the reconstruction, most notably in missing the secondary peak of the differential rotation.
Noisy data.
Finally, Figure 2 reports the results for $(\omega,m)=(1,2)$ with 20 % data leakage and different measurement noise levels. At 1% noise, the recovered parameters remain very close to the truth, as evidenced by small relative errors ($\||p-p_{\text{true}}\|_{L^{2}}/\|p_{\text{true}}\|_{L^{2}}$). At 5% noise, reconstruction quality is slightly reduced, while at 20% noise, the outcome degrades significantly.
7 Conclusions and Outlook
In this paper, we examined the modeling, as well as the associated forward and inverse problems, for inertial waves on the surface of the Sun. Formulating the dynamics in terms of a stream function leads to a fourth-order elliptic differential equation on the sphere. We rigorously proved well-posedness of this equation, including explicit smallness conditions for uniqueness, as well as conditions for non-uniqueness via analytic Fredholm theory, both in agreement with empirical findings in helioseismology. Our studies represent the starting point for further theoretical and numerical studies of resonances and inertial modes, as well as the associated inverse problems.
For the inverse problem of simultaneous identification of differential rotation and viscosity, we established convergence guarantees for the iterative reconstruction process via the tangential cone condition. Furthermore, we established local unique identifiability of rotation profile when viscosity is known, and vice versa. For the case of partial observations, the tangential cone condition required the additional assumption of exact observations of Cauchy boundary data. It is natural to ask whether this assumption can be relaxed, e.g., by including the Cauchy data in the range of the operator. We point out that at least for a finite number of $m$ and $\omega$, uniqueness of rotation at the unobserved latitudes seems unrealistic by dimensionality arguments.
Next steps include the incorporation of physical effects such as radial motion, pressure, and gravity [14] into the model and inversions with solar-like parameters, including realistic convection noise [32], as well as real observational data from the Helioseismic and Magnetic Imager. Another promising avenue involves leveraging data-driven model discovery techniques [1, 30] to refine existing theoretical models by learning hidden laws.
Appendix A Differential geometry and function spaces on spheres
Let $\mathbb{S}^{2}:=\{\textbf{{r}}\in\mathbb{R}^{3}:|\textbf{{r}}|=1\}$ denote the unit sphere in $\mathbb{R}^{3}$ and ${r{\mathbb{S}^{2}}}=\{r\textbf{{r}}:\textbf{{r}}\in\mathbb{S}^{2}\}=\{\textbf{{r}}\in\mathbb{R}^{3}:|\textbf{{r}}|=r\}$ the sphere of radius $r>0$. In this appendix, we recall some differential operators on $r\mathbb{S}^{2}$ and their properties as well as function spaces on $r{\mathbb{S}^{2}}$}, which are used throughout this paper.
Spherical coordinates.
We use $\textbf{{r}}(r,\theta,\phi):=(r\sin\theta\cos\phi,r\sin\theta\sin\phi,r\cos\theta)$ as definition of spherical coordinates with radial distance $r>0$, polar angle $\theta\in(0,\pi)$ and azimuthal angle $\phi\in(0,2\pi)$. In our notation, we do not distinguish between a function $f:\mathbb{R}^{3}\to\mathbb{R}$ and its counterpart $f\circ\textbf{{r}}:(0,\infty)\times(0,\pi)\times(0,2\pi)\to\mathbb{R}$ in spherical coordinates. Note that the vectors $\{\mathbf{e}_{r}(\textbf{{r}}),\mathbf{e}_{\theta}(\textbf{{r}}),\mathbf{e}_{\phi}(\textbf{{r}})\}$ defined by $\mathbf{e}_{r}(\textbf{{r}}):=\frac{\partial\textbf{{r}}}{\partial r}/|\frac{\partial\textbf{{r}}}{\partial r}|$ and similarly for $\mathbf{e}_{\theta}$ and $\mathbf{e}_{\phi}$ form a local orthonormal basis for all r.
Differential operators on the sphere.
A vector-valued function $\textbf{{f}}:{r{\mathbb{S}^{2}}}\to\mathbb{R}^{3}$ is called tangential if $\textbf{{f}}(\textbf{{r}})\cdot\mathbf{e}_{r}(\textbf{{r}})=0$ for all r. For a tangential vector field f and a scalar $f$ on $r{\mathbb{S}^{2}}$}, consider any smooth extensions $\tilde{\textbf{{f}}}$ and $\tilde{f}$ to a neighborhood of $r{\mathbb{S}^{2}}$} in $\mathbb{R}^{3}$. Then it is possible to define the surface (or horizontal) gradient $\operatorname{\mathbf{grad}_{\mathrm{h}}}f$, the surface divergence $\operatorname{div_{\mathrm{h}}}\textbf{{f}}$, and the scalar and vectorial surface curl $\operatorname{curl_{\mathrm{h}}}\textbf{{f}}$ and $\operatorname{\mathbf{curl}_{\mathrm{h}}}f$ such that
$\begin{aligned} &\operatorname{\mathbf{grad}_{\mathrm{h}}}f=\big{(}\operatorname{\mathbf{grad}}\tilde{f}-\frac{\partial\tilde{f}}{\partial r}\mathbf{e}_{r}\big{)}|_{{r{\mathbb{S}^{2}}}},&&\operatorname{div_{\mathrm{h}}}\textbf{{f}}=\operatorname{div}\tilde{\textbf{{f}}}|_{{r{\mathbb{S}^{2}}}},\\ &\operatorname{\mathbf{curl}_{\mathrm{h}}}f=\operatorname{curl}(\tilde{f}\mathbf{e}_{r})|_{{r{\mathbb{S}^{2}}}},&&\operatorname{curl_{\mathrm{h}}}\textbf{{f}}=(\operatorname{curl}\tilde{\textbf{{f}}}\cdot\mathbf{e}_{r})|_{{r{\mathbb{S}^{2}}}}\end{aligned}$ (48)
independent of the choices of $\tilde{f}$ and $\tilde{\textbf{{f}}}$ (see [28, p. 72–73]). Here $\operatorname{div_{\mathrm{h}}}\textbf{{f}}$ and $\operatorname{curl_{\mathrm{h}}}\textbf{{f}}$ are scalar functions, while $\operatorname{\mathbf{grad}_{\mathrm{h}}}f$ and $\operatorname{\mathbf{curl}_{\mathrm{h}}}v$ are tangential vector fields on $r{\mathbb{S}^{2}}$}. In spherical coordinates, writing $\textbf{{f}}=f_{\theta}\mathbf{e}_{\theta}+f_{\phi}\mathbf{e}_{\phi}$ with $f_{\theta}:=\textbf{{f}}\cdot\mathbf{e}_{\theta}$ and $f_{\phi}:=\textbf{{f}}\cdot\mathbf{e}_{\phi}$, we have
$\begin{aligned} &\operatorname{\mathbf{grad}_{\mathrm{h}}}f=\frac{1}{r}\frac{\partial f}{\partial\theta}\mathbf{e}_{\theta}+\frac{1}{r\sin\theta}\frac{\partial f}{\partial\phi}\mathbf{e}_{\phi},&&\operatorname{div_{\mathrm{h}}}\textbf{{f}}=\frac{1}{r\sin\theta}\left(\frac{\partial}{\partial\theta}(f_{\theta}\sin\theta)+\frac{\partial f_{\phi}}{\partial\phi}\right),\\ &\operatorname{\mathbf{curl}_{\mathrm{h}}}f=\frac{1}{r\sin\theta}\frac{\partial f}{\partial\phi}\mathbf{e}_{\theta}-\frac{1}{r}\frac{\partial f}{\partial\theta}\mathbf{e}_{\phi},&&\operatorname{curl_{\mathrm{h}}}\textbf{{f}}=\frac{1}{r\sin\theta}\left(\frac{\partial}{\partial\theta}(f_{\phi}\sin\theta)-\frac{\partial f_{\theta}}{\partial\phi}\right).\end{aligned}$ (49)
These differential operators satisfy the identities ([28, Thm. 2.5.19]):
$\begin{split}&\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}=0,\qquad\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}=0,\\ &\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=-\int_{{r{\mathbb{S}^{2}}}}f\operatorname{div_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s,\,\,\,\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=\int_{{r{\mathbb{S}^{2}}}}f\operatorname{curl_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s.\end{split}$ (50)
The Laplace-Beltrami operator on $r{\mathbb{S}^{2}}$} is given by
$\Delta_{\mathrm{h}}f=\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f=-\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}f,$ (51)
and the vector Laplacian or Hodge operator acting on tangential vector fields f is defined as
$\bm{\Delta}_{\mathrm{h}}\textbf{{f}}=\operatorname{\mathbf{grad}_{\mathrm{h}}}\operatorname{div_{\mathrm{h}}}\textbf{{f}}-\operatorname{\mathbf{curl}_{\mathrm{h}}}\operatorname{curl_{\mathrm{h}}}\textbf{{f}}.$ (52)
We also need a decomposition of $L^{2}_{\mathrm{tan}}({r{\mathbb{S}^{2}}}):=\{\textbf{{f}}\in L^{2}({r{\mathbb{S}^{2}}},\mathbb{R}^{3}:\textbf{{f}}\cdot\mathbf{e}_{r}\equiv 0\}$ into two orthogonal subspaces. Since $r{\mathbb{S}^{2}}$} is simply connected, there exist for each tangential vector field $\textbf{{f}}\in L^{2}_{\mathrm{tan}}({r{\mathbb{S}^{2}}})$ two functions $\varphi,\psi\in H^{1}({r{\mathbb{S}^{2}}})$ such that
$\textbf{{f}}=\operatorname{\mathbf{grad}_{\mathrm{h}}}\varphi+\operatorname{\mathbf{curl}_{\mathrm{h}}}\psi,$ (53)
and $\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\varphi,\operatorname{\mathbf{curl}_{\mathrm{h}}}\psi\rangle_{L^{2}}=0$ (see [28, Eq. (5.6.24)]).
Function spaces.
We define Sobolev spaces $H^{s}({r{\mathbb{S}^{2}}})$ as Hilbert scale generated by the operator $(I-\Delta_{\rm h})^{1/2}$:
$H^{s}({r{\mathbb{S}^{2}}}):=\mathrm{dom}((I-\Delta_{\rm h})^{s/2})\qquad\mbox{with}\quad\|f\|_{H^{s}}:=\|(I-\Delta_{\rm h})^{s/2}f\|_{L^{2}}$
for $s\geq 0$. It is straightforward that these are Hilbert spaces, and that $H^{0}({r{\mathbb{S}^{2}}}):=L^{2}({r{\mathbb{S}^{2}}})$. Moreover, we set $H^{-s}({r{\mathbb{S}^{2}}}):=H^{s}({r{\mathbb{S}^{2}}})^{*}$ with the natural norm. We further introduce subspaces of functions with mean 0 by
$L^{2}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in L^{2}({r{\mathbb{S}^{2}}}):\langle f,1\rangle=0\},\quad H^{s}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in H^{s}:\langle f,1\rangle=0\}$ (54)
for $s\in\mathbb{R}$. These spaces are also complete with the same norms, but we will later define a slightly more convenient equivalent norm on $H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$.
The following embeddings hold true (see [36, Propositions 3.2-3.3], [17, Theorem 3.5]):
$H^{s}({r{\mathbb{S}^{2}}})\hookrightarrow H^{k}({r{\mathbb{S}^{2}}})$ compact for $s>k$ (55a)
$H^{s}({r{\mathbb{S}^{2}}})\hookrightarrow L^{p}({r{\mathbb{S}^{2}}})$ continuous for $s\geq\max(0,1-\tfrac{2}{p})$ and $p\in[1,\infty)$ (55b)
$H^{s}({r{\mathbb{S}^{2}}})\hookrightarrow C^{k}({r{\mathbb{S}^{2}}})$ $\text{compact for $s>k+1$ with $k\geq 0$}.$ (55c)
Azimuthal Fourier transform.
Writing a function $f\in L^{2}({r{\mathbb{S}^{2}}})$ in spherical coordinates $f=f(\theta,\phi)$, we can perform a Fourier transform in $\phi$ to expand $f$ into a Fourier series
$f(\theta,\phi)=\frac{1}{\sqrt{2\pi}}\sum_{m=-\infty}^{m}\hat{f}_{m}(\theta)e^{im\phi}.$ (56)
Let us define $e_{m}(\phi):=(2\pi)^{-1/2}e^{im\phi}$ and $(\hat{f}_{m}\otimes e_{m})(\theta,\phi):=\hat{f}_{m}(\theta)e^{im\phi}$. Since $\hat{f}_{m}\otimes e_{m}\perp\hat{f}_{n}\otimes e_{n}$ in $L^{2}({r{\mathbb{S}^{2}}})$ and in $L^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ for $m\neq n$, the expression (56) induces orthogonal decompositions
$L^{2}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}({r{\mathbb{S}^{2}}};m),\quad L^{2}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ (57)
with $\hat{f}_{m}\otimes e_{m}\in L^{2}({r{\mathbb{S}^{2}}};m)$. As the spherical harmonics $\{Y_{lm}:l\in\mathbb{N}_{0},m\in\mathbb{Z},|m|\leq l\}$ form a complete orthonormal basis of $L^{2}({r{\mathbb{S}^{2}}})$ and $Y_{lm}\in L^{2}({r{\mathbb{S}^{2}}};m)$, we have
$L^{2}({r{\mathbb{S}^{2}}};m)=L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)=\overline{\mathrm{span}\{Y_{lm}:l\geq|m|\}}^{L^{2}},\qquad m\in\mathbb{Z}\setminus\{0\},$
$L^{2}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\}}^{L^{2}},\quad L^{2}_{\diamond}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\setminus\{0\}\}}^{L^{2}},$
where $\mathrm{span}$ denotes the set of finite linear combinations. The Laplace-Beltrami operator is diagonal with respect to the decompositions (57). More precisely, it follows from (48) and (51) that
$\Delta_{\mathrm{h}}f=\sum_{m=-\infty}^{\infty}(\Delta_{m}\hat{f}_{m})\otimes e_{m}\quad\mbox{with}\quad\Delta_{m}:=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\sin\theta\frac{d}{d\theta}\right)-\frac{m^{2}}{r^{2}\sin^{2}\theta}\,$ (58)
where by abuse of notation, we denote $\Delta_{m}(\hat{f}_{m}\otimes e_{m}):=(\Delta_{m}\hat{f}_{m})\otimes e_{m}$, thus also
$\operatorname{\mathbf{grad}_{\mathit{m}\!}}:=\left(\frac{1}{r}\frac{d}{d\theta};\frac{im}{r\sin\theta}\right).$ (59)
We obtain $H^{s}$-orthogonal decompositions
$H^{s}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H^{s}({r{\mathbb{S}^{2}}};m)\quad\mbox{and}\quad H^{s}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H_{\diamond}^{s}({r{\mathbb{S}^{2}}};m)$ (60)
for any $s\in\mathbb{R}$, where $H^{s}({r{\mathbb{S}^{2}}};m)$ and $H_{\diamond}^{s}({r{\mathbb{S}^{2}}};m)$ are the closures of $\mathrm{span}\{Y_{lm}:l\geq|m|\}$ in $H^{s}({r{\mathbb{S}^{2}}})$ and $H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$, respectively. Again, $H^{s}({r{\mathbb{S}^{2}}};m)=H_{\diamond}^{s}({r{\mathbb{S}^{2}}};m)$ for $m\in\mathbb{Z}\setminus\{0\}$ and $H^{s}({r{\mathbb{S}^{2}}};0)=H_{\diamond}^{s}({r{\mathbb{S}^{2}}};0)\oplus\mathrm{span}\{1\}$. As $\Delta_{\mathrm{h}}Y_{lm}(\cdot/r)=\frac{l(l+1)}{r^{2}}Y_{lm}(\cdot/r)$, and $Y_{00}(\cdot/r)\notin L^{2}_{\diamond}({r{\mathbb{S}^{2}}})$, it follows that $\langle-\Delta_{\mathrm{h}}f,f\rangle\geq 2/r^{2}\|f\|_{L^{2}}^{2}$ for all $f\in L^{2}_{\diamond}({r{\mathbb{S}^{2}}})$. Consequently,
$\|f\|_{H_{\diamond}^{s}}:=\|(-\Delta_{\mathrm{h}})^{s/2}f\|_{L^{2}},\qquad s\geq 0$
defines a norm on $H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$ which is equivalent to the $\|\cdot\|_{H^{s}}$-norm. We will consider this as the standard norm on $H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$ as it has the convenient properties
$\begin{aligned} &\|f\|_{H_{\diamond}^{2}}=\|\Delta_{\mathrm{h}}f\|_{L^{2}}\quad\mbox{and}\quad\\ &\|f\|_{H_{\diamond}^{1}}^{2}=\langle-\Delta_{\mathrm{h}}f,f\rangle=\langle-\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f,f\rangle=\|\operatorname{\mathbf{grad}_{\mathrm{h}}}f\|_{L^{2}}^{2}.\end{aligned}$ (61)
Defining $H^{-s}_{\diamond}({r{\mathbb{S}^{2}}}):=H^{s}_{\diamond}({r{\mathbb{S}^{2}}})^{\prime}$ for $s>0$ we have that
$\Delta_{\mathrm{h}}:H^{t}_{\diamond}({r{\mathbb{S}^{2}}})\to H^{t-2}_{\diamond}({r{\mathbb{S}^{2}}})\quad\text{is an isometric isomorphism for all $t\in\mathbb{R}$.}$ (62)
Defining $L^{p}({r{\mathbb{S}^{2}}};m)$ as the closure of $\mathrm{span}\{Y_{lm}:l\geq|m|\}$ in $L^{p}({r{\mathbb{S}^{2}}})$, the embeddings (55) implies the embeddings in the decomposed spaces
$H^{s}({r{\mathbb{S}^{2}}};m)\hookrightarrow H^{k}({r{\mathbb{S}^{2}}};m)$ compact for $s>k$ (63a)
$H^{s}({r{\mathbb{S}^{2}}};m)\hookrightarrow L^{p}({r{\mathbb{S}^{2}}};m)$ continuous for $s\geq\max(0,1-\tfrac{2}{p})$, $p\in[1,\infty)$. (63b)
Appendix B Boundary conditions
The following lemma describes homogeneous boundary conditions satisfied by the $\theta$-dependent factors of separated smooth solutions:
Lemma 18.
Let $m\in\mathbb{Z}$ and $k\in\mathbb{N}_{0}=\{0,1,2,\dots\}$. A function of the form $u(\theta,\phi)=u_{m}(\theta)\exp(im\phi)$ in spherical coordinates belongs to $C^{k}({r{\mathbb{S}^{2}}})$ if and only if the following three conditions are satisfied:
  1. (i)
    $u_{m}\in C^{k}([0,\pi])$,
  2. (ii)
    $u_{m}^{(j)}(\theta)=0$ for $\theta\in\{0,\pi\}$ and all $j\in\mathbb{N}_{0}$ with $j\leq k$ and $(-1)^{|m|+j}=-1$,
  3. (iii)
    $u_{m}^{(j)}(\theta)=0$ for $\theta\in\{0,\pi\}$ and all $j\in\mathbb{N}_{0}$ with $j<m$.
Proof.
It is clear that $u_{m}\in C^{k}([0,\pi])$ is equivalent to $u\in C^{k}({r{\mathbb{S}^{2}}}\setminus\{P_{+},P_{-}\})$ with the poles $P_{\pm}:=(0,0,\pm 1\}$. Moreover, as $u_{m}(\theta)=u(\theta,0)$, we have $u_{m}\in C^{k}([0,\pi])$ if $u\in C^{k}({r{\mathbb{S}^{2}}})$.
To study regularity at the pole $P_{+}$, consider the smooth map $M:D:=\{(x,y)\in\mathbb{R}^{2}:x^{2}+y^{2}\leq r^{2}/2\}\to{r{\mathbb{S}^{2}}}$ of $r{\mathbb{S}^{2}}$} defined by $M(x,y):=\sqrt{r^{2}-x^{2}-y^{2}}$. By definition, $u$ is $k$-times differentiable at $P_{+}$ if and only if $u\circ M$ is $k$ times differentiable at 0. And this is the case if and only if there exists a polynomial $p(x,y)=\sum_{j=0}^{k}\sum_{n=0}^{j}c_{j,n}x^{n}y^{j-n}$ of degree $\leq k$ such that
$\left|(u\circ M-p)(x,y)\right|=o\left(\sqrt{x^{2}+y^{2}}^{k}\right)\qquad\text{as }(x,y)\to 0.$ (64)
Writing $x=\frac{\rho}{2}(e^{i\phi}+e^{-i\phi})$ and $y=\frac{i\rho}{2}(e^{i\phi}-e^{-i\phi})$ in polar coordinates and using binomial expansions, we obtain
$\sum_{n=0}^{j}c_{j,n}x^{n}y^{j-n}=\rho^{j}\sum_{l=0}^{j}d_{j,l}e^{i(j-2l)\phi}$
for some coefficients $d_{j,l}\in\mathbb{C}$. As $(u\circ M)(\rho\cos\phi,\rho\sin\phi)=u_{m}(\rho)\exp(m\phi)$, the coefficients $d_{j,l}$ must vanish unless j2l=mj-2l=m, so
$u_{m}(\rho)=\sum_{l\geq 0,m+2l\leq k}d_{m+2l,l}\rho^{m+2l}+o(\rho^{k})\qquad\text{as }\rho\to 0.$
This shows that $u$ is $k$-times differentiable at $P_{+}$ if and only if $u_{m}$ satisfies the conditions ii and iii at $\theta=0$. Writing derivatives of $u$ of order $\leq k$ in terms of derivatives of $u_{m}$, we see these derivatives are continuous if $u_{m}\in C^{k}([0,\pi])$. The argument for the pole $P_{-}$ is analogous, with replacing the map $M$ by M-M. ∎
With the help of the following lemma, it can be shown that classical solutions to (9) are also weak solutions:
Lemma 19.
For all $m\in\mathbb{Z}$ the operator $\Delta_{m}^{2}$ is symmetric on the set
$\left\{u\in C^{4}([0,\pi]):\Gamma_{m}u=0,\Delta_{m}^{2}u\in L^{2}([0,\pi],r^{2}\sin)\right\}\subset L^{2}([0,\pi],r^{2}\sin).$
Proof.
Recalling the separated Laplacian in (58) and the weighted $L^{2}_{\diamond}$-inner product allow us to perform the integration by part as
$\left\langle\Delta_{m}^{2}u,v\right\rangle$ $=-\left\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Delta_{m}u,\operatorname{\mathbf{grad}_{\mathit{m}\!}}v\right\rangle+\left[\sin\theta\,\partial_{\theta}(\Delta_{m}u)\bar{v}\right]_{\theta\to 0}^{\theta\to\pi}$
$=\left\langle\Delta_{m}u,\Delta_{m}v\right\rangle+\left[\sin\theta\,\partial_{\theta}(\Delta_{m}u)\bar{v}-\sin\theta\,\Delta_{m}u\,\partial_{\theta}\bar{v}\right]_{\theta\to 0}^{\theta\to\pi}$
$=\left\langle\Delta_{m}u,\Delta_{m}v\right\rangle+\frac{1}{r^{2}}\sum\nolimits_{j=1}^{5}[a_{j}(\theta)]_{\theta\to 0}^{\theta\to\pi}+\frac{m^{2}}{r^{2}}\sum\nolimits_{j=1}^{3}[b_{j}(\theta)]_{\theta\to 0}^{\theta\to\pi}$
with
$a_{1}=-\cos\theta\partial_{\theta}u\partial_{\theta}\bar{v},$ $a_{2}=-\sin\theta\partial_{\theta}^{2}u\partial_{\theta}\bar{v},$ $a_{3}=-\frac{1}{\sin\theta}\partial_{\theta}u\bar{v},$ $a_{4}=\cos\theta\partial_{\theta}^{2}u\bar{v},$
$a_{5}=\sin\theta\partial_{\theta}^{3}u\bar{v},$ $b_{1}=\frac{-1}{\sin\theta}\partial_{\theta}u\bar{v},$ $b_{2}=\frac{2\cos\theta}{\sin^{2}\theta}u\bar{v},$ $b_{3}=\frac{1}{\sin\theta}u\partial_{\theta}\bar{v}$
for $u,v\in H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$. Denoting the poles by $p\in\{0,\pi\}$ and considering $v$ having the same boundary conditions as $u$, we can see as follows that the boundary terms vanish for all $m\in\mathbb{Z}$:
$|m|\geq 2$:
Here $u(p)=u^{\prime}(p)=0=v(p)=v^{\prime}(p)$ such that the limits for all the $a_{j}$-terms vanish. Moreover, we observe from l’Hospital’s rule that $\lim_{\theta\to p}b_{2}(\theta)=2u^{\prime}(p)\bar{v}^{\prime}(p)=0$ and $\lim_{\theta\to p}b_{j}(\theta)=0$ for $j\in\{1,3\}$ such that all $b_{j}$-limits vanish.
$|m|=1$:
Here $u(p)=u^{\prime\prime}(p)=0=v(p)=v^{\prime\prime}(p)$, and hence $\lim_{\theta\to p}a_{j}(p)=0$ for $j\in\{2,4,5\}$. Moreover, $\lim_{\theta\to p}(a_{1}+a_{3})(\theta)=-2\cos(p)u^{\prime}(p)\bar{v}^{\prime}(p)=-\lim_{\theta\to p}b_{2}(\theta)$, and $\lim_{\theta\to p}(b_{1}+b_{3})(\theta)=0$ by l’Hospital.
$m=0$:
Here $u^{\prime}(p)=u^{\prime\prime\prime}(p)=0=v^{\prime}(p)=v^{\prime\prime\prime}(p)$, and we only need to discuss the $a_{j}$-terms, which vanish in the limit $\theta\to p$ for $j\in\{1,2,5\}$. Moreover, by l’Hospital $\lim_{\theta\to p}a_{3}(\theta)=-\cos(p)u^{\prime\prime}(p)\bar{v}(p)=-\lim_{\theta\to p}a_{4}(p)$. ∎
Remark 20 (Restricted domain – Cauchy boundary).
It is clear that on the restricted domain $I^{\prime}$ as in Section 3.5 with boundary points $p\in\partial I^{\prime}=\{\epsilon,\pi-\epsilon\}$, the symmetry property remains valid with Cauchy boundary conditions $u(p)=u^{\prime}(p)=0$.

### v3

@@
-Viscous-inertial waves on the surface of the Sun: modeling, forward and inverse problems
+Linear toroidal-inertial waves on a differentially rotating sphere with application to helioseismology: Modeling, forward and inverse problems
 Abstract
-This paper develops a mathematical framework for studying the newly discovered solar inertial oscillations, offering promising new avenues for exploring the Sun’s dynamics.
-Under the assumption of purely toroidal motions, the stream function of the flow satisfies a fourth-order scalar equation governing inertial waves on the rotating Sun.
+This paper develops a mathematical framework for interpreting observations of solar inertial waves in an idealized setting.
+Under the assumption of purely toroidal linear waves on the sphere, the stream function of the flow satisfies a fourth-order scalar equation.
 We prove well-posedness of wave solutions under explicit conditions on differential rotation.
-Moreover, we study the inverse problem of simultaneously reconstructing viscosity and differential rotation parameters from either complete or partial surface observations.
-To this end, we verify the tangential cone condition, ensuring the convergence of iterative regularization methods, furthermore the local unique identifiability of the unknown parameters.
-Numerical experiments employing the Nesterov-Landweber iteration confirm robustness of the reconstruction across different observation schemes and noise levels.
+Moreover, we study the inverse problem of simultaneously reconstructing viscosity and differential rotation parameters from either complete or partial surface data.
+We establish convergence guarantee of iterative regularization methods by verifying the tangential cone condition, and prove local unique identifiability of the unknown parameters.
+Numerical experiments with Nesterov-Landweber iteration confirm reconstruction robustness across different observation strategies and noise levels.
 Keywords.
-inverse problems, differential equations on manifolds, forth order elliptic differential equations, helioseismology, inertial waves, iterative regularization, tangential cone condition
+Inverse problems, differential equations on manifolds, fourth-order elliptic differential equations, helioseismology, inertial waves, iterative regularization, tangential cone condition.
 Acknowledgement.
 Support from Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1456/432680300 Mathematics of Experiment, Project C04 is gratefully acknowledged.
+LG and DF acknowledge funding from the ERC Synergy Grant WHOLESUN 810218.
 1 Introduction
 Helioseismology is the study of the internal structure and dynamics of the Sun from observations of solar oscillations on the surface.
-Such inference is primarily done by analyzing the observed pressure wave (p-modes) and has led to many achievements including the determination of the Sun’s differential rotation [37].
-However, the sensitivity of p-modes decreases with depth exacerbating the ill-posedness of the inversion.
-It is this ill-posedness that motivates the present article, and its accompanying analysis of regularized recovery of rotation and viscosity on the Sun.
-The recent discovery of numerous inertial modes on the surface of the Sun [15] opens new possibilities for helioseismology.
-Inertial oscillations, whose frequencies are in the range of hundreds of nHz, are caused by the Coriolis force, exhibiting periodic behavior on the timescale of the solar rotation cycle.
-As a consequence, these modes require very long observation time to be identified.
-Indeed, their discovery relies on 10 years of data acquisition by the Helioseismic and Magnetic Imager on board the Solar Dynamics Observatory [34].
-As many of these modes have maximum energy at the base of the convection zone [15], i.e., at a depth of about 0.71 solar radii, their study promises to give new insight into the dynamics of the deeper layers of the Sun.
-We will use these newly discovered waves as data of the inversion process.
-Most of the observed modes are quasi-toroidal, i.e.
-almost divergence-free, with very small radial velocity component, giving rise to the use of a 2D model for inertial waves on a rotating sphere.
-In this framework, we can reduce the vectorial wave equation into a scalar equation employing a stream function, the divergence-free component of the Helmholtz decomposition of the flow.
-This idea was proposed in [40] and extended in [7] to take into account viscous stress, an important component in solar applications [12].
-Inspired by these works, we will model the propagation of inertial-viscous waves on the surface of the rotating Sun by the stream function.
-The vectorial motion equation driven by Coriolis force and viscous stress will be turn into a fourth-order scalar equation in a low-dimensional setting.
-Moreover, we derive the separated equations for each longitudinal-wavenumber $m$ with realistic and physical meaningful $m$-dependent boundary conditions; this serves as the underlying model for the inverse parameter identification problem.
-The problem of solar inversion using inertial waves is completely novel as these modes have been discovered only recently [15].
-Here, we take a first step towards recovering the surface differential rotation and the viscosity from Dopplergrams.
-Doppler data are usually not used directly in helioseismology due to the high level of convective noise; indeed, solar excitations can be viewed as random turbulence, resulting in random oscillations.
-However, some inertial modes have been detected directly in the Doppler signal [14] motivating this study using direct measurements of inertial waves as input data.
-Helioseismology predominantly solves linear inverse problem, under the assumption that perturbations are relatively small, as it is the case, for example, for meridional flows [13].
-However, for parameters such as viscosity, where no suitable initial guess is available, this linearization assumption may no longer be valid.
-Here, following a related approach for pressure waves and correlation data in [27], we solve the nonlinear inversion for rotation and viscosity using inertial waves and iterative regularization reconstruction methods.
+Such inference is primarily done by analyzing the observed pressure waves (p-modes) with periods of order five minutes.
+Acoustic-mode helioseismology based on the interpretation of the mode frequencies has led to many achievements, including the determination of the Sun’s differential rotation as a function of radius and unsigned latitude [Thompson03].
+Inversions of the two-point correlations of the wave field at the surface can also be used to infer 3D perturbations in sound speed and flows in the solar interior [GizonBirchReview, Bjoern23].
+The recent discovery of global Rossby modes on the Sun [Loptien2018] has opened a new research topic: inertial-mode helioseismology.
+Solar inertial oscillations are modes restored by the Coriolis force, with periods spanning several weeks, comparable to solar rotation period of $\sim 27$ days at the equator.
+These modes require very long observations from space-based and ground-based observatories to be studied with sufficient frequency resolution; 50 years of data are now available [Gizon2024, Liang2025].
+Since many of these modes have maximum kinetic energy deep in the convective envelope of the Sun [GizonInertial21], their study promises to reveal new insights into the physics and dynamics of the solar interior.
+Current efforts focus on the development of simplified physical models of these inertial modes to investigate their unique sensitivity to internal properties of the Sun, such as differential rotation at very high latitudes and the internal turbulent viscosity.
+Several highly idealized 2D models focusing on purely toroidal modes have been developed for both $\beta$-plane [Gizon2020] and spherical geometries [Damien22].
+The dynamics of these retrograde-propagating modes is governed by the functional form of the background latitudinal differential rotation, with slower rotation at high latitudes than at the equator.
+Particularly, 2D models have demonstrated the existence of viscous critical latitudes for most inertial modes.
+In this article, by employing stream function formulation in a 2D spherical framework, we reduce the vectorial viscous-inertial wave equation to a fourth-order scalar equation of Orr-Sommerfeld type [Gizon2020, Damien22].
+Additionally, we derive the separated equations for each longitudinal-wavenumber $m$ with realistic and physically meaningful $m$-dependent boundary conditions at the poles.
+The separated equations serve as the underlying models for our inverse problem: the retrieval of differential rotation and turbulent viscosity.
+Since solar inertial wave inversion is completely novel in helioseismology, this work establishes the first step towards extracting these quantities from observed surface horizontal velocities.
 Contributions.
-This article provides a complete treatment – modeling, forward problem and inverse problem – for the newly observed inertial waves on the Sun:
-After a rigorous derivation of the forward model already used in [12], we prove existence, uniqueness and stability of both, the full and separated the inertial wave equations.
-This well-posedness is derived under explicit conditions on the relative smallness of variations of the the latitudinal-dependent rotation $\Omega$ compared to the product of the frequency and the third power of the viscous parameter $\gamma$.
-In passing, these results also allow us to apply analytic Fredholm theory, providing mathematical foundations to the recent study of inertial modes.
-The results of our analysis, both on the conditions for uniqueness and on the isolated occurrence of inertial modes, agree with state-of-the-art interpretations in inertial mode helioseismology.
-For the inverse problem of the simultaneous identification of viscosity and differential rotation, we lay out foundations of iterative regularization methods by characterizing adjoint operators on a continuous level.
-Crucially, we provide convergence guarantees for these methods – for full and, to a certain extent, also for restricted observations.
-We achieve this by establishing the important tangential cone condition via a lifted regularity strategy, accounting for realistic $L^{2}$-data.
-In addition, we prove local unique identifiability of $\Omega$ when $\gamma$ is known, and vice versa, from full measurement.
+This article provides a comprehensive treatment – modeling, forward problem and inverse problem – of purely toroidal inertial modes governed by a time harmonic fourth-order equation on a differentially rotating sphere.
+After a derivation of the forward model used in [Damien22], we prove existence, uniqueness and stability of full and separated inertial wave equations.
+Well-posedness is guaranteed under explicit conditions that the latitudinal-dependent rotation $\Omega$ is small relative to the product of the frequency and the third power of the viscousity $\gamma$.
+These results enable the application of analytic Fredholm theory, providing a mathematical foundation for the model [Damien22], and rigorously validating structure of discrete isolated inertial modes.
+We develop a regularization framework for the inverse problem of simultaneously reconstructing viscosity and differential rotation.
+The framework is built on characterization of continuous adjoint operators and convergence guarantees for iterative regularization methods – a result we obtain for both full and, under certain condition, for restricted observations.
+Central to our convergence analysis is the establishment of the tangential cone condition through a lifted regularity strategy, accommodating realistic $L^{2}$-data.
+Furthermore, we prove local unique identifiability: $\Omega$ is uniquely identifiable when $\gamma$ is known, and vice versa under full measurement.
 Structure.
+The article is organized as follows.
 Section 2 introduces the modeling framework for inertial oscillations and outlines different observation strategies for solar data.
-The well-posedness and regularity of the resulting wave solutions are derived in Section 3.
-Section 4 addresses the inverse problem for viscosity and differential rotation, including sensitivity analysis and derivation of associated adjoint operators.
-In Section 5, we establish convergence guarantees of iterative reconstruction schemes and local unique identifiability of the unknown parameters based on the tangential cone condition.
-Finally, Section 6 presents numerical reconstruction results under different scenarios, and Section 7 concludes our findings with a summary and an outlook on future research directions.
+Section 3 establish well-posedness and regularity of the resulting wave solutions.
+Section 4 addresses the inverse problem for viscosity and differential rotation, including sensitivity analysis and adjoint operator derivation.
+In Section 5, we develop convergence guarantees for iterative reconstruction methods and prove local unique identifiability via the tangential cone condition.
+Section 6 demonstrate numerical reconstruction performance across different measurement scenarios.
+Finally, Section 7 concludes our findings and future directions.
 2 Model and observation of inertial waves
+We begin with deriving a model for dynamics of linear inertial waves on the Sun in a rotating frame, incorporating latitudinal differential rotation and eddy viscosity.
+The theoretical framework is developed in Sections 2.12.2, while observations of solar data are discussed in Section 2.3.
 2.1 Reduced-order modeling of inertial waves
-In this section, we derive a physical model for dynamics of waves in the presence of latitudinal differential rotation and eddy viscosity on the Sun in a rotating frame.
 Linearized Navier Stokes equations.
-We begin with the equation of motion (momentum equation), in which the dynamics of a moving particle of density $\rho$ in an incompressible fluid is subject to linear viscous stress $\mathbf{\bm{\tau}}$ with viscosity $\gamma$, acoustic pressure $p$, gravity $\mathbf{g}$, and external source $\mathbf{f}$:
+We begin with the equation of motion (momentum equation), in which the dynamics of a moving particle of density $\rho$ is subject to linear viscous stress $\mathbf{\bm{\tau}}$ with viscosity $\gamma$, acoustic pressure $p$, gravity $\mathbf{g}$, and external source $\mathbf{f}$:
 $\begin{aligned} &\rho\left(\frac{\partial\textbf{{v}}}{\partial t}+\nabla_{\!\textbf{{v}}}\textbf{{v}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{v}}\right)=\operatorname{\mathbf{div}}(\rho\gamma\mathbf{\bm{\tau}})-\operatorname{\mathbf{grad}}p+\rho\mathbf{g}+\rho\mathbf{f},\\ &\mathbf{\bm{\tau}}:=\operatorname{\mathbf{grad}}\textbf{{v}}+(\operatorname{\mathbf{grad}}\textbf{{v}})^{\top}.\end{aligned}$ (1a)
 Here, $\textbf{{v}}=\textbf{{v}}(\textbf{{r}},t)\in\mathbb{R}^{3}$ is the vector velocity of a particle at position $\textbf{{r}}\in\mathbb{R}^{3}$ and time $t\in(0,T)$, while $\nabla_{\!\textbf{{v}}}:=\sum_{i=1}^{3}v_{i}\frac{\partial}{\partial x_{i}}$ applies componentwise.
-We work in the frame rotating at the angular frequency $\Omega_{\mathrm{{ref}}}$ around a fixed axis $\mathbf{\hat{a}}$ and set $\mathbf{\Omega}_{\mathrm{ref}}=\Omega_{\mathrm{{ref}}}\,\mathbf{\hat{a}}$.
+We work in the frame rotating at the constant angular frequency $\Omega_{\mathrm{{ref}}}$ around a fixed axis $\mathbf{\hat{a}}$ and set $\mathbf{\Omega}_{\mathrm{ref}}=\Omega_{\mathrm{{ref}}}\,\mathbf{\hat{a}}$.
 Equation (1a) corresponds to the Navier-Stokes equation in a rotating frame.
@@
 The other additional fictitious forces, Euler force $\rho\left(\frac{d\mathbf{\Omega}_{\mathrm{ref}}}{dt}\times\textbf{{r}}\right)$ and centrifugal force $\rho\mathbf{\Omega}_{\mathrm{ref}}\times(\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{r}})$, are ignored as $\Omega_{\mathrm{{ref}}}$ is uniform in time and assumed to be small.
-For solar applications, one usually chooses the Carrington frame as reference frame with the angular velocity $\Omega_{\mathrm{{ref}}}=14.7^{0}/$day and the rotation axis $\mathbf{e}_{z}=[0,0,1]^{\top}$.
+For solar applications, one usually chooses the Carrington frame as reference frame with the angular velocity $\Omega_{\mathrm{{ref}}}=14.7$ deg/day and the rotation axis $\mathbf{e}_{z}=[0,0,1]^{\top}$.
 Equation (1a) is complemented by the continuity equation (conservation of mass):
@@
 We linearize Eqs. (1) around a stationary background (unperturbed, equilibrium, mean) medium characterized by $\textbf{{u}}_{0}$}, $p_{0}$, $\rho_{0}$, and $\mathbf{g}_{0}$ such that
-$\textbf{{v}}(\textbf{{r}},t)={\textbf{{u}}_{0}}(\textbf{{r}})+\textbf{{u}}(\textbf{{r}},t),\quad\rho(\textbf{{r}},t)=\rho_{0}(\textbf{{r}})+\rho^{\prime}(\textbf{{r}},t),$
-$p(\textbf{{r}},t)=p_{0}(\textbf{{r}})+p^{\prime}(\textbf{{r}},t),\quad\mathbf{g}(\textbf{{r}},t)=\mathbf{g}_{0}(\textbf{{r}})+\mathbf{g}^{\prime}(\textbf{{r}},t),$
+$\textbf{{v}}(\textbf{{r}},t)={\textbf{{u}}_{0}}(\textbf{{r}})+\textbf{{u}}(\textbf{{r}},t),\qquad\rho(\textbf{{r}},t)=\rho_{0}(\textbf{{r}})+\rho^{\prime}(\textbf{{r}},t),$
+$p(\textbf{{r}},t)=p_{0}(\textbf{{r}})+p^{\prime}(\textbf{{r}},t),\qquad\mathbf{g}(\textbf{{r}},t)=\mathbf{g}_{0}(\textbf{{r}})+\mathbf{g}^{\prime}(\textbf{{r}},t),$
 where u is the (perturbed) wave velocity, and ${\textbf{{u}}_{0}}(\textbf{{r}}):=(\mathbf{\Omega}(\textbf{{r}})-\mathbf{\Omega}_{\mathrm{ref}})\times\textbf{{r}}$ is assumed to be a differential rotation field with angular velocity $\mathbf{\Omega}(\textbf{{r}})=\Omega(r,\theta)\mathbf{e}_{z}$ depending only on radius $r$ and colatitude $\theta$ in spherical coordinates $r;\theta;\phi]\in[0,\infty)\times[0,\pi)\times[0,2\pi)$.
+This linearization is possible when no linearly unstable mode exists.
+In this regime, there is no growing (in time) solutions to Eq.
+(1).
+Note that such solutions exists in nonlinear simulations with solar-like differential rotation [Bekki2024], however this regime is not included in this study.
+Background medium.
 We assume that the background medium solves the second-order quasilinear elliptic boundary value problem
@@
 To construct such a background model for the Sun, it is usually assumed that the background flows are weak and the background medium is built from the hydrostatic equilibrium, that is $\operatorname{\mathbf{grad}}p_{0}=\rho_{0}\mathbf{g}_{0}$.
-This assumption leads to background coefficients that depend only on depth and are given by standard solar models such as model S [8].
+This assumption leads to background coefficients that depend only on depth and are given by standard solar models such as Model S [modelS].
 Starting from a radial model for density, it is possible to solve Poisson equation to obtain the gravitational potential and deduce the background pressure.
+First-order equations.
+The first-order the system (1) then becomes
+$\rho_{0}\left(\frac{\partial\textbf{{u}}}{\partial t}+\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}+\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{u}}\right)=-\operatorname{\mathbf{grad}}p^{\prime}+\rho^{\prime}\mathbf{g}_{0}+\rho_{0}\mathbf{g}^{\prime}+\operatorname{\mathbf{div}}\left[\rho_{0}\gamma\bm{\tau}^{\prime}\right]+\rho_{0}\mathbf{f},$ (3a)
+$\frac{\partial\rho^{\prime}}{\partial t}+\mathrm{div}(\rho^{\prime}\mathbf{u}_{0})+\mathrm{div}(\rho_{0}\mathbf{u})=0,$ (3b)
+where $\bm{\tau}^{\prime}:=\operatorname{\mathbf{grad}}\textbf{{u}}+(\operatorname{\mathbf{grad}}\textbf{{u}})^{\top}$. It is complemented by the linearized equation of state
+$\frac{p^{\prime}}{p_{0}}=\gamma\frac{\rho^{\prime}}{\rho_{0}}+\frac{s^{\prime}}{c_{v}},$
+where $\gamma=5/3$ is the ratio of specific heats, the quantity $c_{v}$ is the specific heat at constant volume and $s^{\prime}$ denotes the perturbed entropy.
+The first two terms of the right hand side of Eq.
+3a can be rewritten as
+$\operatorname{\mathbf{grad}}p^{\prime}+\rho^{\prime}\mathbf{g}_{0}=-\rho_{0}\operatorname{\mathbf{grad}}\left(\frac{p^{\prime}}{\rho_{0}}\right)+p^{\prime}\frac{N^{2}}{\mathbf{g}},$ (4)
+where we used hydrostatic equilibrium and defined the buoyancy frequency $N$ such that
+$N^{2}=\mathbf{g}\cdot\left(-\frac{\nabla\rho_{0}}{\rho_{0}}+\frac{\nabla p_{0}}{\gamma p_{0}}\right).$ (5)
 Assumption:
-Waves have small amplitude in comparison to the mean flow, i.e.
-$|\textbf{{u}}|\ll|{\textbf{{u}}_{0}}|$.
-Moreover, the anelastic approximation $\rho^{\prime}\approx 0$ and the Cowling approximation $\mathbf{g}^{\prime}\approx 0$ (see [9]) hold true.
-The anelastic approximation is justified if the flow speed is significantly smaller than the sound speed, which is adequate for inertial modes – a few meters per second versus kilometers per second for the surface sound speed.
-The system (1) thus becomes
-$\rho_{0}\left(\frac{\partial\textbf{{u}}}{\partial t}+\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}+\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{u}}\right)=-\operatorname{\mathbf{grad}}p^{\prime}+\operatorname{\mathbf{div}}\left[\rho_{0}\gamma\bm{\tau}^{\prime}\right]+\rho_{0}\mathbf{f},$ (3a)
+We employ the Cowling approximation with $\mathbf{g}^{\prime}=0$ (see [Cowling1941]), which is classic in helioseismology.
+We further assume the anelastic approximation which is justified if the flow speed is significantly smaller than the sound speed.
+This is adequate for inertial modes – flow speeds of a few meters per second compared to kilometers per second for the surface sound speed – and has been numerically verified [Sup2025].
+Two additional assumptions enable reduction to a 2D problem on the sphere: adiabaticity ($s^{\prime}=0$), and strongly stratified medium (see also [Watson81]) so that the term in $N^{2}$ in Eq.
+4 can be neglected.
+Under these assumptions, the system (3) becomes
+$\frac{\partial\textbf{{u}}}{\partial t}+\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}+\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{u}}=-\operatorname{\mathbf{grad}}\left(\frac{p^{\prime}}{\rho_{0}}\right)+\frac{1}{\rho_{0}}\operatorname{\mathbf{div}}\left[\rho_{0}\gamma\bm{\tau}^{\prime}\right]+\mathbf{f},$ (6a)
 $\operatorname{div}(\rho_{0}\textbf{{u}})=0,$ (6b)
-where $\bm{\tau}^{\prime}:=\operatorname{\mathbf{grad}}\textbf{{u}}+(\operatorname{\mathbf{grad}}\textbf{{u}})^{\top}$.
-The flow field $\textbf{{u}}_{0}$}, equivalently $\Omega$, and the viscosity $\gamma$ will be the quantities of interest in the inverse problem.
-Stratified equations.
-We first rewrite the left-hand side of (3a) in spherical coordinates, i.e.
-in terms of the components $u_{r},u_{\theta},u_{\phi}$ of the expansion $\textbf{{u}}=u_{r}\mathbf{e}_{r}+u_{\theta}\mathbf{e}_{\theta}´+u_{\phi}\mathbf{e}_{\phi}$ into the local orthonormal basis vectors $\mathbf{e}_{r},\mathbf{e}_{\theta}$, and $\mathbf{e}_{\phi}$ as defined in the appendix A.
+where the flow field $\textbf{{u}}_{0}$}, equivalently $\Omega$, and the viscosity $\gamma$ will be the quantities of interest in the inverse problem.
+First-order equations on the sphere.
+We now rewrite the left hand side of (6a) in spherical coordinates using the components $u_{r},u_{\theta},u_{\phi}$ in the expansion $\textbf{{u}}=u_{r}\mathbf{e}_{r}+u_{\theta}\mathbf{e}_{\theta}´+u_{\phi}\mathbf{e}_{\phi}$, where $\mathbf{e}_{r}$, $\mathbf{e}_{\theta}$, $\mathbf{e}_{\phi}$ are the local orthonormal basis vectors defined in the appendix A.
 Noting that ${\textbf{{u}}_{0}}=(\Omega(r,\theta)-\Omega_{\mathrm{{ref}}})\sin\theta\mathbf{e}_{\phi}$ and using the identities
@@
 $=$ $\mathbf{D}_{t}\begin{bmatrix}u_{r}\\ u_{\theta}\\ u_{\phi}\end{bmatrix}+\begin{bmatrix}0&0&-2\Omega\sin\theta\\ 0&0&-2\Omega\cos\theta\\ \left(2\Omega\sin\theta+\frac{\partial\Omega}{\partial r}r\sin\theta\right)&\left(2\Omega\cos\theta+\frac{\partial\Omega}{\partial\theta}\sin\theta\right)&0\end{bmatrix}\begin{bmatrix}u_{r}\\ u_{\theta}\\ u_{\phi}\end{bmatrix}$, (7)
-where $\mathbf{D}_{t}:=\frac{\partial}{\partial t}+(\Omega-\Omega_{\mathrm{{ref}}})\frac{\partial}{\partial\phi}$ is the material derivative.
-We now introduce the $3\times 2$-matrix $\mathbf{R}(\textbf{{r}}):=[\mathbf{e}_{\theta}(\textbf{{r}}),\mathbf{e}_{\phi}(\textbf{{r}})]$.
+where $\mathbf{D}_{t}:=\frac{\partial}{\partial t}+(\Omega-\Omega_{\mathrm{{ref}}})\frac{\partial}{\partial\phi}$ is the material derivative acting on vector fields.
+Define the $3\times 2$-matrix $\mathbf{R}(\textbf{{r}}):=[\mathbf{e}_{\theta}(\textbf{{r}}),\mathbf{e}_{\phi}(\textbf{{r}})]$.
 Then multiplication by the $3\times 3$ matrix $\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}$ is the projection onto the tangential subspace and equivalent to the application of $\textbf{{r}}\times\textbf{{r}}\times$.
 By applying $\textbf{{r}}\times\textbf{{r}}\times$ to (6a), we wish to obtain an approximate equation for the horizontal velocity components $\textbf{{u}}_{\mathrm{h}}(\textbf{{r}}):=-\textbf{{r}}\times\textbf{{r}}\times\textbf{{u}}(\textbf{{r}})$ given by $\textbf{{u}}_{\mathrm{h}}=\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}\textbf{{u}}=u_{\theta}\mathbf{e}_{\theta}+u_{\phi}\mathbf{e}_{\phi}$.
-To this end, we need the following assumption:
+To this end, we require the following assumption:
 Assumption:
@@
 The fluid is strongly stratified in the sense that the radial motion $u_{r}$ and its derivatives are small compared to the horizontal components $u_{\phi},u_{\theta}$.
-Due to this assumption, we can neglect the first column of the matrix in Eq.
+Under this assumption, one can neglect the first column of the matrix in Eq.
 (7), and the application of $\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}$ “kills” the first row of this matrix (the first element of the vector in (7) may be balanced by $\frac{\partial p^{\prime}}{\partial r}$).
@@
 $\mathbf{M}(\textbf{{r}}):=\mathbf{R}(\textbf{{r}})\begin{bmatrix}0&-2\Omega(r,\theta)\cos\theta\\ 2\Omega(r,\theta)\cos\theta+\frac{\partial\Omega}{\partial\theta}(r,\theta)\sin\theta&0\end{bmatrix}\mathbf{R}(\textbf{{r}})^{\top}$
-and obtain the following system of equations for the tangential vector field $\textbf{{u}}_{\mathrm{h}}$ on $r{\mathbb{S}^{2}}$} for a fixed $r>0$ with right hand side $\textbf{{f}}_{\!\mathrm{h}}:=\mathbf{R}\mathbf{R}^{\top}\textbf{{f}}$:
+and obtain the following system of equations for the tangential vector field $\textbf{{u}}_{\mathrm{h}}$ on the sphere $r{\mathbb{S}^{2}}$} for a fixed radius $r>0$ as:
 $-\rho_{0}\gamma\bm{\Delta}_{\mathrm{h}}\textbf{{u}}_{\mathrm{h}}+\rho_{0}\mathbf{D}_{t}\textbf{{u}}_{\mathrm{h}}+\rho_{0}\mathbf{M}\textbf{{u}}_{\mathrm{h}}=\operatorname{\mathbf{grad}_{\mathrm{h}}}p^{\prime}+\rho_{0}\textbf{{f}}_{\!\mathrm{h}},$ (8a)
 $\operatorname{div_{\mathrm{h}}}\textbf{{u}}_{\mathrm{h}}=0.$ (8b)
-employing the fact $\operatorname{\mathbf{grad}}(\rho_{0}(\operatorname{\mathbf{grad}}\textbf{{u}}_{\mathrm{h}})^{\top})=\operatorname{\mathbf{grad}_{\mathrm{h}}}(\rho_{0}\operatorname{div_{\mathrm{h}}}\textbf{{u}}_{\mathrm{h}})=0$ due to (5b).
-Physically, the assumption of small radial motions is motivated by an identification of the observed inertial modes from [15] as the eigenvalues of the 3D linearized equations of momentum, mass and energy [5].
-In this setup, the radial velocities of the eigenfunctions are small compared to the horizontal ones for most of the observed modes.
-This is also the case in nonlinear simulations of rotating convection in a spherical shell [6].
-This assumption could possibly be justify mathematically within an asymptotic analysis as $r\to\infty$ or $\Omega\to 0$.
-This, however, would require to amend the system (3) by an atmospheric model and assumptions on the coefficient functions and is beyond the scope of this paper.
+with the right hand side $\textbf{{f}}_{\!\mathrm{h}}:=\mathbf{R}\mathbf{R}^{\top}\textbf{{f}}$.
+Here, we employ the fact $\operatorname{\mathbf{grad}}(\rho_{0}(\operatorname{\mathbf{grad}}\textbf{{u}}_{\mathrm{h}})^{\top})=\operatorname{\mathbf{grad}_{\mathrm{h}}}(\rho_{0}\operatorname{div_{\mathrm{h}}}\textbf{{u}}_{\mathrm{h}})=0$ due to (8b), with operators $\bm{\Delta}_{\mathrm{h}}$, $\operatorname{\mathbf{grad}_{\mathrm{h}}}$, $\operatorname{div_{\mathrm{h}}}$ defined in (51)-(52) in Appendix A.
+We note that all differential operators with subscript $\rm h$} denote operators on spheres and are defined in Appendix A.
+Physically, the assumption of small radial motions is motivated by an identification made in [GizonInertial21], namely, that the observed inertial modes can be viewed as eigenvalues of the 3D linearized equations of momentum, mass and energy.
+In this setting, the radial velocities of the eigenfunctions are small relative to the horizontal components for most observed modes.
+This observation is also supported by nonlinear simulations of rotating convection in a spherical shell [Bekki2022].
+While this assumption could potentially be justified through asymptotic analysis ($r\to\infty$ or $\Omega\to 0$), such a rigorous treatment would require amending the system (6) with an atmospheric model and explicit assumptions on the coefficient functions – a task beyond this paper’s scope.
 Reformulation as scalar fourth-order equation.
 It follows from the Helmholtz decomposition (56) and Eq. (8b) that there exits a stream function $\Psi$ such that
-$\rho_{0}\textbf{{u}}_{\mathrm{h}}=\operatorname{\mathbf{curl}_{\mathrm{h}}}\Psi.$
+$\rho_{0}\textbf{{u}}_{\mathrm{h}}=\operatorname{\mathbf{curl}_{\mathrm{h}}}(\Psi)$
+with $\operatorname{\mathbf{curl}_{\mathrm{h}}}$ as in (51)-(52).
 In order to obtain an equation for $\Psi$, we apply $\operatorname{curl_{\mathrm{h}}}$ to (8a).
@@
 $\operatorname{curl_{\mathrm{h}}}\rho_{0}\gamma\bm{\Delta}_{\mathrm{h}}\textbf{{u}}_{\mathrm{h}}=-\gamma(\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}})^{2}\Psi=-\gamma\Delta_{\mathrm{h}}^{2}\Psi$ (9)
-$\operatorname{curl_{\mathrm{h}}}\mathbf{M}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}+\frac{1}{r^{2}}\frac{\partial\Omega}{\partial\theta}\frac{\partial^{2}\Psi}{\partial\theta\,\partial\phi},\quad\alpha_{\Omega}(\theta):=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\right)$
+$\operatorname{curl_{\mathrm{h}}}\mathbf{M}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}+\frac{1}{r^{2}}\frac{\partial\Omega}{\partial\theta}\frac{\partial^{2}\Psi}{\partial\theta\,\partial\phi},\qquad\alpha_{\Omega}(\theta):=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\right)$
 with the analog $D_{t}:=\frac{\partial}{\partial t}+(\Omega-\Omega_{\mathrm{{ref}}})\frac{\partial}{\partial\phi}$ of $\mathbf{D}_{t}$, we arrive at the scalar fourth-order equation
@@
 with $f:=\rho_{0}\operatorname{curl_{\mathrm{h}}}\textbf{{f}}_{\!\mathrm{h}}$, and $\alpha_{\Omega}$ is a linear function of $\Omega$ as in (2.1) for convenience.
-If we are looking for time-harmonic solutions $\Psi(t,\textbf{{r}})=\Re\left(e^{i\omega t}\Psi_{\omega}(\textbf{{r}})\right)$, then the space-dependent part has to satisfy the fourth-order elliptic equation
+Lastly, if we are looking for time-harmonic solutions $\Psi(t,\textbf{{r}})=\Re\left(e^{i\omega t}\Psi_{\omega}(\textbf{{r}})\right)$, then the space-dependent part has to satisfy the fourth-order elliptic equation
 $\gamma\Delta_{\mathrm{h}}^{2}\Psi+i\omega\Delta_{\mathrm{h}}\Psi-\beta_{\Omega}\Delta_{\mathrm{h}}\frac{\partial\Psi}{\partial\phi}+\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}=f$ (11)
-with $\beta_{\Omega}(\theta):=\Omega(r,\theta)-\Omega_{\rm ref}$.
+with $\beta_{\Omega}(\theta):=\Omega(r,\theta)-\Omega_{\rm ref}$. The resulting equation (11) serves as the underlying model of the forward problem studied in Section 3
 2.2 Separated equation as inversion model
+To prepare for the inversion analysis, we now derive the separated version of (11) that will serve as our model for parameter identification.
 Expanding $\Psi_{\omega}(\theta,\phi)=\frac{1}{\sqrt{2\pi}}\sum_{m=-\infty}^{\infty}\widehat{\Psi}_{\omega,m}(\theta)e^{im\phi}$ into a Fourier series in azimuth, the coefficients $\widehat{\Psi}_{\omega,m}$ satisfy the ordinary boundary value problems
-$\begin{cases}\gamma\Delta_{m}^{2}\widehat{\Psi}_{\omega,m}(\theta)+i\omega\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)-im\beta_{\Omega}(\theta)\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)+im\alpha_{\Omega}(\theta)\widehat{\Psi}_{\omega,m}(\theta)=\hat{f}_{m}(\theta)\\ (\Gamma_{m}\widehat{\Psi}_{\omega,m})(\theta)=0,\qquad\theta\in P:=\{0,\pi\}\end{cases}$ (9)
+$\begin{cases}\gamma\Delta_{m}^{2}\widehat{\Psi}_{\omega,m}(\theta)+i\omega\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)-im\beta_{\Omega}(\theta)\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)+im\alpha_{\Omega}(\theta)\widehat{\Psi}_{\omega,m}(\theta)=\hat{f}_{m}(\theta)\\ (\Gamma_{m}\widehat{\Psi}_{\omega,m})(\theta)=0,\hskip 18.49988pt\theta\in P:=\{0,\pi\}\end{cases}$ (12)
 with $\Delta_{m}$ defined by (61) in Appendix A, and the following boundary value operators derived in Appendix B
-$\Gamma_{0}\Psi:=\begin{pmatrix}\Psi^{\prime}|_{P}\\ \Psi^{\prime\prime\prime}|_{P}\end{pmatrix},\quad\Gamma_{\pm 1}\Psi):=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime\prime}|_{P}\end{pmatrix},\quad\Gamma_{m}\Psi:=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime}|_{P}\end{pmatrix},\;|m|\geq 2.$ (10)
+$\Gamma_{0}\Psi:=\begin{pmatrix}\Psi^{\prime}|_{P}\\ \Psi^{\prime\prime\prime}|_{P}\end{pmatrix},\qquad\Gamma_{\pm 1}\Psi:=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime\prime}|_{P}\end{pmatrix},\qquad\Gamma_{m}\Psi:=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime}|_{P}\end{pmatrix},\;|m|\geq 2.$ (13)
 We remark that this equation reduces to Eq.
-(2.9) in [40] if $\gamma=0$ and to Eq.
-(8) in [12] if $\hat{f}_{m}=0$.
-However, vanishing Cauchy data are imposed at the poles in [12], which for $|m|\leq 1$ may yield solutions that do not correspond to the unseparated equation (8).
-The system (9)-(10) is the underlying model for the parameter inversion problem considered in Section 4.
-Solving this inverse problem requires measured data, which is the newly discovered inertial waves in this study.
-The acquisition process will be detailed in the forthcoming section.
-2.3 Observation of solar oscillations
-The basic input of helioseismology are measurements of line-of-sight velocity based on Doppler shifts of certain spectral lines.
-These so-called Dopplergrams acquired since 1996 by ground- or space-based telescopes have high-spatial resolution, of at least $1000\times 1000$ pixels, and a temporal cadence of at least $1$ minute.
-A Fourier transform of the Dopplergram data $v_{\rm dop}(\theta,\phi,t)$ in longitude $\phi$ and time $t$ yields data $v_{\rm dop}^{m,\omega}(\theta)$.
-Due to the limited data coverage, as only at most half of the Sun is visible at a given time, the transform in $\phi$ can only be integrated on $0,\pi]$ instead of $0,2\pi]$, leading to a leakage of the signal; see, e.g.
-the review in [4].
-Moreover an operator should be applied to convert the simulated $\hat{\Psi}$ into $v_{\rm dop}$ using for example that $v_{\rm dop}=\mathbf{l}\cdot\mathbf{u}$, where $\mathbf{l}$ is the line-of-sight vector.
-On a formal point of view, this could be written as a (linear) operator that transforms $\Psi$ onto $v_{\rm dop}$.
-We will not consider this aspect in this article, simply assuming that reprocessed signal $y^{\delta}$ is available and described by
-$y^{\delta}(\theta)=\Psi(\theta)+\mathrm{noise}(\theta)\qquad\theta\in I:=(0,\pi)\quad\text{or}\quad\theta\in I^{\prime}:=(\epsilon,\pi-\epsilon),\epsilon\geq 0,$ (11)
-here $\delta=\|\mathrm{noise}\|$ denotes the noise level.
-The case $\epsilon=0$ corresponds to full surface coverage, while $\epsilon>0$ takes into account that higher latitudes cannot easily be observed due to the line-of-sight projection.
-It is usually not possible to work with the Doppler signal directly due to the high level of convective noise.
-However, some inertial modes of large amplitudes have been directly detected [14], enabling us to consider the recovery of viscosity and rotation from the measurements of $y^{\delta}$.
+(2.9) in [Watson81] if $\gamma=0$ and to Eq.
+(8) in [Damien22] if $\hat{f}_{m}=0$.
+However, vanishing Cauchy data are imposed at the poles in [Damien22], which for $|m|\leq 1$ may yield solutions that do not correspond to the unseparated equation (11).
+The system (12)–(13) provides the forward model for the parameter inversion problem addressed in Section 4.
+Solving this inverse problem requires inertial wave observations.
+The data acquisition process is described in the following section.
+2.3 Observation of solar inertial modes
+Solar physicists routinely measure the horizontal flow field $\mathbf{u}_{h}(\textbf{{r}},t)$ at the solar surface, in the rotating frame $\Omega_{\rm ref}$ at intermediate spatial scales with a typical temporal cadence of one day [GizonInertial21].
+These measurements are obtained using local correlation tracking (LCT), which infers large-scale flows from the motion of small-scale convective structures (granules) observed in intensity images.
+High-resolution imaging from instruments such as HMI on the SDO spacecraft is required to resolve these granules.
+The stream function can be computed from $\Psi=-\Delta_{h}^{-1}(\operatorname{curl_{\mathrm{h}}}(\rho_{0}\mathbf{u}_{h})),$ then Fourier transformed in time and longitude to obtain $\Psi^{m,\omega}(\theta)$, which provides data for the inverse problems.
+Due to the limited observation coverage (only half of the Sun is visible at any time), the azimuthal Fourier transform integrates over $0,\pi]$ instead of $0,2\pi]$, causing leakage of the signal between the harmonic degrees $m$; see [Basu2016] for a review.
+We will not concern ourselves with this aspect, but instead focus on the observational limitation: higher latitudes cannot easily be observed due to the line-of-sight projection.
+Thus, we assume to observe
+$y^{\delta}(\theta):=\Psi^{m,\omega}(\theta)+\mathrm{noise}(\theta),\qquad\theta\in I:=(0,\pi)\qquad\text{or}\qquad\theta\in I^{\prime}:=(\epsilon,\pi-\epsilon),\epsilon\geq 0,$ (14)
+where $\delta=\|\mathrm{noise}\|$ denotes the noise level. The case $\epsilon=0$ corresponds to full surface coverage, while $\epsilon>0$ takes into account the missing data at higher latitudes.
 3 Inertial waves on the manifold $r{\mathbb{S}^{2}}$}
-We first show the well-posedness of Eq. (8) associated to the differential operator
+With the model established, we now address the forward problem by proving existence, uniqueness and stability of wave solutions, for both the full equation (11) and its separated counterpart (12).
+We first show the well-posedness of Eq.
+(11) associated to the differential operator
 $\mathbf{B}(\Psi):=\gamma\Delta_{\mathrm{h}}^{2}\Psi+i\omega\Delta_{\mathrm{h}}\Psi-\beta_{\Omega}\Delta_{\mathrm{h}}\frac{\partial\Psi}{\partial\phi}+\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}$ (15)
 with recalling that $\beta_{\Omega}(\theta)=\Omega(\theta)-\Omega_{\mathrm{ref}}$ and $\alpha_{\Omega}(\theta)=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\big(\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\big)$.
-We will work in the space $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ of functions in $H^{2}({r{\mathbb{S}^{2}}})$ with zero mean introduced in the Appendix A.
-By partial integration with taking into account symmetry of the bi-Laplacian and Laplacian operators (see Appendix, Lemma 19), the corresponding sesquilinear form $\mathbf{A}:H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\times H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\to\mathbb{C}$ defined by $\mathbf{A}[\Psi,\psi]:=\langle\mathbf{B}(\Psi),\psi\rangle_{H_{\diamond}^{-2},H_{\diamond}^{2}}$ is given by
-$\begin{aligned} \mathbf{A}[\Psi,\psi]&=\gamma\left\langle\Delta_{\mathrm{h}}\Psi,\Delta_{\mathrm{h}}\psi\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\Psi,\operatorname{\mathbf{grad}_{\mathrm{h}}}\psi\right\rangle+(\mathbf{A}^{\alpha}+\mathbf{A}^{\beta})[\Psi,\psi],\\ &\mathbf{A}^{\alpha}[\Psi,\psi]:=-\left\langle\alpha_{\Omega}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle,\qquad\mathbf{A}^{\beta}[\Psi,\psi]:=\left\langle\beta_{\Omega}\Delta_{\mathrm{h}}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle.\end{aligned}$ (13)
+Consider the space $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ of functions in $H^{2}({r{\mathbb{S}^{2}}})$ with zero mean defined in Appendix A.
+By partial integration with taking into account symmetry of the bi-Laplacian and Laplacian operators (see Appendix, Lemma 19), the corresponding sesquilinear form $\mathbf{A}:H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\times H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\to\mathbb{C}$ defined as $\mathbf{A}[\Psi,\psi]:=\langle\mathbf{B}(\Psi),\psi\rangle_{H_{\diamond}^{-2},H_{\diamond}^{2}}$ is given by
+$\begin{aligned} \mathbf{A}[\Psi,\psi]&=\gamma\left\langle\Delta_{\mathrm{h}}\Psi,\Delta_{\mathrm{h}}\psi\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\Psi,\operatorname{\mathbf{grad}_{\mathrm{h}}}\psi\right\rangle+(\mathbf{A}^{\alpha}+\mathbf{A}^{\beta})[\Psi,\psi],\\ &\mathbf{A}^{\alpha}[\Psi,\psi]:=-\left\langle\alpha_{\Omega}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle,\hskip 18.49988pt\mathbf{A}^{\beta}[\Psi,\psi]:=\left\langle\beta_{\Omega}\Delta_{\mathrm{h}}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle.\end{aligned}$ (16)
 Since the coefficients $\alpha_{\Omega}$ and $\beta_{\Omega}$ are independent of $\phi$, this sesquilinear form is block diagonal with respect to the decomposition $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus_{m\in Z}^{\infty}V_{m}$ with the subspaces
@@
 $\mathbf{A}_{m}[\Psi_{m},\psi_{m}]$ (18)
-$\quad=\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi_{m},\operatorname{\mathbf{grad}_{\mathit{m}\!}}\psi_{m}\right\rangle+(\mathbf{A}^{\alpha}_{m}+\mathbf{A}^{\beta}_{m})[\Psi_{m},\psi_{m}],$
-$\mathbf{A}^{\alpha}_{m}[\Psi_{m},\psi_{m}]:=im\left\langle\alpha_{\Omega}\Psi_{m},\psi_{m}\right\rangle,\qquad\;\mathbf{A}^{\beta}_{m}[\Psi_{m},\psi_{m}]:=-im\left\langle\beta_{\Omega}\Delta_{m}\Psi_{m},\psi_{m}\right\rangle.$
+$\qquad=\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi_{m},\operatorname{\mathbf{grad}_{\mathit{m}\!}}\psi_{m}\right\rangle+(\mathbf{A}^{\alpha}_{m}+\mathbf{A}^{\beta}_{m})[\Psi_{m},\psi_{m}],$
+$\mathbf{A}^{\alpha}_{m}[\Psi_{m},\psi_{m}]:=im\left\langle\alpha_{\Omega}\Psi_{m},\psi_{m}\right\rangle,\hskip 18.49988pt\;\mathbf{A}^{\beta}_{m}[\Psi_{m},\psi_{m}]:=-im\left\langle\beta_{\Omega}\Delta_{m}\Psi_{m},\psi_{m}\right\rangle.$
 with $\mathbf{A}^{\alpha}_{m},\mathbf{A}^{\beta}_{m}:V_{m}\times V_{m}\to\mathbb{C}$ as restrictions of $\mathbf{A}^{\alpha}$, $\mathbf{A}^{\beta}$ defined in (16).
-Here, $\Delta_{m}$, $\operatorname{\mathbf{grad}_{\mathit{m}\!}}$ are the decomposed operators as defined in (58), (59) in Appendix A.
+Here, $\Delta_{m}$, $\operatorname{\mathbf{grad}_{\mathit{m}\!}}$ are the decomposed operators as defined in (61), (62).
 3.1 Boundedness
-It is straightforward to show that $\mathbf{A}^{\alpha}$ and $\mathbf{A}^{\alpha}_{m}$ are bounded if $\Omega\in W^{2,\infty}(0,\pi)$.
-However, we wish to reduce the regularity assumption on the unknown rotation and only assume
+It is straightforward to verify that $\mathbf{A}^{\alpha}$ and $\mathbf{A}^{\alpha}_{m}$ are bounded if $\Omega\in W^{2,\infty}(0,\pi)$.
+However, to reduce the regularity assumption on the unknown rotation, we only constrain
 $\Omega\in H^{1}({r{\mathbb{S}^{2}}};m=0)\,.$ (19)
 Under this reduced regularity assumption, establishing boundedness requires additional technical analysis.
-For later use in $\mathrm{G\mathring{a}rding}$ inequality, we also provide a bound by a weaker norm.
 In the following, we will frequently employ continuous embeddings between function spaces $X,Y$ denoted the embedding constants $C_{X\to Y}$, i.e., $\|\cdot\|_{Y}\leq C_{X\to Y}\|\cdot\|_{X}$.
@@
 $\geq\frac{3\gamma}{4}\|\Psi\|^{2}_{H_{\diamond}^{2}}+|\omega|\|\Psi\|^{2}_{H_{\diamond}^{1}}-\frac{2}{\gamma}(C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}})^{2}\big(\|\Omega-\Omega_{\mathrm{{ref}}}\|^{2}_{H_{\diamond}^{1}}+9\|\Omega\|^{2}_{H_{\diamond}^{1}}\big)\|\Psi\|^{2}_{H_{\diamond}^{3/2}}.$
-Since the embeddings $H^{1}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ and $H^{3/2}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ are compact, $\mathbf{B}$ is a Fredholm operator of index 0; see, e.g., [33, Sec.
+Since the embeddings $H^{1}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ and $H^{3/2}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ are compact, $\mathbf{B}$ is a Fredholm operator of index 0; see, e.g., [Renardy1992, Sec.
 8.2.4].
@@
 $\sqrt{2}|\mathbf{A}[\Psi,\Psi]|$
-$\quad\geq\frac{\gamma}{2}\|\Psi\|^{2}_{H_{\diamond}^{2}}+\left(|\omega|\!-\!\frac{4}{\gamma^{3}}(C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}})^{4}\big{(}\|\Omega\!-\!\Omega_{\mathrm{{ref}}}\|^{2}_{H^{1}}\!+\!9\|\Omega\|^{2}_{H^{1}}\big{)}^{2}\right)\|\Psi\|^{2}_{H_{\diamond}^{1}}.$
-If $|\omega|$ satisfies (23), then $\mathbf{A}$ is coercive, so $\mathbf{B}$ is boundedly invertible by the Lax-Milgram theorem [11, Section 6.2, Theorem 1]. ∎
+$\qquad\geq\frac{\gamma}{2}\|\Psi\|^{2}_{H_{\diamond}^{2}}+\left(|\omega|\!-\!\frac{4}{\gamma^{3}}(C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}})^{4}\big(\|\Omega\!-\!\Omega_{\mathrm{{ref}}}\|^{2}_{H^{1}}\!+\!9\|\Omega\|^{2}_{H^{1}}\big)^{2}\right)\|\Psi\|^{2}_{H_{\diamond}^{1}}.$
+If $|\omega|$ satisfies (26), then $\mathbf{A}$ is coercive, so $\mathbf{B}$ is boundedly invertible by the Lax-Milgram theorem [Evans, Section 6.2, Theorem 1]. ∎
 We now indicate the dependence of $\mathbf{B}$ on the frequency $\omega$ explicitely by $\mathbf{B}_{\omega}$ and also consider complex values of $\omega$.
@@
 There exists a discrete set $\Lambda\subset\mathbb{C}$ without accumulation points such that $\mathbf{B}_{\omega}$ is boundedly invertible in $L(H_{\diamond}^{2}({r{\mathbb{S}^{2}}}),H_{\diamond}^{-2}({r{\mathbb{S}^{2}}}))$ for $\omega\in\mathbb{C}\setminus\Lambda$, and $\dim\operatorname{ker}\mathbf{B}_{\omega}\in\mathbb{N}$ for $\omega\in\Lambda$.
-The elements of $\operatorname{ker}\mathbf{B}_{\omega}$ for $\omega\in\Lambda$ are called inertial modes and have been studied intensively in [12] and other publications.
+The elements of $\operatorname{ker}\mathbf{B}_{\omega}$ for $\omega\in\Lambda$ are called inertial modes and have been studied intensively in [Damien22] and other publications.
 Proof of Corollary 4.
@@
 Moreover, $\omega\mapsto\mathbf{C}(\omega)$ is affinely linear and in particular holomorphic.
-Therefore, we are in the framework of the analytic Fredholm theorem (see [33, Thm.
+Therefore, we are in the framework of the analytic Fredholm theorem (see [Renardy1992, Thm.
 7.92]).
@@
 Let us summarize some consequences for the separated operators $\mathbf{B}_{m}:V_{m}\to V_{m}^{*}=H^{-2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ defined by
-$\left\langle\mathbf{B}_{m}\Psi_{m},\psi_{m}\right\rangle=\mathbf{A}_{m}[\Psi_{m},\psi_{m}],\qquad\Psi_{m},\psi_{m}\in V_{m}.$ (24)
+$\left\langle\mathbf{B}_{m}\Psi_{m},\psi_{m}\right\rangle=\mathbf{A}_{m}[\Psi_{m},\psi_{m}],\hskip 18.49988pt\Psi_{m},\psi_{m}\in V_{m}.$ (27)
 Corollary 5.
@@
 $\Re\big(im\big\langle\beta_{\Omega}\Delta_{m}\Psi,\Psi\big\rangle\big)$
-$\quad\leq|m|\left|\Im\big{\langle}\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega}\Psi)\big{\rangle}\right|=|m|\left|\Im\big{\langle}\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega})\Psi\big{\rangle}\right|$
-$\quad\leq|m|\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi\|_{L^{6}}\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\beta_{\Omega}\|_{L^{2}}\|\Psi\|_{L^{3}}\leq\frac{|m|}{r}C_{H^{2}_{\diamond}\to L^{3}_{\diamond}}C_{H^{1}_{\diamond}\to L^{6}_{\diamond}}\|\Omega^{\prime}\|_{L^{2}({r{\mathbb{S}^{2}}})}\|\Psi\|_{H^{2}_{\diamond}}^{2}.$
+$\qquad\leq|m|\left|\Im\big\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega}\Psi)\big\rangle\right|=|m|\left|\Im\big\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega})\Psi\big\rangle\right|$
+$\qquad\leq|m|\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi\|_{L^{6}}\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\beta_{\Omega}\|_{L^{2}}\|\Psi\|_{L^{3}}\leq\frac{|m|}{r}C_{H^{2}_{\diamond}\to L^{3}_{\diamond}}C_{H^{1}_{\diamond}\to L^{6}_{\diamond}}\|\Omega^{\prime}\|_{L^{2}({r{\mathbb{S}^{2}}})}\|\Psi\|_{H^{2}_{\diamond}}^{2}.$
 Therefore, $\Re\mathbf{A}_{m}[\Psi,\Psi]\geq c\|\Delta_{m}\Psi\|^{2}_{L^{2}({r{\mathbb{S}^{2}}})}$ with $c:=\gamma-\frac{|m|}{r}C_{H^{2}_{\diamond}\to L^{3}_{\diamond}}C_{H^{1}_{\diamond}\to L^{6}_{\diamond}}\|\Omega^{\prime}\|_{L^{2}}>0$ ensured by the assumption (28).
@@
 3.4 Additional regularity
+We have established unique existence of wave solutions.
+However, we proceed further with proving higher regularity of the wave solutions, particularly for the separated equations.
+This regularity result is not merely a technical refinement; it is essential for convergence guarantee of iterative regularization methods later in Section 5.
 Proposition 8 (Lifted regularity).
@@
 Further assume that source and rotation have regularity
-$f_{m}\in L^{2}({r{\mathbb{S}^{2}}};m),\qquad\Omega\in H^{2}({r{\mathbb{S}^{2}}};m=0).$ (26)
+$f_{m}\in L^{2}({r{\mathbb{S}^{2}}};m),\hskip 18.49988pt\Omega\in H^{2}({r{\mathbb{S}^{2}}};m=0).$ (29)
 Then the solution $\Psi_{m}:=\mathbf{B}_{m}^{-1}f_{m}$ as well as $\tilde{\Psi}_{m}:=(\mathbf{B}_{m}^{\star})^{-1}f_{m}$ with the adjoint operator $\mathbf{B}_{m}^{\star}:H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)\to H^{-2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ have higher regularity
@@
 The solution $\Psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ satisfies
-$\begin{aligned} \forall\psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m):\quad&\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle=\left\langle\tilde{f}_{m},\psi_{m}\right\rangle\\ &\mbox{with }\tilde{f}_{m}:=f_{m}+i\omega\Delta_{m}\Psi_{m}-im\beta_{\Omega}\Delta_{m}\Psi_{m}+im\alpha_{\Omega}\Psi_{m}.\end{aligned}$ (28)
+$\begin{aligned} \forall\psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m):\qquad&\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle=\left\langle\tilde{f}_{m},\psi_{m}\right\rangle\\ &\mbox{with }\tilde{f}_{m}:=f_{m}+i\omega\Delta_{m}\Psi_{m}-im\beta_{\Omega}\Delta_{m}\Psi_{m}+im\alpha_{\Omega}\Psi_{m}.\end{aligned}$ (31)
 We shall claim that $\tilde{f}_{m}\in L^{2}({r{\mathbb{S}^{2}}};m)$.
@@
 The proof of the analogous statement for $\mathbf{B}_{m}^{\star}$ only requires a different treatment of the term involving $\beta_{\Omega}$.
-Here, we use the fact that $H^{2}({r{\mathbb{S}^{2}}})$ is a Banach algebra (see [2, p.
-115] or [3, Thm 1.4]) to show that $\|\beta_{\Omega}\Psi_{m}\|_{H^{2}}\leq C\|\beta_{\Omega}\|_{H^{2}}\|\Psi_{m}\|_{H^{2}}$.
+Here, we use the fact that $H^{2}({r{\mathbb{S}^{2}}})$ is a Banach algebra (see [adams:75, p.
+115] or [BadrBernicotRuss, Thm 1.4]) to show that $\|\beta_{\Omega}\Psi_{m}\|_{H^{2}}\leq C\|\beta_{\Omega}\|_{H^{2}}\|\Psi_{m}\|_{H^{2}}$.
 Concerning the boundary values, we use the continuous embedding $H^{4}({r{\mathbb{S}^{2}}})\hookrightarrow C^{2}({r{\mathbb{S}^{2}}})$ in (58c) and Lemma 18 to show that $\Gamma_{m}\hat{\Psi}_{m}$ is well-defined and vanishes.
@@
 3.5 Well-posedness on the restricted domain
-Inspired by the practical measurement scenario in Section 2.3, where only data on the restricted latitude range $I^{\prime}=(\epsilon,\pi-\epsilon)$ can be observed, this section derives well-posedness, regularity of the wave operators on this restricted domain.
-This result will be later employed in Section 5.2 for convergence guarantee of reconstruction method under restricted measurements.
-We sketch how an analysis similar to that of Propositions 7 and 8 can be carried out, whose key elements are symmetry of the bi-Laplacian and Laplacian operators and partial integrations.
+Motivated by the practical measurement scenario in Section 2.3, where observations are limited to a restricted latitude range $I^{\prime}=(\epsilon,\pi-\epsilon)$, this section derives well-posedness and regularity of the wave operators on this restricted domain.
+These results are subsequently employed in Section 5.2 for convergence analysis of reconstruction methods under restricted observations.
+We outline how an analysis analogous to that of Propositions 7 and 8 can be conducted.
+The key elements are symmetry of the bi-Laplacian and Laplacian operators and strategic use of partial integrations.
 According to Remark 20, symmetry of Laplacian operators on function spaces supported on $I^{\prime}$ is guaranteed with homogeneous Cauchy boundary data, i.e.
@@
 Given $\gamma>0$, and $\Omega\in H^{1}({r{\mathbb{S}^{2}}};m=0)$ satisfying (26) or (28). Assume for any $m\in\mathbb{Z}$ the Cauchy boundary condition
-$\Psi|_{\partial I^{\prime}}=u_{1}\in\mathbb{C}^{2}\quad\text{and}\quad\Psi^{\prime}|_{\partial I^{\prime}}=u_{2}\in\mathbb{C}^{2}.$ (31)
+$\Psi|_{\partial I^{\prime}}=u_{1}\in\mathbb{C}^{2}\qquad\text{and}\qquad\Psi^{\prime}|_{\partial I^{\prime}}=u_{2}\in\mathbb{C}^{2}.$ (34)
 Let $u\in C^{\infty}(I)$ with $u|_{\partial I^{\prime}}=u_{1}$, $u^{\prime}|_{\partial I^{\prime}}=u_{2}$ and $g:=\mathbf{B}_{m}u$. We claim:
@@
 Using the fact the weak solutions of the bi-Laplace equations are also strong solutions, one obtains boundedness of the fourth order term $\|\Delta_{m}^{2}\Psi_{m}\|_{L^{2}}$, given higher regularity of $\Omega$ and $f$ as in Proposition 8.
-Alternatively, one can apply general elliptic regularity results as, e.g., in [38, Sec.
+Alternatively, one can apply general elliptic regularity results as, e.g., in [Triebel, Sec.
 5.3.4].
 ∎
-4 Inverse problems for viscosity and rotation
+4 Inverse problem for viscosity and rotation
+With the forward problems established, we now address the inverse problem: recovering the differential rotation function $\Omega$ and the viscosity parameter $\gamma$ in the separated equations under various data measurement strategies.
 4.1 Inversion with different observation strategies
-In this section, we address the inverse problem of recovering the differential rotation function $\Omega$ and the viscosity coefficient $\gamma$ in the separated equations.
-To achieve this, we require additional information, namely full or partial measurements of the state, as described in Section 2.3:
-$\begin{array}[]{l}L:V_{m}\to Y\\ (L\Psi)(\theta):=\Psi(\theta)\end{array}\qquad Y=\begin{cases}L^{2}(I,r^{2}\sin),&\text{full measurements}\\ L^{2}(I^{\prime},r^{2}\sin),&\text{partial measurements}\end{cases}$ (34)
-with the weighted $L^{2}$-data spaces. We formulate the parameter identification problem within the Hilbert space framework established by the well-posedness results in Sections 3.2-3.3:
-$\gamma\in\mathbb{R}^{+},\quad\Omega\in X:=H^{1}({r{\mathbb{S}^{2}}},m=0),\quad\Psi\in V_{m}:=H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m),\quad y\in Y,$ (35)
-By introducing the parameter-to-state map
+As described in Section 2.3, data is acquired under two scenarios: full or partial measurements of the wave solution. The measurement operator is hence defined as
+$\begin{array}[]{l}L:V_{m}\to Y\\ (L\Psi)(\theta):=\Psi(\theta)\end{array}\hskip 18.49988ptY=\begin{cases}L^{2}(I,r^{2}\sin),&\text{full measurements}\\ L^{2}(I^{\prime},r^{2}\sin),&\text{partial measurements}\end{cases}$ (37)
+where $Y$ are weighted $L^{2}$-data spaces. Within the Hilbert space framework established by the well-posedness results in Sections 3.23.3, that is
+$\gamma\in\mathbb{R}^{+},\qquad\Omega\in X:=H^{1}({r{\mathbb{S}^{2}}},m=0),\qquad\Psi\in V_{m}:=H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m),\qquad y\in Y,$ (38)
+we introduce the parameter-to-state map
 […]
-together with the measurement operator we can formulate the forward operator $F$ for the parameter identification as
-$F:\mathcal{D}(S)\to Y\qquad F:=L\circ S.$ (37)
+Together with the measurement operator $L$ in (37), the forward operator $F$ for the parameter identification is formulated as
+$F:\mathcal{D}(S)\to Y\hskip 18.49988ptF:=L\circ S.$ (40)
 Note that $X$ is the space of real-valued functions, while $V_{m}$ and $Y$ are spaces of complex-valued functions; together, they form Gelfand triples.
+Compactness of the embedding $V_{m}\hookrightarrow Y$ yields compactness of the measurement operator $L$, thus of the forward operator $F$.
+In such cases, the inverse parameter problem becomes ill-posed, necessitating regularization to ensure stable reconstruction of the target coefficients.
 4.2 Sensitivity and adjoints
-Compactness of the embedding $V_{m}\hookrightarrow L^{2}(I,r^{2}\sin)$ yields compactness of the measurement operator $L$, thus of the forward operator $F$.
-In such cases, the inverse coefficient problem becomes ill-posed, necessitating regularization for stable reconstruction.
-Given the inherent nonlinearity of the problem, iterative methods such as Landweber-type and Newton-type algorithms are typically employed for reconstruction (see Section 5).
-For noisy data $y^{\delta}\in Y$, the former is given by
-$p_{k+1}=p_{k}-F^{\prime}(p_{k})^{*}(F(p_{k})-y^{\delta})\qquad k\leq K(y^{\delta},\delta),$
-where the unknown parameter is $p:=(\gamma,\Omega)$.
+Given the inherent nonlinearity of the inverse problem, iterative regularization methods such as Landweber-type and Newton-type algorithms are typically employed (see Section 5).
+For noisy data $y^{\delta}\in Y$, the Landweber iteration is given as
+$p_{k+1}=p_{k}-F^{\prime}(p_{k})^{*}(F(p_{k})-y^{\delta})\hskip 18.49988ptk\leq K(y^{\delta},\delta),$
+for the unknown parameter $p:=(\gamma,\Omega)$.
 The stopping index $K(y^{\delta},\delta)$ serves as the regularization parameter and is determined according to the discrepancy principle; see Section 6 for further details.
-This formulation highlights that the main components of iterative gradient-based methods: sensitivity analysis and the derivation of the adjoint; these are the primary focus of this section.
-To this end, we will add the dependence on the unknown parameters $p=(\gamma,\Omega)\in\mathcal{D}(S)$ in parenthesis, i.e.
-$\mathbf{B}_{m}(p)=\mathbf{B}_{m}:V_{m}\to V_{m}^{*}.$ Since the mapping $p\mapsto\mathbf{B}_{m}(p)$ is affine linear, its derivative with respect to $p$ is independent of $p$ and we can omit it in our notation.
-Hence, we simply denote the derivative of $p\mapsto\mathbf{B}_{m}(p)$ by
-$\mathbf{B}_{m}^{\prime}:\mathbb{R}\times X\to\mathcal{L}(V_{m},V_{m}^{*}),\qquad\mathbf{B}_{m}^{\prime}(\delta\gamma,\delta\Omega)\Psi:=(\delta\gamma)\Delta_{m}^{2}\Psi+im\alpha_{\delta\Omega}\Psi-im(\delta\Omega)\Psi.$
+This formulation highlights the main components of gradient-based regularization schemes: sensitivity analysis and adjoint derivation; these are the main focus of this section.
+For clarity, we will explicitly denote the dependence of $\mathbf{B}_{m}$ on the unknown parameters $p=(\gamma,\Omega)\in\mathcal{D}(S)$, writing $\mathbf{B}_{m}(p):V_{m}\to V_{m}^{*}.$ As the mapping $p\mapsto\mathbf{B}_{m}(p)$ is affine linear, the derivative is independent of $p$; we thus suppress this argument in the derivative and write
+$\mathbf{B}_{m}^{\prime}:\mathbb{R}\times X\to\mathcal{L}(V_{m},V_{m}^{*}),\hskip 18.49988pt\mathbf{B}_{m}^{\prime}(\delta\gamma,\delta\Omega)\Psi:=(\delta\gamma)\Delta_{m}^{2}\Psi+im\alpha_{\delta\Omega}\Psi-im(\delta\Omega)\Psi.$
 More precisely, $\mathbf{B}_{m}^{\prime}(\delta\gamma,\delta\Omega)$ is the operator induced by the sesquilinear form $\mathbf{A}^{\prime}_{m}(\delta\gamma,\delta\Omega):V_{m}\times V_{m}\to\mathbb{C}$ defined by
@@
 Proof.
-The result follows from the differentiability statement in the implicit function theorem (see, e.g., [41, §4.7]) applied to the operator $G:\mathcal{D}(S)\times V_{m}\to V_{m}^{*}$, $G(p,\Psi):=\mathbf{B}_{m}(p)\Psi-f$.
+The result follows from the differentiability statement in the implicit function theorem (see, e.g., [zeidler1986, §4.7]) applied to the operator $G:\mathcal{D}(S)\times V_{m}\to V_{m}^{*}$, $G(p,\Psi):=\mathbf{B}_{m}(p)\Psi-f$.
 Here, one interprets $S$ as the implicitly defined function $G(p,S(p))=0$ and employs linearity of $G$ in the second argument as well as bounded invertibility of $\mathbf{B}_{m}(p):V_{m}\to V_{m}^{*}$.
 ∎
-Differentiability of $F$ is straightforward from that of $S$ and boundedness of the linear measurement operator $L$ defined in (34). We now derive the adjoint for different observation senarios.
+Differentiability of $F$ is straightforward from that of $S$ and boundedness of the linear measurement operator $L$. We now derive the adjoint for different observations.
 Proposition 11 (Adjoint – full observation).
@@
 The Hilbert space adjoint of the derivative of $F$ in (40) is given by
-$F^{\prime}[p]^{*}:Y\to\mathbb{R}\times X\qquad F^{\prime}[p]^{*}y=-\left[\begin{smallmatrix}1&0\\ 0&I_{X}\end{smallmatrix}\right]\,\Re\left([\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}(\mathbf{B}_{m}(p)^{-1})^{\star}y\right)$ (39)
+$F^{\prime}[p]^{*}:Y\to\mathbb{R}\times X\hskip 18.49988ptF^{\prime}[p]^{*}y=-\left[\begin{smallmatrix}1&0\\ 0&I_{X}\end{smallmatrix}\right]\,\Re\left([\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}(\mathbf{B}_{m}(p)^{-1})^{\star}y\right)$ (42)
 where $\Psi:=\mathbf{B}_{m}(p)^{-1}f$ is again the solution to the separated equation (12) and
-$[\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}:V_{m}\to\mathbb{R}\times X^{*},\quad(\mathbf{B}_{m}(p)^{-1})^{\star}:V_{m}^{*}\to V_{m}$ (40)
+$[\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}:V_{m}\to\mathbb{R}\times X^{*},\qquad(\mathbf{B}_{m}(p)^{-1})^{\star}:V_{m}^{*}\to V_{m}$ (43)
 denotes the Banach space adjoints of $\mathbf{B}_{m}^{\prime}(\cdot)\Psi:\mathbb{R}\times X\to V_{m}^{*}$ and $\mathbf{B}_{m}(p)^{-1}:V_{m}^{*}\to V_{m}$.
@@
 $\left\langle\alpha_{\delta\Omega}\,,\,\overline{\Psi}z\right\rangle$
-$\quad=\int_{0}^{\pi}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\right)\Psi\overline{z}\,d\theta=\int_{0}^{\pi}\frac{-1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\frac{d}{d\theta}\left(\Psi\overline{z}\right)\,d\theta$
-$\quad=\int_{0}^{\pi}(\delta\Omega)(\theta)\sin^{2}\theta\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\Psi\overline{z}\right)\right)\,d\theta=\left\langle\delta\Omega\,,\,\frac{\sin}{r^{2}}\frac{d}{d\theta}\left(\frac{1}{\sin}\frac{d}{d\theta}\left(\overline{\Psi}z\right)\right)\right\rangle$
+$\qquad=\int_{0}^{\pi}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\right)\Psi\overline{z}\,d\theta=\int_{0}^{\pi}\frac{-1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\frac{d}{d\theta}\left(\Psi\overline{z}\right)\,d\theta$
+$\qquad=\int_{0}^{\pi}(\delta\Omega)(\theta)\sin^{2}\theta\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\Psi\overline{z}\right)\right)\,d\theta=\left\langle\delta\Omega\,,\,\frac{\sin}{r^{2}}\frac{d}{d\theta}\left(\frac{1}{\sin}\frac{d}{d\theta}\left(\overline{\Psi}z\right)\right)\right\rangle$
 Inserting these formulas into (42) yields (44). ∎
@@
 Note that the partial observation operator is the restriction operator $R\Psi:=\Psi|_{I^{\prime}}$. Its Hilbert space adjoint is the extension-by-zero operator
-$R^{*}:L^{2}(I^{\prime},r^{2}\sin)\to L^{2}(I,r^{2}\sin)\qquad R^{*}\Psi)(\theta):=\begin{cases}\Psi(\theta),&\theta\in I^{\prime},\\ 0,&\theta\in I\setminus I^{\prime}.\end{cases}$
-The corresponding forward operator relates to the forward operator with full measurement via $F_{\mathrm{part}}[p]=RF_{\mathrm{full}}[p]$.
+$R^{*}:L^{2}(I^{\prime},r^{2}\sin)\to L^{2}(I,r^{2}\sin)\hskip 18.49988ptR^{*}\Psi)(\theta):=\begin{cases}\Psi(\theta),&\theta\in I^{\prime},\\ 0,&\theta\in I\setminus I^{\prime}.\end{cases}$
+The corresponding forward operator relates to that of full measurement via $F_{\mathrm{part}}[p]=RF_{\mathrm{full}}[p]$.
 Thus, $F_{\mathrm{part}}^{\prime}[p]^{*}=F_{\mathrm{full}}^{\prime}[p]^{*}R^{*}$ with $F_{\mathrm{full}}^{\prime}[p]^{*}$ as in Corollary 12.
@@
 5 Tangential cone condition
-As discussed in Section 4, the inverse problem under consideration is ill-posed, necessitating the use of regularization to achieve stable reconstruction of the target coefficients.
-Iterative gradient-based regularization methods, such as Landweber iteration, Newton-type schemes, and their variants, fundamentally rely on three core elements: the well-posedness of the forward operator $F$ (Section 3), adjoint of the linearized operator (Section 4.2), and guarantees of convergence.
-This final requirement is typically ensured by making specific structural assumptions about $F$, particularly regarding its nonlinearity and the uniform boundedness of its derivative; these issues are the main focus of this section.
-To address this, we consider the tangential cone condition (TCC), a celebrated criterion first introduced in [16] that ensures local convergence of iterative regularization algorithms.
+As discussed in Section 4, ill-posedness of our inverse problem necessitates regularization strategies for stable reconstruction.
+Gradient-based regularization methods, including Landweber, Newton-type schemes, and their variants, rely on three core elements: well-posedness of the forward operator $F$ (Section 3), the adjoint of the linearized operator (Section 4.2), and convergence guarantees.
+The third requirement, which forms the primary focus of this section, is ensured by structural assumptions on $F$ – particularly, condition on nonlinearity and uniform boundedness of its derivative.
+We address this through the tangential cone condition (TCC), a celebrated criterion first introduced in [HNS95] that ensures local convergence of iterative regularization algorithms.
 Intuitively, if the forward map $F$ is excessively nonlinear, there is no general guarantee that gradient descent steps will remain within the vicinity of the true solution.
-The tangential cone condition offers a quantitative tool for assessing the nonlinearity of $F$, especially in the context of compact operators, thereby paving the way for the successful application of iterative regularization methods.
-Notably, satisfaction of this condition also implies a characterization of the solution set via null space of the linearized operator, uniqueness of the minimum-norm solution, and, combined with the null space condition, ensures that the iterative reconstruction sequence converges to this solution [22].
-For a general model $F:X\ni p\mapsto y\in Y$, TCC is stated in the following (strong) form, that is the inequality
+The tangential cone condition offers a quantitative tool for assessing nonlinearity of $F$, particularly of compact operators, thereby enabling application of regularization methods to realistic problems.
+Notably, this condition also implies: characterization of the solution set via null space of the linearized operator, uniqueness of minimum-norm solutions, and local convergence of reconstruction sequences towards minimum-norm solutions [KalNeuSch08].
+For a general model $F:X\ni p\mapsto y\in Y$, the TCC states that the following inequality holds:
 $\|F(p)-F(\tilde{p})-F^{\prime}(p)(p-\tilde{p})\|_{Y}\leq C_{tc}\|p-\tilde{p}\|_{X}\|F(p)-F(\tilde{p})\|_{Y}\,\,\,\text{for all }p,\tilde{p}\in B^{X}_{R}(p^{\dagger})$
-holds for some constant $C_{tc}>0$ and radius $R>0$ of the ball around a ground truth $p^{\dagger}\in\mathcal{D}(F)$; see also [25] for the weak forms. In our setting, TCC reads as
+where $C_{tc}>0$ is the TCC constant, and $B^{X}_{R}(p^{\dagger})$ is the ball of radius $R$ around a ground truth $p^{\dagger}\in\mathcal{D}(F)$. In our setting, the TCC reads as
 $\|F(\gamma,\Omega)-F(\tilde{\gamma},\tilde{\Omega})-F^{\prime}[\gamma,\Omega]((\gamma,\Omega)-(\tilde{\gamma},\tilde{\Omega}))\|_{Y}$ (45)
@@
 $\hskip 227.62204pt\text{for all }(\gamma,\Omega),(\tilde{\gamma},\tilde{\Omega})\in B_{R}^{\mathbb{R}\times X}(\gamma^{\dagger},\beta^{\dagger}).$
-It is worth noting that, in addition to the TCC, several other conditions can be employed to characterize the structure of the forward map and provide convergence guarantees.
-Notable examples include range-invariance [24], convexity, and the Polyak-Lojasiewicz condition [31].
-Although the tangential cone condition is a powerful tool, it can be challenging to verify in practice; for a general verification strategy with full data, we refer the reader to [23].
-Contributions in this area span a wide range of applications, including elliptic inverse problems [18], elastography [21, 20], electrical impedance tomography [26], full waveform inversion [10], and more recently, neural network-based inverse problems [35, 1].
-In this work, we contribute to the literature by establishing the tangential cone condition for full measurements, introducing a lifted regularity strategy developed in Section 3.4 to account for realistic measured data in $L^{2}$.
-Furthermore, we extend our verification to the setting of restricted measurements.
+Weak variants of the TCC are discussed in [Kindermann17].
+Beyond the TCC, other structural conditions can also ensure convergence guarantees: range-invariance [Kaltenbacher23-rangeinv], convexity, and the Polyak–Łojasiewicz condition [Nguyen24]
+Although the TCC is a powerful tool, verification in practice can be challenging, even for specific examples.
+For a general verification strategy with full data, we refer the reader to [TCC21].
+Applications of TCC verification span a wide range: elliptic inverse problems [HoffmanWaldNguyen:2021], elastography [Nakamura:MRE21, HubmerScherzer:TCC18], electrical impedance tomography [Kindermann21], full waveform inversion [EllerRolandRieder24], and neural network-based inverse problems [ScherzerHofmannNashed, AarsetHollerNguyen23].
+Our contribution is the TCC for inertial wave inversion under full measurement by introducing a lifted regularity strategy (Section 3.4) that handles realistic $L^{2}$-data, and extending the verification to the restricted measurement regime, which is novel in this context.
+As a consequence, we obtain local unique identifiability of the unknown parameters, a property well-known to be important in ill-posed inverse problems.
 5.1 Full measurement
@@
 In this step, we observe the importance of the lifted regularity laid out in Section 3.4, which is the key to Lemma 14. This proves the TCC (45) in $B^{X}_{R}(p^{\dagger})$. ∎
-With TCC, we further obtain unique identifiability of $\Omega$ given $\gamma$ and vice versa.
-To this end, we denote by $\mathrm{Arg}(z)\in[0,2\pi)$ the principal value of the argument of a complex number $z$.
+With the TCC, we further obtain unique identifiability of $\Omega$ given $\gamma$ and vice versa. We denote by $\mathrm{Arg}(z)\in[0,2\pi)$ the principal value of the argument of a complex number $z$.
 Corollary 16 (Unique identifiability).
@@
     Assume $\gamma$ is known. Suppose that
-    $\Delta_{m}\Psi\neq 0\text{ a.e.~on }I,\quad\mathrm{Arg}(\Psi)-\mathrm{Arg}(\Delta_{m}\Psi)\notin\{-\pi,0,\pi\}\text{ a.e.~on }I.$ (44)
+    $\Delta_{m}\Psi\neq 0\text{ a.e.~on }I,\qquad\mathrm{Arg}(\Psi)-\mathrm{Arg}(\Delta_{m}\Psi)\notin\{-\pi,0,\pi\}\text{ a.e.~on }I.$ (47)
     Then $\Omega$ is uniquely determined locally in $B_{R}^{X}(\Omega^{\dagger})$.
@@
 Thus, $\mathcal{N}(F^{\prime}(\Omega^{\dagger}))=\{0\}$.
-From this, along with the characterization of the solution set given the TCC in [16, Proposition 2.1], we deduce $\Omega^{*}-\Omega^{\dagger}\in\mathcal{N}(F^{\prime}(\Omega^{\dagger}))=\{0\}$ for any solution $\Omega^{*}$, $\Omega^{\dagger}$ of the inverse problem.
+From this, along with the characterization of the solution set given the TCC in [HNS95, Proposition 2.1], we deduce $\Omega^{*}-\Omega^{\dagger}\in\mathcal{N}(F^{\prime}(\Omega^{\dagger}))=\{0\}$ for any solution $\Omega^{*}$, $\Omega^{\dagger}$ of the inverse problem.
 We thus claim the unique identifiability of $\Omega$ given full measurement.
@@
 5.2 Partial measurement including Cauchy data
-We now turn to the case of restricted measurements, where observations are limited to $I^{\prime}=(\epsilon,\pi-\epsilon)$ as described in Section 2.3.
-As a simplification, we additionally assume that Cauchy data at the boundary of $I^{\prime}$ can be measured exactly.
-We follow the analysis for the full Sun, while specifically leveraging the well-posedness and regularity results for the restricted domain established in Section 3.5, Proposition 9.
+In the last part of the analysis, we consider the restricted measurement scenario where observational latitude range is limited to $I^{\prime}=(\epsilon,\pi-\epsilon)$ as described in Section 2.3.
+For simplification, we additionally assume that Cauchy data at the boundary of $I^{\prime}$ can be measured exactly.
+We follow the analysis developed for the full observation, but specifically leveraging the well-posedness and regularity results for the restricted domain from Proposition 9 in Section 3.5.
 Theorem 17 (TCC – partial measurement).
 Consider the setting in Proposition 9, (ii). If one imposes the constraint
-$\Psi|_{\partial I^{\prime}}=\Psi^{\dagger}|_{\partial I^{\prime}}\quad\text{and}\quad\Psi^{\prime}|_{\partial I^{\prime}}=(\Psi^{\dagger})^{\prime}|_{\partial I^{\prime}},$ (46)
+$\Psi|_{\partial I^{\prime}}=\Psi^{\dagger}|_{\partial I^{\prime}}\qquad\text{and}\qquad\Psi^{\prime}|_{\partial I^{\prime}}=(\Psi^{\dagger})^{\prime}|_{\partial I^{\prime}},$ (49)
 where $\Psi^{\dagger}$ is the exact state, then for $m\neq 0$, the forward operator $F$ with partial measurements satisfies the TCC (45).
@@
 for any $p$, $p+h$ in the ball $B^{X}_{R}(p^{\dagger})$. ∎
-It is worth remarking that the constraint assumption in (46), which requires knowledge of the exact solution on the boundary, is an essential component to the proof of Theorem 17; relaxing this condition poses an interesting question for future research.
+The boundary data assumption in (49), requiring knowledge of the exact state at $\partial I^{\prime}$, is important for the proof of Theorem 17.
+Relaxing this condition presents an interesting question for future research.
+With this theoretical discussion concluded, we now turn to the numerical experiments.
 6 Numerical experiments
-We implemented the accelerated Nesterov-Landweber method to simultaneously reconstruct the scalar viscosity and latitudinal differential rotation parameters, $(\gamma,\Omega)$.
-A convergence analysis of this method has been carried out in [19, 29] under the tangential cone condition, which we have been verified for our problem in Section 5.
+This section is dedicated to numerical experiments on synthetic data, using analytical ground truths.
+We implemented the accelerated Nesterov-Landweber method to simultaneously reconstruct the scalar viscosity and latitudinal differential rotation parameters $(\gamma,\Omega)$.
+A convergence analysis of this method has been carried out in [HubmerRamlau17, Neubauer17] under the tangential cone condition, which we have been verified in Section 5.
 The Nesterov-Landweber iteration proceeds as follows: initiate $p_{-1}=p_{0}:=(\gamma_{\mathrm{int}},\Omega_{\mathrm{int}})$, then run
-$z_{k}$ $=p_{k}+\frac{k-1}{k+\alpha-1}\left(p_{k}-p_{k-1}\right)\qquad$ $\alpha=3$ (47)
+$z_{k}$ $=p_{k}+\frac{k-1}{k+\alpha-1}\left(p_{k}-p_{k-1}\right)\hskip 18.49988pt$ $\alpha=3$ (50)
 $p_{k+1}$ $=z_{k}-\mu_{k}F^{\prime}(z_{k})^{*}\left(F(z_{k})-y^{\delta}\right)$ $k\leq K(y^{\delta},\delta).$
@@
 The Hilbert space adjoints $F^{\prime}(z_{k})^{*}$ are derived in Section 4.2.
-The inversion takes place in the following function spaces: $\gamma\in\mathbb{R}$, $\Omega\in H^{2}({r{\mathbb{S}^{2}}},m=0)$, thus ensuring convergence of the iterative reconstruction.
-The numerical wave solution uses a fourth-order finite difference scheme [39] on a uniform grid of 100 points spanning the latitude range $(0,\pi)$.
-To generate ground truths while avoiding inverse crimes, we explicitly choose analytic $\Psi_{m}$$\gamma$ and $\Omega$ with Neumann (Fig.
+The inversion takes place in the function spaces: $\gamma\in\mathbb{R}$, $\Omega\in H^{2}({r{\mathbb{S}^{2}}},m=0)$, ensuring convergence of the iterative reconstruction.
+The numerical wave solver uses a fourth-order finite difference scheme [VersyptBraatz14] on a uniform grid of 100 points spanning the latitude range $(0,\pi)$.
+To generate ground truths while avoiding inverse crimes, we choose analytic $\Psi_{m}$$\gamma$, and $\Omega$ with Neumann (Fig.
 1) or Dirichlet (Fig.
 2) boundary, then explicitly obtain the corresponding source $f$.
-The discrete measurement $\underline{y}^{\delta}$ is corrupted by independent additive complex Gaussian noise with variance $\sigma^{2}$ on 100 equidistant measurement points and $\sigma$ chosen to have relative noise levels $\|\underline{y}^{\delta}-\underline{y}\|_{2}/\|\underline{y}\|_{2}\in\{0.01,0.05,0.2\}$.
-We initialize the algorithm with $\gamma_{\mathrm{int}}=-3\gamma_{\mathrm{true}}$ and $\Omega_{\mathrm{int}}=0$, corresponding to no prior knowledge of the ground truth, and iterate with a step size $\mu_{k}$ determined by backtracking line search.
-The stopping criterion is the first instance when the residual falls below the threshold $\tau\delta$ (with $\tau=1.03$), which in practice yields better results than choosing $\tau>2$ [19].
-The algorithm typically completes 200 iterations within one second on an i7-1255U CPU (4.70 GHz).
+The discrete measurement $\underline{y}^{\delta}$ is corrupted by independent additive complex Gaussian noise with variance $\sigma^{2}$ on 100 equidistant measurement points and $\sigma$ chosen to have relative noise levels $\|\underline{y}^{\delta}-\underline{y}\|_{2}/\|\underline{y}\|_{2}\in\{0.01,0.05,0.1,0.2\}$.
+We initialize the algorithm with $\gamma_{\mathrm{int}}=-3\gamma_{\mathrm{true}}$, $\Omega_{\mathrm{int}}=0$ corresponding to no prior knowledge of ground truths, and iterate with a step size $\mu_{k}$ determined by backtracking line search.
+Following a discrepancy principal, the stopping criterion is the first instance when the residual falls below the threshold $\tau\delta$ with $\tau=1.03$; this in our practice yields better results than the theoretical $\tau>2$ in [HubmerRamlau17].
+The algorithm typically completes 200 iterations within one second on an i7-1255U CPU (4.70 GHz) [^1] Code is available on Github at https://github.com/TramNguyenAca/Inertial_Waves..
 Figure 1: 1% data noise. Top to bottom: full and leaked data (filled area) at different levels: 20%, 50%, missing imaginary part.
@@
 Despite of an initial guess far from the ground truth, the reconstruction closely approximates the exact parameters.
-The reconstructed wave field also matches the true state in both real and imaginary parts.
+The approximated wave field also matches the true state in both real and imaginary parts.
 Leaked data at poles.
@@
 In our numerical experiments, we did not impose additional Cauchy boundary data in Theorem 17, and simply computed the adjoint as described in Remark 13.
-With 20% of the data leaked, the reconstruction quality remains nearly as high as with full data.
-For 50% data loss, the recovery of $\Omega$ moderately deteriorates.
+With 20% of the data leaked (row 2), the reconstruction quality remains nearly as high as with full data.
+For 50% data loss (row 3), the recovery of $\Omega$ moderately deteriorates.
 Nonetheless, the algorithm exhibits robust performance, consistent with our convergence results in Section 5.2.
 Real part measurement.
-Figure 1, row 4 considers the scenario where only the real part of $\Psi$ is measured.
-Compared to row 3 (which also has 50% data leakage, but full complex data), the absence of the imaginary component further impairs the reconstruction, most notably in missing the secondary peak of the differential rotation.
+Figure 1, row 4 presents the case where only the real part of $\Psi$ can be measured.
+Compared to row 3 (which also has 50% data leakage, but full complex data), the absence of the imaginary component further deteriorates the reconstruction.
 Noisy data.
-Finally, Figure 2 reports the results for $(\omega,m)=(1,2)$ with 20 % data leakage and different measurement noise levels.
-At 1% noise, the recovered parameters remain very close to the truth, as evidenced by small relative errors ($\||p-p_{\text{true}}\|_{L^{2}}/\|p_{\text{true}}\|_{L^{2}}$).
-At 5% noise, reconstruction quality is slightly reduced, while at 20% noise, the outcome degrades significantly.
+Finally, Figure 2 reports the results for $(\omega,m)=(1,2)$ with 20 % data leakage and different noise levels.
+At 1% noise (row 1), the recovered parameters remain very close to the truth, as evidenced by small relative errors ($\||p-p_{\text{true}}\|_{L^{2}}/\|p_{\text{true}}\|_{L^{2}}$).
+At 5% noise (row 2), reconstruction quality is slightly reduced, while at 20% noise (row 4), the outcome degrades significantly.
 7 Conclusions and Outlook
-In this paper, we examined the modeling, as well as the associated forward and inverse problems, for inertial waves on the surface of the Sun.
-Formulating the dynamics in terms of a stream function leads to a fourth-order elliptic differential equation on the sphere.
-We rigorously proved well-posedness of this equation, including explicit smallness conditions for uniqueness, as well as conditions for non-uniqueness via analytic Fredholm theory, both in agreement with empirical findings in helioseismology.
+In this paper, we investigated the modeling, as well as the associated forward and inverse problems for inertial waves on the surface of the Sun.
+Formulating the dynamics in terms of a stream function leads to a fourth-order elliptic differential equation on the sphere and its separated version for each longitudinal-wavenumber.
+We rigorously proved well-posedness of these equations, including explicit smallness conditions for uniqueness, as well as conditions for non-uniqueness via analytic Fredholm theory, both in agreement with empirical findings in helioseismology.
 Our studies represent the starting point for further theoretical and numerical studies of resonances and inertial modes, as well as the associated inverse problems.
@@
 We point out that at least for a finite number of $m$ and $\omega$, uniqueness of rotation at the unobserved latitudes seems unrealistic by dimensionality arguments.
-Next steps include the incorporation of physical effects such as radial motion, pressure, and gravity [14] into the model and inversions with solar-like parameters, including realistic convection noise [32], as well as real observational data from the Helioseismic and Magnetic Imager.
-Another promising avenue involves leveraging data-driven model discovery techniques [1, 30] to refine existing theoretical models by learning hidden laws.
+Future extensions of this work include incorporating realistic stochastic sources of solar excitation [PhilidetGizon23] within a passive imaging framework.
+Additionally, we will investigate the nonlinear equation for finite amplitude waves, which is required to interpret the time evolution of linearly unstable inertial modes [Bekki2024].
+Another promising avenue involves leveraging data-driven model discovery techniques [AarsetHollerNguyen23, nguyen25] to refine existing theoretical inertial wave models by learning hidden laws.
 Appendix A Differential geometry and function spaces on spheres
@@
 $\begin{aligned} &\operatorname{\mathbf{grad}_{\mathrm{h}}}f=\big(\operatorname{\mathbf{grad}}\tilde{f}-\frac{\partial\tilde{f}}{\partial r}\mathbf{e}_{r}\big)|_{{r{\mathbb{S}^{2}}}},&&\operatorname{div_{\mathrm{h}}}\textbf{{f}}=\operatorname{div}\tilde{\textbf{{f}}}|_{{r{\mathbb{S}^{2}}}},\\ &\operatorname{\mathbf{curl}_{\mathrm{h}}}f=\operatorname{curl}(\tilde{f}\mathbf{e}_{r})|_{{r{\mathbb{S}^{2}}}},&&\operatorname{curl_{\mathrm{h}}}\textbf{{f}}=(\operatorname{curl}\tilde{\textbf{{f}}}\cdot\mathbf{e}_{r})|_{{r{\mathbb{S}^{2}}}}\end{aligned}$ (51)
-independent of the choices of $\tilde{f}$ and $\tilde{\textbf{{f}}}$ (see [28, p.
+independent of the choices of $\tilde{f}$ and $\tilde{\textbf{{f}}}$ (see [nedelec:01, p.
 72–73]).
@@
 $\begin{aligned} &\operatorname{\mathbf{grad}_{\mathrm{h}}}f=\frac{1}{r}\frac{\partial f}{\partial\theta}\mathbf{e}_{\theta}+\frac{1}{r\sin\theta}\frac{\partial f}{\partial\phi}\mathbf{e}_{\phi},&&\operatorname{div_{\mathrm{h}}}\textbf{{f}}=\frac{1}{r\sin\theta}\left(\frac{\partial}{\partial\theta}(f_{\theta}\sin\theta)+\frac{\partial f_{\phi}}{\partial\phi}\right),\\ &\operatorname{\mathbf{curl}_{\mathrm{h}}}f=\frac{1}{r\sin\theta}\frac{\partial f}{\partial\phi}\mathbf{e}_{\theta}-\frac{1}{r}\frac{\partial f}{\partial\theta}\mathbf{e}_{\phi},&&\operatorname{curl_{\mathrm{h}}}\textbf{{f}}=\frac{1}{r\sin\theta}\left(\frac{\partial}{\partial\theta}(f_{\phi}\sin\theta)-\frac{\partial f_{\theta}}{\partial\phi}\right).\end{aligned}$ (52)
-These differential operators satisfy the identities ([28, Thm. 2.5.19]):
-$\begin{split}&\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}=0,\qquad\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}=0,\\ &\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=-\int_{{r{\mathbb{S}^{2}}}}f\operatorname{div_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s,\,\,\,\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=\int_{{r{\mathbb{S}^{2}}}}f\operatorname{curl_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s.\end{split}$ (50)
+These differential operators satisfy the identities ([nedelec:01, Thm. 2.5.19]):
+$\begin{split}&\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}=0,\hskip 18.49988pt\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}=0,\\ &\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=-\int_{{r{\mathbb{S}^{2}}}}f\operatorname{div_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s,\,\,\,\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=\int_{{r{\mathbb{S}^{2}}}}f\operatorname{curl_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s.\end{split}$ (53)
 The Laplace-Beltrami operator on $r{\mathbb{S}^{2}}$} is given by
@@
 $\textbf{{f}}=\operatorname{\mathbf{grad}_{\mathrm{h}}}\varphi+\operatorname{\mathbf{curl}_{\mathrm{h}}}\psi,$ (56)
-and $\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\varphi,\operatorname{\mathbf{curl}_{\mathrm{h}}}\psi\rangle_{L^{2}}=0$ (see [28, Eq. (5.6.24)]).
+and $\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\varphi,\operatorname{\mathbf{curl}_{\mathrm{h}}}\psi\rangle_{L^{2}}=0$ (see [nedelec:01, Eq. (5.6.24)]).
 Function spaces.
 We define Sobolev spaces $H^{s}({r{\mathbb{S}^{2}}})$ as Hilbert scale generated by the operator $(I-\Delta_{\rm h})^{1/2}$:
-$H^{s}({r{\mathbb{S}^{2}}}):=\mathrm{dom}((I-\Delta_{\rm h})^{s/2})\qquad\mbox{with}\quad\|f\|_{H^{s}}:=\|(I-\Delta_{\rm h})^{s/2}f\|_{L^{2}}$
+$H^{s}({r{\mathbb{S}^{2}}}):=\mathrm{dom}((I-\Delta_{\rm h})^{s/2})\hskip 18.49988pt\mbox{with}\qquad\|f\|_{H^{s}}:=\|(I-\Delta_{\rm h})^{s/2}f\|_{L^{2}}$
 for $s\geq 0$.
@@
 We further introduce subspaces of functions with mean 0 by
-$L^{2}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in L^{2}({r{\mathbb{S}^{2}}}):\langle f,1\rangle=0\},\quad H^{s}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in H^{s}:\langle f,1\rangle=0\}$ (54)
+$L^{2}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in L^{2}({r{\mathbb{S}^{2}}}):\langle f,1\rangle=0\},\qquad H^{s}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in H^{s}:\langle f,1\rangle=0\}$ (57)
 for $s\in\mathbb{R}$. These spaces are also complete with the same norms, but we will later define a slightly more convenient equivalent norm on $H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$.
-The following embeddings hold true (see [36, Propositions 3.2-3.3], [17, Theorem 3.5]):
+The following embeddings hold true (see [Taylor, Propositions 3.2-3.3], [Hebey, Theorem 3.5]):
 $H^{s}({r{\mathbb{S}^{2}}})\hookrightarrow H^{k}({r{\mathbb{S}^{2}}})$ compact for $s>k$ (58a)
@@
 Since $\hat{f}_{m}\otimes e_{m}\perp\hat{f}_{n}\otimes e_{n}$ in $L^{2}({r{\mathbb{S}^{2}}})$ and in $L^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ for $m\neq n$, the expression (59) induces orthogonal decompositions
-$L^{2}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}({r{\mathbb{S}^{2}}};m),\quad L^{2}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ (57)
+$L^{2}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}({r{\mathbb{S}^{2}}};m),\qquad L^{2}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ (60)
 with $\hat{f}_{m}\otimes e_{m}\in L^{2}({r{\mathbb{S}^{2}}};m)$.
 As the spherical harmonics $\{Y_{lm}:l\in\mathbb{N}_{0},m\in\mathbb{Z},|m|\leq l\}$ form a complete orthonormal basis of $L^{2}({r{\mathbb{S}^{2}}})$ and $Y_{lm}\in L^{2}({r{\mathbb{S}^{2}}};m)$, we have
-$L^{2}({r{\mathbb{S}^{2}}};m)=L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)=\overline{\mathrm{span}\{Y_{lm}:l\geq|m|\}}^{L^{2}},\qquad m\in\mathbb{Z}\setminus\{0\},$
-$L^{2}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\}}^{L^{2}},\quad L^{2}_{\diamond}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\setminus\{0\}\}}^{L^{2}},$
+$L^{2}({r{\mathbb{S}^{2}}};m)=L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)=\overline{\mathrm{span}\{Y_{lm}:l\geq|m|\}}^{L^{2}},\hskip 18.49988ptm\in\mathbb{Z}\setminus\{0\},$
+$L^{2}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\}}^{L^{2}},\qquad L^{2}_{\diamond}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\setminus\{0\}\}}^{L^{2}},$
 where $\mathrm{span}$ denotes the set of finite linear combinations.
@@
 More precisely, it follows from (51) and (54) that
-$\Delta_{\mathrm{h}}f=\sum_{m=-\infty}^{\infty}(\Delta_{m}\hat{f}_{m})\otimes e_{m}\quad\mbox{with}\quad\Delta_{m}:=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\sin\theta\frac{d}{d\theta}\right)-\frac{m^{2}}{r^{2}\sin^{2}\theta}\,$ (58)
+$\Delta_{\mathrm{h}}f=\sum_{m=-\infty}^{\infty}(\Delta_{m}\hat{f}_{m})\otimes e_{m}\qquad\mbox{with}\qquad\Delta_{m}:=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\sin\theta\frac{d}{d\theta}\right)-\frac{m^{2}}{r^{2}\sin^{2}\theta}\,$ (61)
 where by abuse of notation, we denote $\Delta_{m}(\hat{f}_{m}\otimes e_{m}):=(\Delta_{m}\hat{f}_{m})\otimes e_{m}$, thus also
@@
 We obtain $H^{s}$-orthogonal decompositions
-$H^{s}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H^{s}({r{\mathbb{S}^{2}}};m)\quad\mbox{and}\quad H^{s}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H_{\diamond}^{s}({r{\mathbb{S}^{2}}};m)$ (60)
+$H^{s}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H^{s}({r{\mathbb{S}^{2}}};m)\qquad\mbox{and}\qquad H^{s}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H_{\diamond}^{s}({r{\mathbb{S}^{2}}};m)$ (63)
 for any $s\in\mathbb{R}$, where $H^{s}({r{\mathbb{S}^{2}}};m)$ and $H_{\diamond}^{s}({r{\mathbb{S}^{2}}};m)$ are the closures of $\mathrm{span}\{Y_{lm}:l\geq|m|\}$ in $H^{s}({r{\mathbb{S}^{2}}})$ and $H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$, respectively.
@@
 Consequently,
-$\|f\|_{H_{\diamond}^{s}}:=\|(-\Delta_{\mathrm{h}})^{s/2}f\|_{L^{2}},\qquad s\geq 0$
+$\|f\|_{H_{\diamond}^{s}}:=\|(-\Delta_{\mathrm{h}})^{s/2}f\|_{L^{2}},\hskip 18.49988pts\geq 0$
 defines a norm on $H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$ which is equivalent to the $\|\cdot\|_{H^{s}}$-norm.
 We will consider this as the standard norm on $H^{s}_{\diamond}({r{\mathbb{S}^{2}}})$ as it has the convenient properties
-$\begin{aligned} &\|f\|_{H_{\diamond}^{2}}=\|\Delta_{\mathrm{h}}f\|_{L^{2}}\quad\mbox{and}\quad\\ &\|f\|_{H_{\diamond}^{1}}^{2}=\langle-\Delta_{\mathrm{h}}f,f\rangle=\langle-\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f,f\rangle=\|\operatorname{\mathbf{grad}_{\mathrm{h}}}f\|_{L^{2}}^{2}.\end{aligned}$ (61)
+$\begin{aligned} &\|f\|_{H_{\diamond}^{2}}=\|\Delta_{\mathrm{h}}f\|_{L^{2}}\qquad\mbox{and}\qquad\\ &\|f\|_{H_{\diamond}^{1}}^{2}=\langle-\Delta_{\mathrm{h}}f,f\rangle=\langle-\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f,f\rangle=\|\operatorname{\mathbf{grad}_{\mathrm{h}}}f\|_{L^{2}}^{2}.\end{aligned}$ (64)
 Defining $H^{-s}_{\diamond}({r{\mathbb{S}^{2}}}):=H^{s}_{\diamond}({r{\mathbb{S}^{2}}})^{\prime}$ for $s>0$ we have that
-$\Delta_{\mathrm{h}}:H^{t}_{\diamond}({r{\mathbb{S}^{2}}})\to H^{t-2}_{\diamond}({r{\mathbb{S}^{2}}})\quad\text{is an isometric isomorphism for all $t\in\mathbb{R}$.}$ (62)
+$\Delta_{\mathrm{h}}:H^{t}_{\diamond}({r{\mathbb{S}^{2}}})\to H^{t-2}_{\diamond}({r{\mathbb{S}^{2}}})\qquad\text{is an isometric isomorphism for all $t\in\mathbb{R}$.}$ (65)
 Defining $L^{p}({r{\mathbb{S}^{2}}};m)$ as the closure of $\mathrm{span}\{Y_{lm}:l\geq|m|\}$ in $L^{p}({r{\mathbb{S}^{2}}})$, the embeddings (58) implies the embeddings in the decomposed spaces
@@
 And this is the case if and only if there exists a polynomial $p(x,y)=\sum_{j=0}^{k}\sum_{n=0}^{j}c_{j,n}x^{n}y^{j-n}$ of degree $\leq k$ such that
-$\left|(u\circ M-p)(x,y)\right|=o\left(\sqrt{x^{2}+y^{2}}^{k}\right)\qquad\text{as }(x,y)\to 0.$ (64)
+$\left|(u\circ M-p)(x,y)\right|=o\left(\sqrt{x^{2}+y^{2}}^{k}\right)\hskip 18.49988pt\text{as }(x,y)\to 0.$ (67)
 Writing $x=\frac{\rho}{2}(e^{i\phi}+e^{-i\phi})$ and $y=\frac{i\rho}{2}(e^{i\phi}-e^{-i\phi})$ in polar coordinates and using binomial expansions, we obtain
@@
 for some coefficients $d_{j,l}\in\mathbb{C}$. As $(u\circ M)(\rho\cos\phi,\rho\sin\phi)=u_{m}(\rho)\exp(m\phi)$, the coefficients $d_{j,l}$ must vanish unless j2l=mj-2l=m, so
-$u_{m}(\rho)=\sum_{l\geq 0,m+2l\leq k}d_{m+2l,l}\rho^{m+2l}+o(\rho^{k})\qquad\text{as }\rho\to 0.$
+$u_{m}(\rho)=\sum_{l\geq 0,m+2l\leq k}d_{m+2l,l}\rho^{m+2l}+o(\rho^{k})\hskip 18.49988pt\text{as }\rho\to 0.$
 This shows that $u$ is $k$-times differentiable at $P_{+}$ if and only if $u_{m}$ satisfies the conditions ii and iii at $\theta=0$.